### v1

# TabSTAR: A Foundation Tabular Model With Semantically Target-Aware Representations

Abstract
While deep learning has achieved remarkable success across many domains, it has historically underperformed on tabular learning tasks, which remain dominated by gradient boosting decision trees (GBDTs). However, recent advancements are paving the way for Tabular Foundation Models, which can leverage real-world knowledge and generalize across diverse datasets, particularly when the data contains free-text. Although incorporating language model capabilities into tabular tasks has been explored, most existing methods utilize static, target-agnostic textual representations, limiting their effectiveness. We introduce TabSTAR: a Foundation Tabular Model with Semantically Target-Aware Representations. TabSTAR is designed to enable transfer learning on tabular data with textual features, with an architecture free of dataset-specific parameters. It unfreezes a pretrained text encoder and takes as input target tokens, which provide the model with the context needed to learn task-specific embeddings. TabSTAR achieves state-of-the-art performance for both medium- and large-sized datasets across known benchmarks of classification tasks with text features, and its pretraining phase exhibits scaling laws in the number of datasets, offering a pathway for further performance improvements.
1 Introduction
In recent years, deep learning has profoundly reshaped research and practice in computer vision [48, 63, 32, 16] and natural language processing [53, 5, 76, 15, 12]. This transformation was notably accelerated by the rise of foundation models [7, 84, 4], capable of cross-modal understanding and generalization from massive pretraining across heterogeneous data sources. Importantly, they enabled an end-to-end approach that outperformed previous modular alternatives [68, 1]. Moreover, deep learning models excel at transfer learning [87], generalizing from their pretraining data to new tasks. Their strength, combined with techniques like In-Context Learning (ICL) [12] and Parameter-Efficient Fine-Tuning (PEFT) [40], has enabled rapid adaptation to new tasks with only limited labeled data.
Despite this progress, deep learning has historically lagged behind gradient-boosted decision trees (GBDTs) on tabular data [11, 14, 57], in both classification and regression tasks [62, 10, 30, 52]. The heterogeneity of tabular data, which lacks the spatial locality of images or the sequential order of text, makes it more challenging for deep models to learn. Consequently, GBDTs have remained the de facto standard for tabular learning, offering strong out-of-the-box performance, computational efficiency, and built-in inductive biases (e.g., robustness to skewed feature distributions and automatic feature selection) that make them especially well-suited to heterogeneous datasets [30]. Nonetheless, GBDTs cannot be pretrained to reuse strong representations for downstream tasks. This limitation becomes critical in low-data settings like those often found in healthcare applications [50]. Crucially, they must rely on external embedding models to process unstructured data types like text and images, yielding fixed feature representations that cannot be finetuned for a specific prediction task.
Table 1: A binary classification toy dataset for hospital patient release outcomes. Decision is the target variable. Age (numerical), Department (high-cardinality), and Report (textual) are the features.
Age Department Report Decision
45 Cardiology Mild chest discomfort. Released
62 Neurology Complaints of headache and occasional dizziness. Hospitalized
38 Oncology Completed treatment cycle without adverse reactions. Released
55 Neurology Reports episodes of vertigo and memory lapses. Hospitalized
The emerging field of Tabular Foundation Models (TFMs) has begun addressing these shortcomings, introducing powerful cross-dataset learning strategies [80, 45, 37]. However, the flagship model TabPFN-v2 [37] still handles text inputs no more flexibly than conventional GBDTs. This design choice is not incidental; historically, tabular benchmarks have prioritized numerical datasets without free-text features, largely for ease of modeling and evaluation. A recent study [46] of mainstream tabular datasets benchmarks [24, 22, 83, 52] found that half of these datasets are more than 20 years old, being a poor representation of modern real-world data.
Real-world tabular datasets often include high-cardinality [^1] High-cardinality features are categorical columns with a large number of unique values. and free-text features [13], illustrated by a toy example in Table 1. In such datasets, free-text features (e.g., Report) carry rich semantic information critical for tasks like predicting whether a patient will be discharged from the hospital or require continued care. Yet, most models encode them in a target-agnostic manner, delegating to a generic embedding that fails to capture task-specific nuances for predicting Decision. Crucially, that same embedding would have been used for a different target variable (e.g., Treatment Cost). Similarly, categorical features with dozens of unique values (e.g., Department) are difficult to encode efficiently without external knowledge, making naive approaches brittle and limiting generalization. Importantly, the column names, which could guide the model toward more effective representations, are typically ignored. Addressing these limitations is crucial for developing tabular models that leverage semantic information, transfer knowledge from many datasets, and generalize across domains.
In this paper, we introduce TabSTAR: a novel Tabular Foundation Model with Semantically Target-Aware Representations, [^2] Code is available at https://github.com/alanarazi7/TabSTAR. designed explicitly for end-to-end handling of purely textual features. By integrating an unfrozen text encoder at its core, TabSTAR can optimize free-text feature representations, demonstrating their clear superiority over alternative frozen embedding approaches. Additionally, it introduces a novel approach of target-aware tokens, which inject semantic information about the target variable as part of the input, allowing for efficient parameter sharing and resulting in an architecture with no dataset-specific parameters (see Figure 1). TabSTAR’s training is highly efficient [^3] Pretraining within 48 hours on a single A40 GPU. Finetuning with PEFT for low memory footprint. and its performance steadily improves with more pretraining data. Empirically, TabSTAR achieves state-of-the-art (SOTA) performance on classification benchmarks containing substantial textual content, demonstrating significant advances over GBDTs and leading TFMs.
2 Related Work
This section reviews prior work in five areas relevant to our approach. We begin with deep learning methods tailored for tabular data, which were applied to a single dataset. We then discuss cross-dataset transfer learning techniques that improve generalization by leveraging related datasets. Next, we cover the field of TFMs, which aim to generalize across diverse tasks and datasets through large-scale pretraining. Finally, we review recent work on applying large language models (LLMs) to tabular data and elaborate on existing AutoML [33] multimodal solutions.
Deep Learning on a Single Tabular Dataset
Several architectures have been proposed to enhance deep learning for tabular data [67, 44, 78, 81]. TabNet [3] and TabTransformer [42] introduced attention mechanisms into tabular deep learning, while FT-Transformer [25] and its improvement [26] jointly integrated numerical and categorical features into a transformer [76]. Other novel approaches leveraged inter-example information at inference time, with SAINT [66] proposing row-level attention between examples, Non-Parametric Transformers [47] processing the entire dataset, including labels, in a single forward pass, and TabR [27] combining a k-nearest-neighbor mechanism with a traditional Multi-Layer Perceptron (MLP) architecture. Recent works such as TabM [28] and RealMLP [38] focused on refining MLPs without an attention component. Despite these innovations, single-dataset deep learning models have not yet convincingly outperformed GBDTs [62, 30, 61]. Furthermore, none of them addressed the challenge of modeling tabular datasets with rich textual features.
Cross-Dataset Transfer Learning
Deep learning was proven to shine when performing transfer learning in many machine learning domains [87]. Motivated by this success, [50, 85] proved that cross-dataset learning can boost single-dataset performance, but were limited to strict requirements such as partial overlap of feature names. To address this limitation, TransTab [79] integrated semantic understanding into feature tokenization, and XTab [86] pretrained a transformer backbone with dataset-specific parameters, proving that pretraining contributes to a stronger initialization for a downstream task. Despite their small scale, these studies demonstrated cross-dataset transfer learning’s potential, laying essential groundwork for the rise of TFMs.
Tabular Foundation Models
TFMs represent an emerging paradigm in tabular learning. While the definition is still evolving, we adopt the framing proposed by [74], which identifies key desired characteristics of TFMs: large-scale pretraining with adaptability to downstream tasks, mixed-type column support, cross-domain generalization, use of textual metadata, [^4] Contextual information such as the dataset description, column names and category names. and column-order invariance.
TabPFN [35] is recognized as the first TFM, and its successor TabPFN-v2 [37] currently sets the SOTA in tabular learning, becoming a popular approach for TFMs [58, 20]. TabPFN-v2 was the first model to consistently outperform GBDTs on medium-sized datasets, by pretraining Bayesian Prior-Data Fitted Networks (PFNs) [55] on 130 million synthetic datasets. Using ICL at inference time, it accepts up to 10,000 examples as input and predicts without updating its weights. Nonetheless, similarly to GBDTs, TabPFN-v2 uses off-the-shelf embeddings for text features, limiting its effectiveness.
CM2 [85], CARTE [45] and TP-BERTa [80] represent a shift toward semantic tabular modeling, leveraging textual signals and external knowledge at a greater scale. Unlike prior methods, these models transfer knowledge via language representations. CM2 pretrained on over 2,000 datasets, but did not focus on free-text features and used static word embeddings without further finetuning them. CARTE encodes tables as star-shaped graphs, jointly representing features by their names and values, and applies attention over the graph to capture contextual relations. While effective for high-cardinality features, it lacks exposure to longer free-text fields during pretraining and was proven useful mainly for small datasets. TP-BERTa adapts RoBERTa [51] with intra-feature attention and a tokenization scheme that maps numerical values into discrete relative-magnitude bins, to address the weakness of language models when tokenizing numbers [72]. Although it performs well, its use of dataset-specific output layers limits scalability and complicates multi-task learning. Consequently, they trained two separate models, [^5] One for classification and one for regression. A joint model for both tasks performed significantly worse. wasting potential for better cross-dataset learning. Notably, none of these approaches finetune semantic representations during downstream task training. In our work, we demonstrate that this is critical to align textual and tabular features.
Large Language Models for Tabular Data
The remarkable success of LLMs is unprecedented [12, 56]. During the past years, several research attempts have tried to combine LLMs and tabular data. One line of work is on using LLMs directly for tabular prediction, by converting tabular data into serialized text. TabLLM [34] assessed LLMs under few-shot scenarios, while Tabula-8b [23] finetuned the Llama 3-8B model extensively on tabular data. Although useful for few-shot learning, these models are computationally expensive, [^6] Llama 3-8b has orders of magnitude more parameters than TP-BERTa, which has roughly 110M parameters. suboptimal for numerical features [72, 74], and potentially compromised on widely-used benchmarks due to prior exposure during training [8]. While current generations of LLMs weren’t adopted for tabular learning, their emergent knowledge from their pretraining could be crucial when textual features are present [74, 19]. Additionally, LLMs can be used in multiple aspects of tabular learning, as they seem to be promising synthetic data generators [9, 65], useful data cleaners [6] and clever feature engineers [36].
Multimodal AutoML
Historically, textual tabular datasets have been largely overlooked in classical tabular benchmarks. However, the AutoML [33] community has made significant progress in developing multimodal solutions. In particular, AutoGluon [18] introduced the AutoML Multimodal Benchmark [60], initially focusing on text features and later evolving into AutoGluon-Multimodal [70, 69], which incorporates images as well. This powerful AutoML framework can fuse text and image foundation models with tabular models and ensemble multiple models through a meta-learning approach [21], making it one of the few systems able to refine static textual representations via joint learning. Nevertheless, this line of work should not be seen as a single model but rather as a highly optimized, production-ready system. According to the authors, it is "a collection of tricks that significantly enhance performance" [70], establishing itself as a robust approach for multimodal, multi-model tabular learning. However, this line of work remains somewhat orthogonal to the development of novel TFMs.
Figure 1: The TabSTAR architecture illustrated with our toy dataset. The model processes numerical features, textual features, and all possible target values for classification.
3 TabSTAR
In this section, we introduce TabSTAR: a Tabular Foundation Model with Semantically Target-Aware Representations. Our training framework consists of two stages: (1) Pretraining, where it is pretrained over a corpus of tabular datasets [^7] Ranging from metadata-rich, text-heavy datasets to numeric-only tables lacking column names. in a multi-task regime, mixing classification with regression tasks, then (2) Finetuning, where the pretrained model is further trained with LoRA [41] on a single downstream task. TabSTAR is designed to enable effective cross-dataset learning by applying supervised learning on the target variable in both stages. At its core, is uses an unfrozen encoder-only language model, which can invoke world knowledge acquired during the language model pretraining. [^8] Note that the language-model pretraining occurs before TabSTAR’s pretraining. Unless specified differently, the term pretraining refers to TabSTAR’s pretraining, which assumes the use of a pretrained language model. The encoder is combined with a tabular-specific architecture tailored to structured data, mitigating the known limitations of language models in tabular settings [72, 74].
TabSTAR’s architecture comprises five core modules: (1) Verbalization, mapping every feature into a textual representation composed of both the column name and value, with a special treatment to numerical features for full numerical precision; (2) Encoding, transforming semantic and numerical data into meaningful embeddings of the same dimension; (3) Fusion, integrating textual and numerical representations; (4) Interaction, modeling dependencies and relationships between different elements through self-attention and cross-element interactions; and (5) Prediction, where Interaction’s outputs are projected into a real value for regression or a probability distribution for classification. Figure 1 illustrates the architecture while Appendix A elaborates, and Appendix B discusses the training.
A key innovation of TabSTAR is the introduction of target-aware tokens, a novel approach that integrates the target variable’s identity as an input to the model. Unlike existing TFMs [37, 45, 80, 82, 86, 79], which treat the target value as a mere label, TabSTAR fuses target-awareness from the very beginning. For classification tasks, each target value is verbalized and encoded like any other feature. Then, features and target tokens interact with each other, building representations that are then used for prediction. Crucially, this target-awareness allows parameter sharing between all target tokens, which can later use a shared prediction head that maps tokens to probabilities regardless of the number of classes and their identity. By doing so, TabSTAR eliminates the need for dataset-specific components commonly found in prior work [25, 86, 80]. TabSTAR’s flexible architecture effortlessly scales [^9] Except when the number of features becomes very large, where memory limitations may arise. to any dataset size, and handles any number of classes in multiclass classification tasks.
Table 2: An illustrative verbalization of the first patient of Table 1. Each semantic feature is verbalized with its name and value. The numerical Age value 45 is standardized (mapped into z-scores, e.g., 0.27) and binned (providing a range to the verbalization, e.g., 40-50, and its quantile). The target variable Decision is mapped into its two possible elements, regardless of its original true value.
Name Value Semantic Numerical
Age 45 “Age: 40–50 (Quantile 50–60%)” $0.27$
Department Cardiology “Department: Cardiology” 0-
Report Mild chest discomfort. “Report: Mild chest discomfort.” 0-
Decision Hospitalized “Target. Decision: Hospitalized” 0-
Decision Released “Target. Decision: Released” 0-
Verbalization
All the features and each of the target values are processed into a sequence of elements. Numerical features are processed into two inputs: a numerical one and a semantic one. The numerical input is standardized using z-scores, with outlier clipping at ±3 standard deviations. In addition, they are verbalized using quantile-based binning into 10 bins, a novel approach to mitigate the precision loss inherent in language models [72]. Appendix A.1 shows a precise example and §6 discusses different verbalization strategies. In contrast, semantic features are directly verbalized by concatenating the feature name and textual value, without any numerical representation. The target variable is also included as part of the input: In classification tasks, each of the $C$ possible values is represented by an element, constant for every example, while the true value remains hidden. For regression tasks, a single element is verbalized, carrying only the target name. Table 1 shows a toy dataset of patient records and outcomes and Table 2 shows the verbalization for the first patient.
Encoding
We employ a pretrained e5-small-v2 [77] embedding model for semantic encoding, chosen for its strong performance on the MTEB benchmark [54] with a relatively modest parameter count. By unfreezing half of its layers, the representations are optimized for predicting the target variable, which leads to a significant impact on TabSTAR performance (see §6). Each verbalization element is encoded independently into a semantic representation, with attention applied between tokens within each sequence element. In parallel, we encode standardized numerical values by projecting them into the same dimension using a small MLP. For the patient in Table 2, this results in a numerical embedding for Age alongside semantic representations for each of the five verbalizations.
Fusion
To obtain a unified representation for each sequence element, we apply a fusion block consisting of a single encoder-only transformer layer. For each numerical feature, the block attends over its numerical and semantic embeddings, producing a fused representation. In our running example, the representation of Age now jointly captures both its semantic context (the fact that the value represents age) as well as its numerical value (the patient’s age, 45, or 0.27 after standardization).
Interaction
The fused, semantically-rich and numerically-grounded representations of all elements interact via a 6-layer Transformer encoder [76]. Each input element is now a token, with feature tokens and target tokens all attending to each other. Unlike standard language models which integrate positional encoding, the Interaction module’s inputs are order-invariant, a desiredatum for TFMs, as defined by [74]. The encoder produces contextualized representations for each target value. In our example, this yields dedicated embeddings for the Release and Hospitalization target values. The role of these representations is to carry information about the likelihood of each value to be the true value.
Prediction
TabSTAR is designed for cross-dataset learning, with shared regression and classification heads used during both pretraining and finetuning. For classification, each of the $C$ target tokens is processed independently through the same classification head, which projects them to scores. We then apply a softmax over all the possible values to yield a probability distribution. Crucially, the fact that target tokens for every class in every dataset share the same classification head allows efficient parameter sharing, flexibly supports any number of output classes, and removes any need for dataset-specific parameters. This is not only efficient during pretraining, but also provides a better initialization for finetuning. In our example, both the Released and Hospitalized tokens go through the same classification head, which maps them from representations to logits. Applying softmax yields predicted probabilities. For regression tasks, a single target token is projected into a real value.
4 Experiments
The TabSTAR Pretraining Corpus
While TabSTAR could be pretrained on a massive scale, for this work we limit ourselves to a modest pretraining corpus focusing on classification, as we believe that TabSTAR’s inductive biases are best suited to shine in this task. We manually curate a pretraining corpus of 350 high-quality tabular datasets (253 classification, 97 regression), in a tedious process in which we uncover numerous duplications in the most popular tabular repositories, OpenML [75] and Kaggle, [^10] https://www.kaggle.com/datasets as elaborated by [73]. Our datasets are sourced from popular benchmarks [24, 45, 22, 21, 30, 60, 52, 31], which have almost non-existing representation of textual tabular datasets. Thus, we furthermore increase our corpus, focusing on classification datasets with rich semantic content. See Appendix C for more details.
Benchmark
Tabular datasets with free-text have seen little prior research, and accordingly, benchmarks are incredibly rare. Therefore, we consider all the possible datasets from AutoML Multimodal Benchmark [60], from the analysis about free-text and high-cardinality features by [31] and from the CARTE paper [45]. After a deduplication process we end up with 50 datasets. However, there are two important limitations: first, the benchmark is heavily biased towards regression, with 36 datasets in total. Secondly, 29 out of these 36 datasets were solely contributed by the CARTE benchmark, which focuses more heavily on high-cardinality features as it was pretrained over knowledge graphs. While our main motivation is classification tasks with textual features, we decide nevertheless to evaluate on the full set of 50 datasets although it heavily biases towards regression problems and high-cardinality features, rather than classification and free-text (see Appendix D).
Baselines
We compare TabSTAR against a diverse set of baselines. For tree-based methods, we evaluate CatBoost [57], XGBoost (XGB) [14], and Random Forest (RF) [11] with the default configuration proposed by [25]. For CatBoost and XGBoost we consider a tuned version, where hyperparameters are optimized separately for each task using random search with 5-fold cross-validation under a 4-hour budget on 8 CPU cores. Among TFMs, we evaluate TabPFN-v2 [37] and CARTE [45]. For CARTE, we tune only the learning rate for each task, following the original paper. Since the public TabPFN-v2 model does not support text, we use their closed-sourced API client. [^11] https://github.com/PriorLabs/tabpfn-client For models lacking native support for textual features, we embed text using e5-small-v2 [77], allowing a fair comparison. For more details about the hyperparameters for each baseline as well as exclusion of models such as TP-BERTa due to potential leakage concerns, see Appendix E.
Experimental Setup
Each of the 50 datasets in the benchmarks is evaluated with 10 random train-test splits (90% training, 10% testing), resulting in 500 runs per model. While 30 of the datasets have more than 10,000 examples, the evaluated TFMs have strict limitations. TabPFN-v2 employs ICL and thus receives as input at most 10,000 examples. Although CARTE imposes no size cap, [^12] In their own paper, CARTE was evaluated only over up to 2,048 examples, without scaling guarantees. it suffers from inefficiency and no preset configuration, requiring six learning-rate trials and totalling 6,000 slow GPU runs. Because of these important limitations, we consider two experiment conditions: (1) 10K: Each model is trained [^13] While TabPFN-v2 isn’t technically trained, we adopt this term for conciseness. over at most 10,000 training examples, and (2) Unlimited: We add TabSTAR-Unlimit, CatBoost-Tuned-Unlimit, and XGBoost-Tuned-Unlimit and evaluate them on the full version of the 30 datasets, [^14] We technically cap the amount of examples to 100,000 for computational efficiency. while retaining 10K baselines as weaker reference points. We exclude the untuned GBDTs and keep the same number of models as in the 10K condition.
The TabSTAR Training
To maximize the value of cross-dataset learning, instead of pretraining TabSTAR once, we create five dataset splits. Each variant is pretrained on the 350 pretraining datasets and 40 of the benchmark datasets, while the other 10 serve exclusively as its test set. Crucially, the whole collection was carefully curated to prevent any data leakage from duplicate or overly similar datasets. For finetuning, while dataset-specific hyperparameter tuning can boost performance, we believe that robust out-of-the-box defaults are essential for TFMs and their evaluation, following Tab-PFN-v2’s approach. Therefore, we use a default hyperparameters configuration that was found robust over a disjoint set of tabular datasets, as detailed in Appendix B.2.
5 Results
We evaluate each model using AUROC (classification) and $R^{2}$ (regression) as metrics. Following [37], we normalize scores per dataset split to the $[0,1]$ range, using the best and worst model performance as anchors. [^15] For a single run, the best model gets 1, the worst gets 0 and the rest are linearly scaled accordingly. The normalized scores are averaged across all runs, with 95% CIs. Performance for all models on both conditions are shown in Figure 2 (classification) and Figure 3 (regression). Besides their example limit, TabPFN-v2 cannot process 4 datasets (more than 10 target classes or more than 500,000 cells) and CARTE cannot handle 15 (a bug in their PowerTransformer implementation). Reported averages for these models are computed only over the datasets where evaluation is feasible. Appendix F expands on dataset level performance, head-to-head comparisons and running times.
Figure 2: Comparison of normalized scores with 95% CIs between TabSTAR and baseline models in classification tasks, evaluated on up to 10,000 examples (left) and above 10,000 (right).
In classification problems, TabSTAR consistently achieves SOTA performance. This is evident both when restricting the dataset size to 10,000 examples and when using larger datasets in the unlimited condition. For the 10K condition, TabSTAR achieves a 0.809 score, performing better than TabPFN-v2 (0.783) and significantly better than GBDTs (0.756 CatBoost-Tuned, 0.744 XGB-Tuned). When analyzing head-to-head comparisons (Appendix F.2), TabSTAR outperforms TabPFN-v2 (7/11 datasets), XGB-Tuned (10/14) and CatBoost-Tuned (11/14). For the Unlimited condition, TabSTAR-Unlimit achieves a 0.874 score, significantly above the second-best CatBoost-Tuned-Unlimit with 0.734. Importantly, all Unlimit variants surpass the 10K ones, emphasizing the importance of scaling.
Figure 3: Comparison of normalized scores with 95% CIs between TabSTAR and baseline models in regression tasks, evaluated on up to 10,000 examples (left) and above 10,000 (right).
Although regression is not our main focus, TabSTAR achieves competitive results in the 10K condition, but clearly does not set the SOTA. Surprisingly, while TabPFN-v2 is superior, it significantly underperforms compared to GBDTs which dominate this category. This emphasizes the need for better modeling of textual tabular learning, especially since TabPFN-v2 has shown remarkable performance in non-textual tabular datasets, and CARTE set the SOTA for small datasets. When analyzing the Unlimited variants, TabSTAR scales well, and surpasses other TFMs which cannot scale, but the gap from GBDTs remains significant. §7 discusses this limitation and suggests promising directions for future generations of TabSTAR to achieve SOTA in regression as well.
6 Analysis
We analyze the factors contributing to TabSTAR’s strong performance by addressing three key research questions: Q1: How important is the encoder language model unfreezing? Q2: Does the number of datasets during pretraining contribute to the downstream task performance? and Q3: How do different verbalization methods of numerical features impact performance?
To answer these questions, we pretrain several variants of TabSTAR for each analysis, limiting ourselves to a subset of the tabular datasets used for the main experiment (see §4). Specifically, each variant is pretrained over 256 datasets [^16] Except for variants of Q2, which analyze the effect of number of datasets on pretraining. including 30 datasets from our benchmark, and evaluated over the remaining 20 datasets (12 regression, 8 classification). This reduced setup allows leveraging transfer learning and exploiting our corpus, without the burden of training multiple folds per variant. Results are reported with the same normalized metric used in §5, scaling the performance to the $[0,1]$ range. Appendix G.1 lists the 20 datasets used for evaluation along with per-dataset results.
Figure 4: Performance as a function of the number of encoder layers unfrozen: Validation loss during TabSTAR’s pretraining (left) and normalized scores with 95% CIs on the downstream tasks (right). Unfreezing even a single encoder layer significantly improves the performance of TabSTAR.
Q1: The Role of the Encoder Unfreezing
To investigate whether unfreezing layers of the textual encoder impacts performance, we conduct experiments where we unfreeze varying numbers of the 12 encoder layers during both TabSTAR’s pretraining and finetuning stages. [^17] For each variant, the number of unfrozen layers remains the same in both pretraining and finetuning. Figure 4 shows the validation loss during TabSTAR pretraining (left) and the normalized score on the downstream tasks (right) as a function of the number of unfrozen encoder layers. Notably, unfreezing even a single encoder layer significantly outperforms using static embeddings. Further substantial improvements are observed as more layers are tuned, with the best results achieved when unfreezing 6 layers. While unfreezing 9 layers shows lower performance, it is plausible that adding more datasets to the pretraining phase will affect this finding. See Appendix G.2 for more details.
Table 3: Normalized score with 95% CIs by the number of datasets used during TabSTAR pretraining.
Pretraining Datasets 0 16 64 256
Classification 0.352 ± 0.086 0.450 ± 0.084 0.558 ± 0.086 0.786 ± 0.076
Regression 0.338 ± 0.073 0.395 ± 0.068 0.642 ± 0.066 0.811 ± 0.055
Q2: The Effect of Pretraining
To evaluate the impact of pretraining on TabSTAR’s downstream performance, we compare a pretrained version of TabSTAR with a version that was finetuned from scratch. [^18] Since LoRA underperforms on random weights, we finetune the entire non-pretrained model. In line with previous work [86, 82], the pretrained model performs significantly better, highlighting the critical role of transfer learning for TabSTAR’s success. To further investigate the effect of the number of pretraining datasets on downstream task performance, we train two additional versions: one pretrained on 16 datasets and another on 64 datasets. As shown in Table 3, increasing the number of pretraining datasets consistently improved performance in both classification and regression tasks. Notably, the substantial gain in regression tasks suggests that TabSTAR’s downstream performance on §5 could improve with more pretraining data (see Appendix G.3).
Q3: Numerical Verbalization
A key challenge in integrating language models with numerical data is determining how to best represent numerical values within a linguistic framework. While some semantic tabular methods omit numerical features from the verbalization [79, 82], TP-BERTa [80] introduced Relative Magnitude Tokenization [80], which encode numerical information through non-semantic special bin tokens. In constrast, TabSTAR injects semantic numerical information into the verbalization of numerical features, as illustrated in Table 2. To quantify the effect of our novel verbalization, we explore two thinner variants: (1) Name + Bin, which excludes the quantile information, and (2) Name, which omits numeric information entirely and verbalizes the feature name only. Appendix G.4 shows an illustrative example for each variant and presents the full results. As demonstrated in Table 4, our findings reveal that incorporating numerical information significantly enhances performance, highlighting the importance of balancing numerical precision with a representation format that aligns with the language model’s parametric knowledge.
Table 4: Normalized score with 95% CIs by the numerical verbalization method.
Verbalization Method Name Name + Bin TabSTAR
Classification 0.386 ± 0.095 0.544 ± 0.093 0.593 ± 0.097
Regression 0.386 ± 0.081 0.584 ± 0.076 0.596 ± 0.079
7 Discussion and Conclusion
We introduce TabSTAR, a Tabular Foundation Model with Semantically Target-Aware Representations, which integrates textual features through an unfrozen pretrained encoder. In addition, its novel target-aware tokens enable efficient cross-dataset generalization without dataset-specific parameters. Despite limited pretraining data and a relatively small text encoder [77], TabSTAR sets the SOTA in tabular classification with textual features, significantly surpassing GBDTs and leading TFMs.
Since scaling laws in data and model size have proven themselves for LLMs [43] and TabSTAR improves with the number of pretraining datasets (see §6), future work should scale TabSTAR across both model and data dimensions. For model scaling, we envision a family of model sizes, common for LLMs [29, 71, 49], that will allow a trade-off between quality and costs. Data scaling might leverage self-supervised learning [59] over large-scale table corpora [17], or realistic synthetic tabular data generators [9], which have proven successful [37, 2]. At scale, it could potentially unlock few-shot learning capabilities and develop automatic feature-engineering skills [36].
Beyond scaling, TabSTAR’s semantic approach has tremendous potential to explicitly include world knowledge, by leveraging LLMs which to date have had limited impact on tabular learning. As a few motivating examples, LLMs could improve TabSTAR’s numerical verbalization binning approach by providing semantically informed thresholds, or provide explicit, contextual world knowledge that could be injected as a strong prior in small data scenarios. While these directions seem like plausible research paths, they come with a risk of data leakage due to the memorization properties of LLMs [8]. Evaluating TFMs fairly while keeping benchmarks uncontaminated would be an important enabler for tabular research. As a step in this direction, we are releasing several TabSTAR variants, each with a different dataset withheld during pretraining, ensuring that for every dataset there is a TabSTAR model that has never seen it. We urge fellow researchers to adopt this approach in their own work.
While TabSTAR sets a new bar in classification, its regression results lag behind GBDTs, which outperform other TFMs as well. This gap could be narrowed through additional scaling, and also by exploring regression-via-classification techniques like [37, 2]. Furthermore, TabSTAR has not been extensively evaluated in few-shot scenarios and in purely numerical datasets. [^19] Partly because of the computational burden of tuning baselines, and the lack of objective leaderboards [73]. In addition, it demands more compute compared to GBDTs, and it may struggle with memory constraints on datasets containing hundreds of features. Despite these limitations, TabSTAR offers a promising pathway toward improving performance on tabular datasets with textual fields, common in industries with high social impact (e.g., healthcare, education), or with significant economic value (e.g., banking, manufacturing). We believe TabSTAR paves the way for a new generation of semantically enriched tabular models, and we welcome the research community’s innovations built on this foundation.
Acknowledgments and Disclosure of Funding
We thank Omri Feldman for brainstorming since the very beginning; Elad Hoffer and Ofir Lindenbaum for consulting and feedback; David Holzmüller and Myung Kim for supporting evaluations; and Noah Hollmann, Léo Grinsztajn, and the Prior Labs team for providing extensive access to TabPFN-v2.
Appendix A Architecture
This appendix provides additional technical details for the architecture introduced in §3. First, we discuss the verbalization module; next, we formally describe the architecture step-by-step; and finally, we present selected experiments on the TabSTAR architecture.
A.1 The Verbalization Module
TabSTAR’s verbalization module standardizes heterogeneous tabular inputs by converting each column, whether predictive feature or target variable, into templated text blocks. We first describe the detection of column types, and then detail the processing steps for each type.
Feature Detection
We classify each column as either numerical, referring to quantitative values, or semantic, referring to textual values including categorical and boolean fields encoded as text. We rely on heuristics involving both the primitive data type (e.g., string, float) and human annotation (e.g., OpenML metadata). However, real-world datasets pose challenges, as numerical features can often be stored as strings (e.g., “35 years”, “unknown age”) or may lack inherent order (e.g., country calling codes). Leveraging LLMs for contextualized data cleaning can be a promising direction [6].
A special case is the handling of timestamp and date columns. Similarly to [39], we rely on skrub’s DatetimeEncoder [^20] https://skrub-data.org/ to detect datetime columns and decompose each one of them into a set of new features. Each extracted feature then undergoes its own processing: For example, the weekday is treated as semantic, while the total seconds since the Unix epoch is treated as numerical. Integrating date features more holistically remains an open research question.
Table 5: Illustrative verbalization of a numerical feature (Age) with 10 bins. Examples outside the range and missing values are considered as special bins.
Bin Range Example Value Illustrative Verbalization
Lower than 18 17 Age: Lower than 18 (Quantile 0%)
1 18–23 20 Age: 18–23 (Quantile 0–10%)
2 23–27 25 Age: 23–27 (Quantile 10–20%)
3 27–31 29 Age: 27–31 (Quantile 20–30%)
4 31–35 33 Age: 31–35 (Quantile 30–40%)
5 35–40 38 Age: 35–40 (Quantile 40–50%)
6 40–45 42 Age: 40–45 (Quantile 50–60%)
7 45–51 48 Age: 45–51 (Quantile 60–70%)
8 51–58 55 Age: 51–58 (Quantile 70–80%)
9 58–67 63 Age: 58–67 (Quantile 80–90%)
10 67–87 83 Age: 67–87 (Quantile 90–100%)
Higher than 87 93 Age: Higher than 87 (Quantile 100%)
Unknown Age: Unknown Value
Numerical Features
[…]
Semantic Features
Semantic features are sanitized (e.g., normalizing whitespaces) and verbalized using the template presented in Table 6. Missing values are mapped to “Unknown Value”, just like for numerical features. If a text exceeds the model’s context window (512 tokens for e5-small-v2), it is naively truncated to fit it. This limitation is far more pronounced for methods that serialize the entire example into a single textual sequence [80], thereby dramatically reducing the effective context size.
Target Variables
The verbalization templates for the target values are prepended to every example. For classification tasks, each possible label is verbalized, while for regression we verbalize a single element consisting solely of the feature name. Employing a binning strategy to treat regression as a classification task is a future work direction, as discussed in §7. For regression tasks, target values go through the same standardization with outlier clipping as numerical features, being used solely as the ground truth without going through the input.
Table 6: Verbalization templates for semantic features and target values.
Element Type Verbalization Template
Predictive feature "Predictive Feature: {feature_name}"
"Feature Value: {feature_value}"
Classification target "Target Feature: {target_name}"
"Feature Value: {target_value}"
Regression target "Numerical Target Feature: {target_name}"
A.2 The Annotated TabSTAR
Table 7 describes the number of parameters per component in the TabSTAR architecture, when using e5-small-v2 [77] as the text encoder. It has approximately 47.26M parameters, most of which come from the text encoder. When unfreezing 6 layers of the text encoder, about 24.70M parameters are tuned, with the remaining 11.92M embedding parameters and 10.65M layer ones being kept frozen.
Table 7: Parameter counts for TabSTAR components
Module # Parameters
Encoding: Semantic 33,360,000
Encoding: Numerical 296,832
Fusion 1,774,464
Interaction 10,646,784
Prediction 1,185,794
To describe the architecture more precisely, we start by defining the dataset formally. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ denote a tabular dataset with $n$ examples. Each example $x_{i}=[x_{i1},\dots,x_{im}]$ has $m$ features. The target variable $y_{i}$ is either continuous (regression) or discrete (classification) taking one of $C$ classes. For simplicity, we describe the architecture at the example level, though all computations are actually carried out on mini-batches of size $B$. The batches are always drawn from a single dataset in both pretraining and finetuning, removing any need for padding.
Verbalization
We denote by $t$ the number of target entries where $t=C$ for classification and $t=1$ for regression. We then form a raw sequence of length $e=t+m$ by listing the $t$ target values, followed by the $m$ feature entries. Each element $j$ in this sequence is then verbalized into a semantic string $s_{j}$ and a numerical value $n_{j}$, set to be the clipped z-score for numerical non-missing features, and zero otherwise. The example is thus represented by parallel sequences $(\mathbf{s},\mathbf{n})$ of length $e$.
Encoding
Each semantic string $s_{j}$ and numerical value $n_{j}$ are projected into a $d$-dimensional vector. Semantic strings are encoded with an encoder-only language model (e5-small-v2 [77]). Each string is tokenized, passed through the model, and pooled to produce its final embedding. This process is independent between elements, i.e. the attention is at the token level within a single element. In parallel, each numeric value is fed through a two-layer MLP that first projects from 1 to $2d$ dimensions, applies a ReLU and dropout, and then projects back down to $d$. This produces matching $d$-dimensional embeddings for each of the $e$ elements, ready to be fused.
Fusion
To unify semantic and numerical embeddings into a single representation, we apply a single-layer Transformer Encoder [^21] With 2 attention heads, a feed-forward hidden size of $4d$, dropout 0.1, and ReLU activation. over each element’s pair of vectors. Concretely, for each element we stack its $d$-dimensional text and numeric embeddings and feed them through the encoder layer. For every element, the attention is applied between its two representations, and we average the two outputs to produce one fused $d$-dimensional embedding. This yields a final sequence of length $e$ and dimension $d$, which will serve as tokens for the Interaction block.
Interaction
The fused sequence of $e$ tokens is processed by a standard Transformer Encoder with model dimension $d=384$, $L=6$ layers, $6$ attention heads per layer, feed-forward size $4d$, dropout $0.1$, ReLU activation and using a pre-norm configuration. Unlike in language modeling, feature ordering is irrelevant, so no positional encodings are used. The encoder produces contextualized embeddings for every position, and we retain the $t$ target embeddings for the prediction.
Prediction
We initialize two identical MLP heads, one for regression and one for classification. Each of them consists of a hidden layer of size $4d$ (with ReLU activation) followed by a linear projection to a single output. For each dataset, we choose the relevant head and process the $t$ target token embeddings. For classification, we independently feed each one of the $t=C$ target tokens to the classification head to obtain a score (logit) per class. Notably, the same head is shared across classes and datasets. We apply softmax over these scores, yielding a probability distribution regardless of the number of classes. For regression, the single target token is projected into a real value. Note that the heads is shared between datasets, as regression outputs are always clipped z-scores.
A.3 Architecture Experiments
In this section we explore the effect of different design choices for TabSTAR’s architecture. For each experiment, we only vary the parameter of interest, keeping everything else fixed. We follow the same pretraining regime as in Appendix B.1, except that for computational efficiency we train only 25 epochs (instead of 50) with 128 pretraining datasets (instead of 390). We evaluate each variant relying solely on pretraining performance, as an approximation for downstream task performance. We acknowledge that our conclusions might depend on this limited scale, hence we discuss a subset of the experiments briefly to reflect the depth of our work and inspire future research.
The Fusion Block’s Mechanism
For the fusion block, we consider two simpler alternatives to the attention mechanism, both of then underperforming: (1) Concatenation, by concatenating the semantic and numerical $d$-dimensional vectors into a $2d$-dimensional vector, and projecting them back via an MLP, and (2) Multiplication, by multiplying the semantic representation directly with the numerical value [^22] After rescaling it to be centered around 1 rather than 0, using a learned scaling factor. in a straightforward, parameter-free manner as in [79, 82].
The Number of Interaction Layers
We experiment with the number of encoder layers, and observe that 3 yields anecdotally worse performance than 6, with lower parameter count. Nevertheless, we prioritize a deeper network as for datasets with very complex relationships we believe that this might be beneficial. Additionally, we try a 9-layer variant which performs significantly worse, while also increasing the parameter count.
Row-Level Attention
We experiment with adopting the architecture proposed by SAINT [66], which adds row-level attention to each encoder layer. Similar concepts are also employed by models that get the input the whole dataset, labels included [37, 47]. We run experiments with 2, 4 and 6 layers as they are roughly equivalent in parameter count to 3, 6, and 9 layers without row-attention. We observe no substantial gain, and thus we prioritize the simpler solution, as row-level attention is sensitive to the batch size and adds complexity to inference time.
Appendix B Training
In this section we elaborate on the two stages of TabSTAR’s training: pretraining and finetuning, presented in §3. As in Appendix A.3, we summarize key pretraining experiments.
B.1 Pretraining
TabSTAR is pretrained employing supervised learning in a multi-task regime, jointly learning regression, binary and multiclass classification tasks. The parameters of the architecture are fully-shared, without any need for dataset-specific parameters. Every example during the pretraining updates all the model’s weights, with the sole exception of the prediction head, for which every example uses its respective head depending on the nature of the task (classification or regression).
Sampling
For computational efficiency, each dataset is subsampled once before the pretraining. At the example level, we sample up to 300,000 examples from each dataset, stratified by the target variable for classification tasks. Since we only use a fraction of each dataset for each pretraining epoch, this decision has negligible influence. In addition, we randomly sample up to 200 features per dataset. While straightforward, this decision is suboptimal as feature importance isn’t taken into consideration. As this work does not focus on wide-feature datasets, we consider this trade-off acceptable. Importantly, this setup is enabled during finetuning as the TabSTAR architecture is agnostic to the number of features. We split each dataset into train-validation splits (95%-5%), [^23] We choose only 5% for efficiency, as we use hundreds of pretraining datasets. without any need for test splits, and cap the validation set at a maximum of 1,000 examples used for evaluating pretraining performance.
Batching
Every epoch, we randomly sample up to 2,048 examples from each dataset in mini-batches of 32, and shuffle all the batches. We conduct gradient accumulation and update the model every 4 steps to reduce the chances of a single update being dominated by a single dataset, so the global batch size is effectively 128. Appendix B.3.1 elaborates on the effect of batch size.
Metrics
Our loss function is cross-entropy for classification, and MSE for regression. With standardized targets, $R^{2}\approx 1-\text{MSE}$, although this equivalence is degraded by clipping targets in preprocessing. We train with mixed-precision and apply gradient clipping to stabilize training without task-specific weights, with Appendix B.3.2 discussing the limitations of this approach. We use as metrics AUROC for classification and $R^{2}$, so for each task the optimal metric value is 1. We average performance across all datasets into a single metric that reflects the pretraining performance.
Training
We pretrain for 50 epochs with the OneCycleLR [64] optimizer, with warmup during the first 5 epochs (10%) and cosine annealing. Early stopping is conducted after 3 epochs without improvement on the pretraining metric. The weight decay is set to 0.001, and a max learning rate of $lr=5\times 10^{-5}$ is applied uniformly across all layers. Appendix B.3.3 discusses experiments with differential learning rate. Pretraining running time varies depending on the number of epochs and the included datasets. The full models (390 datasets) reported in §5 train for less than 48 hours on a single NVIDIA A40 GPU (48GB memory), and we believe that this could be optimized much further.
B.2 Finetuning
We finetune downstream tasks using LoRA’s implementation of the peft package [^24] https://github.com/huggingface/peft. We use a rank of $r=32$, set $\alpha=2r=64$ and $dropout=0.1$. We employ the same scheduler as in the pretraining phase, with the only difference being that we set $lr=0.001$, and increase the patience parameter for early stopping to 5. We apply a train-test split of 90%-10% and sample a validation set of 10%. As opposed to the pretraining, all batches are drawn from the same dataset. Therefore, we observe no effect from changing the mini-batch size when keeping the global batch size fixed to 128. We tune 1,597,440 out of TabSTAR’s 47,263,874 parameters (3.4%).
Finetuning hyperparameters are selected by pretraining TabSTAR over 256 datasets and performing grid-search over a held-out set of 25 downstream tasks disjoint from the 50 datasets in the benchmark evaluated in §4. The search space is presented in Table 8, and we observe that average performance is relatively robust across this space. An interesting observation is that decreasing the number of parameters by setting $r=16$ mildly hurts performance, but it has upsides on memory and latency aspects, allowing a future trade-off exploration. As a final note, we argue that providing a strong default configuration for TFMs is crucial for evaluating them, but for real-world applications, it is still recommendable to find the best hyperparameters tailored to the downstream task.
Table 8: LoRA hyperparameter tuning grid search for TabSTAR’s finetuning.
Hyper-parameter Search Space
LoRA rank ($r$) 16, 32, 64
Learning Rate 0.0005, 0.001, 0.002, 0.005, 0.01
Dropout 0, 0.1
The only experiment in this paper where we employ full finetuning instead of LoRA is for the non-pretrained variant discussed in the analysis in §6. For this variant we fully finetune the pretrained model on each downstream task. Compared to the pretraining, we use $lr=2.5\times 10^{-5}$ and increase the patience to 5. These hyperparameters are lightly tuned using the same procedure as for LoRA, and we observe that fully finetuning the model achieves comparable performance, except for small datasets, where training is more prone to overfitting.
B.3 Pretraining Experiments
In this section, we briefly elaborate on some experiments performed over TabSTAR’s pretraining protocol. As in Appendix A.3, we highlight only a subset of them in a high-level manner.
B.3.1 Batch Size
During pretraining, we use a mini-batch size of 32, each of them drawn from a single dataset. Since we train with gradient accumulation and a global batch size of 128, varying the batch size affects the diversity of a single model update: lower batch sizes are likely to be exposed to more datasets. We decrease the batch size to 16 and 8 and observe an improvement at the cost of slower training. An interesting direction for future work is moving to mixed-datasets batches, which require more complex implementation but might benefit from more regularized learning. Such approach, however, goes against row-level attention methods and ICL, as discussed in Appendix A.3.
B.3.2 Loss Weights
Pretraining the model over hundreds of datasets in a multi-task regime presents a key challenge: the loss scale of each dataset can vary substantially, depending on task difficulty or task type. For example, a multiclass classification task with dozens of classes will naturally yield a higher average loss than a binary task. These dynamics can also shift during training. Our default approach naively averages the loss across all datasets, which risks over-weighting tasks for potentially arbitrary reasons.
To address this, we explore two alternative weighting strategies: (1) Assigning a constant weight per task type, accounting for the number of classes in classification tasks, and (2) Normalizing each dataset’s contribution by the best loss achieved by CatBoost [57] when being fitted to that dataset. While these strategies better reflect task-specific characteristics, they hardly impact performance and introduce additional complexity. Notably, adjusting loss weights across tasks impacts metric interpretability, as each weighting scheme implicitly optimizes a different objective.
We do not explore more sophisticated methods such as learning per-dataset weights, as these often require mixed-dataset batches and introduce additional learnable parameters. We believe, however, that multi-task pretraining over tabular datasets remains an open and important research question.
B.3.3 Differential Learning Rate
TabSTAR’s weights initialization is not balanced: the textual encoder is a pretrained embedding model while the rest of the architecture parameters are randomly initialized. To counteract this imbalance, we experiment with using differential learning rates for the textual encoder layers, and experiment with scaling it by a factor of 0.5 and of 0.75. To our surprise, this decision hurts performance, so we stick to a uniform learning rate across all layers.
Appendix C Training Datasets
In this appendix we briefly expand on the pretraining corpus elaborated in §4. It is composed of 350 datasets, spanning all the datasets appearing in AMLB [24], OpenML-CTR23 [22], TabZilla [52] and the ones presented by Grinsztajn [30]. After deduplication, [^25] And the exclusion of the fifa dataset, which is included in the benchmark. this results in 152 datasets (94 classification, 58 regression). Interestingly, only 6 of these 152 datasets have free-text or high-cardinality features. We manually add datasets from OpenML [75] and Kaggle, as well as from the AutoML-Benchmark-Train [21] corpus, and achieve a total of 350 datasets, with 49 textual datasets.
Table LABEL:tab:cls_pretrain_metadata details the 253 classification datasets and Table LABEL:tab:reg_pretrain_metadata the 97 regression ones. We elaborate the Dataset name, the number of examples $n$, the number of features $m$, and the number of classes $C$ for classification. In addition, we mark datasets that belong to one of the benchmarks, and the ones that have text features. Importantly, the textual flag is quite permissive, as it includes features with relatively short texts or potentially low predictive power (e.g., people names or addresses).
Table 9: The 253 Classification Datasets of the Pretraining Corpus, with their $n$ examples, $m$ features, $C$ classes, presence in a benchmark ($B$) and whether they are textual ($T$).
Dataset $n$ $m$ $C$ B T
KDDCup99 4,898,422 40 20
mimic_extract_los_3 4,155,270 17 68
Online-P2P-Lending 2,875,146 16 5
sf-police-incidents 2,215,023 8 2
physionet_sepsis 1,552,210 42 2
poker-hand 1,025,009 10 10
Higgs 1,000,000 28 2
BAF_base 1,000,000 30 2
Credit_Card_Fraud_ 1,000,000 7 2
Harry-Potter-fanfiction-data 648,493 13 4
porto-seguro 595,212 57 2
covertype 581,012 54 7
AVIDa-hIL6 573,891 3 2
airlines 539,383 7 2
HolisticBias 472,991 14 4
albert 425,240 78 2
DBPedia 342,781 3 219
hcdr_main 307,511 120 2
Mental_Health_Dataset 292,364 16 5
Kuzushiji-49 270,912 784 49
spoken-arabic-digit 263,256 14 10
cdc_diabetes 253,680 21 2
skin-segmentation 245,057 3 2
LT-Vehicle-Loan-Default-Prediction 233,154 38 2
Churn_Telco_Europa 190,776 17 2
ldpa 164,860 6 11
Give-Me-Some-Credit 150,000 10 2
walking-activity 149,332 4 22
social_bias_frames 144,649 16 3
Wikipedia_Talk_Labels 140,379 12 15
Municipal-Debt-Risk-Analysis 138,509 13 2
MiniBooNE 130,064 50 2
nba-shot-logs 128,069 15 2
college_scorecard 124,699 117 2
drug-directory 120,215 16 7
TVS_Loan_Default 119,528 29 2
road-safety 111,762 32 2
Diabetes130US 101,766 46 3
fars 100,968 29 8
Credit_Score_Classification 100,000 26 3
numerai28.6 96,320 21 2
Run_or_walk_information 88,588 6 2
jannis 83,733 54 4
KDD98 82,318 477 2
APSFailure 76,000 169 2
kick 72,983 32 2
human-choice-prediction 71,579 20 2
Traffic_violations 70,340 20 3
Fashion-MNIST 70,000 784 10
Cardiovascular-Disease-dataset 70,000 11 2
mnist_784 70,000 719 10
connect-4 67,557 42 3
mobile_churn 66,469 63 2
helena 65,196 27 100
LICD 63,634 413 2
CIFAR_10 60,000 3,072 10
REASONER 58,497 34 2
volkert 58,310 147 10
shuttle 58,000 9 7
GTSRB-HueHist 51,839 256 43
okcupid-stem 50,789 19 3
KDDCup09-Upselling 50,000 13,419 2
KDDCup09_appetency 50,000 207 2
adult 48,842 14 2
League-of-Legends-Diamond 48,651 14 2
tamilnadu-electricity 45,781 2 20
bank-marketing 45,211 16 2
meta_stream_intervals 45,164 74 11
jungle_chess 44,819 6 3
Dynamically-Generated-Hate-Speech-Dataset 41,144 8 2
Breast-cancer-prediction 39,998 11 2
Click_prediction_small 39,948 11 2
Hotel-Reviews 38,932 3 2
electricity 38,474 8 2
nomao 34,465 118 2
Employee-Turnover-at-TECHCO 34,452 9 2
Amazon_employee_access 32,769 9 2
Credit-Risk-Dataset 32,581 11 2
Default-of-Credit-Card-Clients-Dataset 30,000 23 2
funpedia 29,819 3 3
credit_risk_china 27,522 27 5
Insurance 23,548 10 2
guillermo 20,000 4,281 2
riccardo 20,000 4,283 2
insurance_dataset 20,000 26 4
letter 20,000 16 26
game-of-thrones-script-all-seasons 16,825 5 43
NewspaperChurn 15,855 16 2
mozilla4 15,545 5 2
pol 15,000 26 11
eeg-eye-state 14,980 14 2
MagicTelescope 13,376 10 2
nursery 12,958 8 4
online-shoppers-intention 12,330 17 2
Disaster-Tweets 11,370 4 2
mammography 11,183 6 2
PhishingWebsites 11,055 30 2
Binary-Dataset-of-Phishing-and-Legitimate-URLs 11,000 14 2
pendigits 10,992 16 10
WBCAtt 10,298 11 5
artificial-characters 10,218 7 10
internet_usage 10,108 71 46
robert 10,000 7,200 10
dilbert 10,000 2,000 5
shrutime 10,000 10 2
JapaneseVowels 9,961 14 9
GesturePhaseSegmentationProcessed 9,873 32 5
FICO-HELOC-cleaned 9,871 23 2
IBRD_Loans_Classification 9,215 6 10
Indian_pines 9,144 220 8
SpeedDating 8,378 120 2
fabert 8,237 795 7
mushroom 8,124 21 2
isolet 7,797 617 26
eye_movements 7,608 23 2
twonorm 7,400 20 2
blastchar 7,043 19 2
musk 6,598 167 2
first-order-theorem-proving 6,118 51 6
HMEQ_Data 5,960 12 2
philippine 5,832 308 2
optdigits 5,620 62 10
BachChoralHarmony 5,586 15 68
page-blocks 5,473 10 5
wall-robot-navigation 5,456 24 4
christine 5,418 1,611 2
phoneme 5,404 5 2
Is_fraud 5,227 19 2
sylvine 5,124 20 2
Satellite 5,100 36 2
Multiclass_Classification_for_Corporate_Credit 5,000 7 10
Personal-Loan-Modeling 5,000 12 2
churn 5,000 20 2
waveform-5000 5,000 40 3
air-quality-and-pollution-assessment 5,000 9 4
Heart_Failure_Prediction 5,000 12 2
compas-two-years 4,966 11 2
wine-quality-white 4,898 11 7
wilt 4,839 5 2
spambase 4,601 57 2
StackOverflow-polarity 4,423 1 3
hiva_agnostic 4,229 1,617 2
Fraud-Detection-Updated 4,156 27 2
ada 4,147 46 2
analcatdata_supreme 4,052 7 10
hypothyroid 3,770 27 3
Bioresponse 3,751 1,776 2
Internet-Advertisements 3,279 1,558 2
led24 3,200 24 10
kr-vs-kp 3,196 36 2
splice 3,190 60 3
dna 3,186 180 3
gina 3,153 970 2
madeline 3,140 259 2
jasmine 2,984 144 2
cjs 2,796 29 6
madelon 2,600 500 2
ozone-level-8hr 2,534 72 2
segment 2,310 16 7
cardiotocography 2,126 23 10
Estimation_of_Obesity_Levels 2,111 16 7
kc1 2,109 21 2
Corporate-Credit-Rating 2,026 30 8
mfeat-factors 2,000 216 10
South_Asian_Churn_dataset 2,000 13 2
mfeat-zernike 2,000 47 10
mfeat-fourier 2,000 76 10
pbcseq 1,945 18 3
steel-plates-fault 1,941 27 7
car 1,728 6 4
GAMETES_Heterogeneity 1,600 20 2
one-hundred-plants-texture 1,599 64 100
audit-data 1,552 35 2
OVA_Breast 1,545 10,935 2
amazon-commerce-reviews 1,500 10,000 50
yeast 1,484 8 10
cmc 1,473 9 3
ibm-employee-attrition 1,470 31 2
pc4 1,458 37 2
Data_Science_Nigeria_Telecoms_Churn 1,400 14 2
hepatitis_c_virus_hcv_for_egyptian_patients 1,385 28 4
Bank-Note-Authentication-UCI 1,372 4 2
baseball 1,340 16 3
Titanic 1,309 13 2
mental-health-in-tech-survey 1,259 26 2
hill-valley 1,212 100 2
Heart-Disease-Dataset-(Comprehensive) 1,190 11 2
volcanoes-e1 1,183 3 5
Airlines-Tweets-Sentiments 1,097 1 3
MiceProtein 1,080 77 8
cnae-9 1,080 856 9
solar_flare 1,058 9 5
qsar-biodeg 1,055 41 2
SOCC 1,043 13 4
rmftsa_sleepdata 1,024 2 4
autoUniv-au1-1000 1,000 20 2
collins 1,000 19 30
credit-g 1,000 20 2
vowel 990 12 11
The-Estonia-Disaster-Passenger-List 989 6 2
xd6 973 9 2
tokyo1 959 42 2
tic-tac-toe 958 9 2
Tour-and-Travels-Customer-Churn-Prediction 954 6 2
acp-breast-cancer 949 1 4
oil_spill 937 48 2
anneal 898 18 5
Cervical_Cancer_Risk_Factors 858 30 5
vehicle 846 18 4
analcatdata_authorship 841 70 4
glioma_grading_clinical_and_mutation_features 839 23 2
analcatdata_dmft 797 4 6
regensburg_pediatric_appendicitis 780 55 3
QSAR_Bioconcentration_classification 779 12 3
Diabetes_Dataset 768 8 2
blood-transfusion-service-center 748 4 2
eucalyptus 736 19 5
breast-w 699 9 2
Australian 690 14 2
soybean 683 35 19
profb 672 8 2
Student_Performance 666 11 4
balance-scale 625 4 3
Loan-Predication 614 11 2
monks-problems-2 601 6 2
synthetic_control 600 60 6
ilpd 583 10 2
micro-mass 571 1,082 20
wdbc 569 30 2
arsenic-male-lung 559 4 2
cylinder-bands 540 34 2
climate-model-simulation-crashes 540 18 2
water-treatment 527 36 2
Early-Stage-Diabetes-Risk-Prediction-Dataset 520 16 2
dresses-sales 500 12 2
irish 500 5 2
arrhythmia 443 262 10
wholesale-customers 440 7 2
vote 435 16 2
cars 406 7 3
chronic-kidney-disease 400 25 2
differentiated_thyroid_cancer_recurrence 383 16 2
colic 368 26 2
breast-cancer 286 9 2
qualitative-bankruptcy 250 6 2
us-2020-presidential-election-speeches 245 5 7
audiology 192 57 8
bone_marrow_transplant_children 187 36 2
darwin 174 450 2
tae 151 5 3
EgyptianSkulls 150 4 5
lymph 148 18 3
arcene 100 9,920 2
Table 9: The 253 Classification Datasets of the Pretraining Corpus.
Table 10: The 93 Regression Datasets of the Pretraining Corpus, with their $n$ examples, $m$ features, presence in a benchmark ($B$) and whether they are textual ($T$).
Dataset $n$ $m$ B T
delays_zurich_transport 5,465,575 14
New-York-Citi-Bike-Trip 4,500,000 7
USA-Airport-Dataset 3,606,803 14
New-York-Taxi-Trip 2,083,778 21
Buzzinsocialmedia_Twitter 583,250 77
nyc-taxi-green-dec-2016 581,835 18
515K-Hotel-Reviews-Data-in-Europe 515,738 16
dionis 416,188 54
Yolanda 400,000 100
Allstate_Claims_Severity 188,318 130
Football_players_Fifa_stats 183,142 37
black_friday 166,821 9
medical_charges 163,065 3
football-manager-data 159,541 87
wave_energy 72,000 32
video_transcoding 68,784 18
dating_profile 59,946 30
diamonds 53,940 9
sarcos 48,933 21
physiochemical_protein 45,730 9
fried 40,768 10
2dplanes 40,768 10
mv 40,768 10
Perth-House-Prices 33,656 17
cps88wages 28,155 6
fps_benchmark 24,624 39
news_popularity2 24,007 4
house_16H 22,784 16
health_insurance 22,272 11
house_sales 21,613 21
superconductivity 21,263 81
california_housing 20,640 8
avocado-sales 18,249 11
Bike_Sharing_Demand 17,379 12
elevators 16,599 18
FIFA20-Players 14,999 72
miami_housing 13,932 15
naval_propulsion_plant 11,934 14
Brazilian_houses 10,692 11
German-House-Prices 10,552 24
sulfur 10,081 5
climate_change_impact 10,000 14
grid_stability 10,000 12
Credit-Card-Dataset-for-Clustering 8,949 16
topo_2_1 8,885 261
yprop_4_1 8,885 212
seoul_bike_sharing_demand_cat 8,760 13
pumadyn32nh 8,192 32
kin8nm 8,192 8
cpu_activity 8,192 21
bank32nh 8,192 32
Pollen-Luxembourg-1992-2018 7,784 36
colleges 7,063 44
wind 6,574 14
QSAR-TID-10980 5,766 1,024
QSAR-TID-11 5,742 1,024
Myanmar-Air-Quality 5,122 10
Santander_transaction_value 4,459 4,735
SAT11-HAND-runtime-regression 4,440 114
Mercedes_Benz_Greener_Manufacturing 4,209 364
abalone 4,177 8
pollen 3,848 4
space_ga 3,107 6
scotch-whiskey-reviews-update-2020 2,247 4
quake 2,178 3
auction_verification 2,043 7
us_crime 1,994 126
airfoil_self_noise 1,503 5
house_prices 1,460 80
house_prices_nominal 1,460 79
NBA-PLAYERS–2016-2019 1,408 43
Insurance-Premium-Data 1,338 6
Moneyball 1,232 14
socmob 1,156 5
MIP-2016-regression 1,090 116
geographical_origin_of_music 1,059 116
concrete_compressive_strength 1,030 8
Household-monthly-electricity-bill 1,000 9
stock 950 9
QSAR_fish_toxicity 908 6
cars 804 17
energy_efficiency 768 8
kdd_el_nino-small 709 8
student_performance_por 649 30
strikes 625 6
sensory 576 11
meta 528 21
forest_fires 517 12
rmftsa_ladata 508 10
boston 506 13
no2 500 7
Diabetes(scikit-learn) 442 10
NBA-2k20-player-dataset 439 14
baseball-hitter 263 22
bodyfat 252 14
Lisbon-House-Prices 246 13
tecator 240 124
Table 10: The 93 Regression Datasets of the Pretraining Corpus.
Appendix D Benchmark Datasets
This appendix elaborates on the benchmark presented in §4. We consider all datasets proposed by AutoML Multimodal Benchmark (SHI) [60], Vectorizing (VEC) [31], and CARTE-Benchmark (CRT) [45], resulting in a final set of 50 datasets. We deduplicate datasets that appear as-is in more than one benchmark. In addition, since CARTE explores the concept of multi-table learning, they introduce highly-overlapping datasets for which we remove one variant (see 4.3 and B.2 in their paper).
Table 11 presents the classification datasets and Table 12 the regression ones. Each table includes an internal ID used for reference, the Dataset name, the number of examples $n$ and of features $m$, and the number of classes $C$ for classification. [^26] We treat ranking problems with up to 10 discrete values as multiclass problems. Finally, we also indicate the benchmark sources where each dataset appears. In addition, Table 13 presents the full benchmark with a short description per dataset, and Table 14 details the datasets removed during the deduplication process. Most of the excluded datasets are regression datasets from the CARTE-Benchmark, because of its high-overlapping nature.
Table 11: The 14 classification datasets of the benchmark, with their $n$ examples, $m$ features, $C$ classes, and presence in the SHI, VEC and CRT benchmarks.
ID Dataset $n$ $m$ $C$ SHI VEC CRT
C01 women_clothing_review 18,788 10 5
C02 us-accidents 7,728,394 42 4
C03 data_scientist_salary 15,841 6 6
C04 imdb_genre_prediction 800 11 2
C05 product_sentiment_machine_hack 5,091 2 4
C06 google_qa_question_type_reason 4,863 39 5
C07 michelin-guide-restaurants-2021 17,735 11 5
C08 fake_job_postings2 12,725 5 2
C09 jigsaw_unintended_bias100K 100,000 40 2
C10 yelp-reviews-dataset 10,000 5 5
C11 news_channel 20,284 17 6
C12 wine_reviews 84,123 5 30
C13 kick_starter_funding 86,502 9 2
C14 melbourne_airbnb 18,316 89 10
Table 12: The 36 regression datasets of the benchmark, with their $n$ examples, $m$ features, and presence in the SHI, VEC and CRT benchmarks.
ID Dataset $n$ $m$ SHI VEC CRT
R01 used-cars-dataset-cardekho 37,814 112
R02 second-hand-mercedes-benz 16,392 7
R03 animeplanet-recommendation 14,391 14
R04 ML/DS-Salaries 119,628 9
R05 Babies-R-Us 5,085 12
R06 employee_salaries 9,228 11
R07 spotify-tracks-dataset 114,000 18
R08 california_house_price 37,951 39
R09 fifa 19,178 28
R10 coffee-scrap-coffeereview 2,440 17
R11 BikeWale 9,003 6
R12 used-car-prices-in-pakistan 72,655 9
R13 bookprice_prediction 4,989 8
R14 ae_price_prediction 22,662 12
R15 Employee-remuneration 44,574 5
R16 filmtv-movies-dataset 41,399 17
R17 free-7-million-company-dataset 7,173,426 7
R18 museums 22,290 21
R19 vivino-wine-data 8,650 6
R20 wikiliq-dataset 12,569 12
R21 beer-profile-and-ratings 3,197 24
R22 korean-drama 1,647 9
R23 videogamesales 16,598 5
R24 zomato-bangalore-restaurants 41,665 15
R25 the-movies-dataset 45,460 20
R26 nba-draft-basketball 1,669 22
R27 Goodreads 3,967 14
R28 Rotten-Tomatoes 7,158 15
R29 saudi-arabia-used-cars-dataset 8,035 12
R30 top-ramen-ratings-2022 4,105 4
R31 Journal-Score-SJR 31,136 21
R32 chocolate-bar-ratings 1,795 8
R33 mercari_price_suggestion100K 100,000 9
R34 wine-price-on-polish-market 2,247 18
R35 clear-corpus 4,724 30
R36 jc_penney_products 10,860 5
Table 13: Benchmark Datasets Description
ID Description
C01 Women Clothing E-Commerce Reviews
C02 US Accidents between 2016 and 2023
C03 Indian Data Scientist Salary Prediction
C04 IMDB Movies Genre Prediction
C05 Product Sentiment Analysis
C06 Google QA Question Type Reason Explanation
C07 Michelin Guide Restaurants Awards
C08 Fake Job Posting Detection
C09 Online Social Media Comments Toxicity
C10 YELP Dataset Reviews
C11 News Channel Prediction
C12 Wine Reviews for Variety Prediction
C13 Kickstarter Funding Prediction
C14 Melbourne AirBnB Listings
R01 User cars and listing price in the website Cardekho
R02 Second-hand cars Mercedes Benz price Italy
R03 Anime-Planet Recommendation Database 2020
R04 Salaries of ML/DS Professionals Worldwide
R05 Prices Prediction for baby product from Babies R Us website
R06 Employee Salary in Montgomery County, MD
R07 Spotify Tracks Popularity
R08 California Houses 2020 Prices
R09 FIFA 2022 Players Wages
R10 Coffee Review Rating
R11 Bike and scooters from bikewale website in India
R12 Used car prices in Pakistan 2021
R13 Book Price Prediction
R14 American Eagle Retailer Price Prediction
R15 Employee Remuneration and Expenses - Vancouver
R16 FilmTV movies ataset rating
R17 Company size prediction
R18 General information on the US museums
R19 Vivino Spanish Wine Data
R20 WikiliQ - Alcohol dataset (May, 2022)
R21 Tasting profiles and consumer reviews for beers
R22 Korean Dramas
R23 Video Games Sales
R24 Zomato Restaurants in Bengaluru
R25 Metadata of movies released until 2017 for box-office revenues
R26 NBA Draft Basketball Player Data 1989-2021
R27 Books ratings
R28 Rotten Tomatoes Movie Ratings
R29 Saudi Arabia Used Cars Price from Syarah Website
R30 Ramen Ratings
R31 Academic impact for Scientific Journals
R32 Chocolate Bar expert ratings
R33 Mercari Online Marketplace Product Prices
R34 Information about wines on the polish market
R35 Readability scores for text passages spanning various genres and time periods
R36 JC Penney Product Prices in Retailer Website
Table 14: Excluded datasets, with their benchmark origin and reason for removal: (1) Duplicate dataset, (2) Unavailable, for datasets with inconvenient or unavailable hosting outside tabular repositories, and (3) Pretraining, for two (regression) datasets mistakenly used for the pretraining.
Dataset Benchmark Reason Duplicate
google_qa_answer_type SHI Duplicate google_qa_question_type
news_popularity2 SHI Pretraining
US Presidential VEC Unavailable
Journal Influence VEC Duplicate Journal-Score-SJR
Buy Buy Baby CRT Duplicate Babies-R-Us
Bikedekho CRT Duplicate BikeWale
Journal Score JCR CRT Duplicate Journal-Score-SJR
Japanese Anime CRT Duplicate animeplanet-recommendation
Mydramalist CRT Duplicate korean-drama
Prescription Drugs CRT Unavailable
Roger Ebert CRT Unavailable
US Presidential CRT Unavailable
Used Cars 24 CRT Duplicate used-car-prices-in-pakistan
Whisky CRT Pretraining
Wine.com CRT Duplicate wine_reviews
WineEnthusiasts CRT Duplicate wine_reviews
Appendix E Baselines
In this appendix we first discuss models excluded from the evaluation due to data leakage concerns, and then cover implementation details for the baselines used in our main experiments §4.
E.1 Excluded Baselines
As opposed to GDBTs or single-dataset deep learning methods, evaluating pretrained tabular models introduces additional complexity. Indeed, leakage can come in multiple forms. When LLMs are involved, there is a risk of memorization [8], and models trained on synthetic datasets [37] which try to mimic real-world distributions, can be unintentionally biased towards popular benchmarks.
While these two forms of leakage are subtle and hard to detect, a more direct form must be strictly avoided: When the same dataset (or a variant of it) is used during pretraining, and then it is evaluated as a downstream task. In such scenario there is inevitable severe data leakage, especially when running with multiple random test splits. The rest of the section explains how both TP-BERTa [80] and CM2 [82] suffer from such contamination with respect to our benchmark. As we briefly mention in §7, we advocate for improving TFM research by encouraging models that are practical to evaluate, by releasing several versions of each model, and providing default hyperparameters.
TP-BERTa
We exclude TP-BERTa from our evaluation for two key reasons. First, their implementation assumes that every example is treated as a serialized single sequence, which allows for a maximum length of 512 tokens, as elaborated in Appendix A.1. While this decision is efficient for datasets with a low amount of features and no free-text presence, around half of the datasets in our benchmark are too long for that limitation, as they either contain too many features or long free-texts. Second, TP-BERTa’s pretraining uses datasets that appear in our evaluation set, as listed in Table 6 of their paper [80]. It is evident that several datasets overlap directly with datasets in our benchmark §4 (e.g., 1510_fifa, 1368_IMDb-Ratings, 1639_Melbourne), disqualifying them for our purposes. Furthermore, we observe a concerning overlap between their pretraining and downstream task datasets (e.g., airlines, sf police and diabetes). We believe that this questions the validity of their evaluation, and that such contamination poses a serious challenge for the TFM community which could be substantially addressed by better tabular data repositories [73].
CM2
CM2 was pretrained over OpenTabs, a compilation of more than 2,000 datasets drawn from public tabular data repositories, including OpenML and Kaggle. While this collection is valuable, pretraining a model over these datasets compromises further evaluation of any of them. Naturally, the overlap with our benchmark here is extremely high, making it infeasible to use as a baseline. Interestingly, their repository [^27] https://github.com/Chao-Ye/CM2 lists TP-BERTa as a method trained on a subset of OpenTabs, reinforcing that the leakage is shared between the models.
E.2 Baselines Implementation and Hyperparameters
This section outlines the implementation and hyperparameter tuning strategy used for the baselines reported in §4. While each baseline has its own model-specific preprocessing pipeline, we apply two shared preprocessing steps to both TabSTAR (as detailed in Appendix A.1) and all the baselines: (1) We perform date preprocessing by using skrub’s DatetimeEncoder, and (2) Apply a clipped z-score transformation for target variables in regression datasets.
Textual Feature Handling
CARTE natively supports textual inputs, and the TabPFN-v2 API client [^28] https://github.com/PriorLabs/tabpfn-client does as well, although its implementation details remain undisclosed. On the other hand, GBDTs do not natively support free-text features, [^29] CatBoost includes a built-in text module, but it underperforms compared to dense text embeddings., so we preprocess these features into fixed-size embeddings using skrub’s TextEncoder, which internally applies a frozen e5-small-v2 encoder to each semantic column. This aligns with the encoder used in TabSTAR, enabling a fair comparison across models. There are, however, two key differences in how TabSTAR handles these embeddings: First, the embeddings are specifically finetuned for the task, contributing significantly to its strong performance as shown in §5 and further analyzed in §G.2. The second detail is that skrub applies dimensionality reduction to 30 dimensions, as proposed by [31]. This compressed representation performs comparably to the full embedding space, while offering improved inference efficiency.
E.2.1 TabPFN-v2
We run TabPFN-v2 using their API client which supports text features. [^30] We use v2.0.8, the latest version available at the time of running the experiments. While the intrinsic details of their textual handling remain undocumented, it’s reasonable to assume that it resembles the processing we apply to GBDTs, as their model leverages ICL and their architecture has no textual encoder.
E.2.2 CARTE
[…]
E.2.3 CatBoost and CatBoost-Tuned
We run CatBoost using the catboost package [^33] https://pypi.org/project/catboost/ and run the default configuration suggested by [25] by setting $early\_stopping\_rounds=50$, $od\_pval=0.001$, $iterations=2000$. For the tuned version, we use the Optuna package [^34] https://pypi.org/project/optuna/ with random search, with a budget of 4 hours for every run and parallelizing trials on 8 CPU cores. We use 5-fold cross-validation and take the best configuration selected based on this mean score. We use it then to retrain the model on the full training data. For hyperparameter space, we follow the hyperparameter search suggested by [37] as detailed in Table 15.
Table 15: CatBoost-Tuned Hyperparameters Search Space
Hyperparameter Search Space
lr $\log\mathcal{U}(e^{-5},1)$
random_strength $\mathcal{U}\{1,2,\ldots,20\}$
l2_leaf_reg $\log\mathcal{U}(1,10)$
bagging_temperature $\mathcal{U}(0.0,1.0)$
leaf_estimation_iterations $\mathcal{U}\{1,2,\ldots,20\}$
iterations $\mathcal{U}\{100,101,\ldots,4000\}$
E.2.4 XGBoost and XGBoost-Tuned
We run XGBoost using the xgboost package [^35] https://pypi.org/project/xgboost/ and follow the same procedure as for CatBoost, except additional preprocessing (e.g., transforming categorical variables into one-hot encodings). For the default configuration, we follow the suggestion of [25] and use: $booster="gbtree"$, $early\_stopping\_rounds=50$, $n\_estimators=2000$. For the tuned variant, we follow the hyperparameter search space suggested by [37], as shown in Table 16.
Table 16: XGBoost-Tuned Hyperparameters Search Space
Hyperparameter Search Space
learning_rate $\log\mathcal{U}(e^{-7},1)$
max_depth $\mathcal{U}\{1,2,\ldots,10\}$
subsample $\mathcal{U}(0.2,1)$
colsample_bytree $\mathcal{U}(0.2,1)$
colsample_bylevel $\mathcal{U}(0.2,1)$
min_child_weight $\log\mathcal{U}(e^{-16},e^{5})$
alpha $\log\mathcal{U}(e^{-16},e^{2})$
reg_lambda $\log\mathcal{U}(e^{-16},e^{2})$
gamma $\log\mathcal{U}(e^{-16},e^{2})$
n_estimators $\mathcal{U}\{100,101,\ldots,4000\}$
E.2.5 Random Forest
We treat Random Forest as a weak baseline to establish a lower-bound reference for each dataset split. We run it with the sklearn package  [^36] https://scikit-learn.org/ and use its default configuration with $n\_estimators=100$.
Appendix F Extended Main Results
In this appendix we provide the main results for the experiment as reported in §5. As elaborated in §4, each model is evaluated on each dataset across 10 splits. Since performance scales vary between datasets, we follow the normalization approach proposed by [37], rescaling all scores to the $[0,1]$ range. The final reported performance for each model is the average over these normalized runs, and we compute 95% confidence intervals using the standard normal approximation: $\hat{\mu}\pm 1.96\,\frac{\hat{\sigma}}{\sqrt{n}}$.
In this section, we often use abbreviated model names for conciseness: We refer to CatBoost variants as CatB and CatB-T, with the latter being the Tuned version. Similarly, we use XGB and XGB-T for XGBoost. We maintain the abbreviation of Random Forest (RF) and shorten TabPFN-v2 to simply TabPFN. For models that run on an unlimited number of training examples, we add a -U suffix, i.e. TabSTAR-U, CatB-T-U, and XGB-T-U.
F.1 Dataset Level Performance
We report AUROC for classification and $R^{2}$ for regression, with 95% CIs computed over the 10 runs for each dataset. Tables  17 and 18 summarize classification performance on datasets with up to 10K and over 10K examples, respectively. Tables 19 and 20 to the same for regression tasks. For conciseness, datasets are referred by their ID from Appendix D.
As discussed in §5, TabPFN-v2 is unable to run over 4 datasets: C12, because it is a multiclass problem with more than 10 classes, and C02, C14 and R01 because they support inference for up to 500,000 cells. Attempts to run the model over a subset of the examples led to a significantly worse performance, and thus we decide not to report them to allow a fair comparison. Furthermore, CARTE is unable to run over 15 of the datasets in the benchmark due a known bug [^37] https://github.com/soda-inria/carte/issues/23 in their implementation for the PowerTransformation, which struggles in the presence of features with too less unique values.
Table 17: Classification performance per dataset (up to 10K). The top performance score is bolded first, and then all scores are rounded. We report average AUROC with 95% CIs.
ID CARTE CatB CatB-T RF TabPFN TabSTAR XGB XGB-T
C01 88.4±0.3 90.2±0.3 90.3±0.4 88.8±0.3 90.3±0.3 90.8±0.3 89.3±0.4 90.2±0.4
C02 97.3±0.5 97.4±0.4 96.3±0.5 97.9±0.5 97.2±0.3 97.6±0.3
C03 82.5±0.3 82.0±0.3 81.2±0.6 77.3±0.3 82.4±0.3 83.0±0.3 80.5±0.3 82.1±0.3
C04 84.5±1.9 85.4±1.8 82.4±3.0 88.3±1.5 83.7±2.3 80.7±2.9 85.4±2.0
C05 88.5±0.9 90.6±1.0 90.9±0.5 88.0±1.1 91.2±0.7 91.3±0.8 88.1±1.1 90.3±0.6
C06 80.9±1.5 81.3±1.3 82.6±1.1 73.6±1.2 87.7±0.5 87.0±0.6 81.5±1.0 83.7±0.9
C07 90.1±0.4 90.3±0.3 90.6±0.3 85.1±0.7 89.8±0.4 91.5±0.3 87.2±0.6 89.3±0.5
C08 90.8±1.6 93.0±0.9 93.2±1.0 90.2±1.3 91.3±1.1 93.5±1.5 91.9±1.1 94.4±0.7
C09 99.8±0.1 99.8±0.1 99.6±0.2 99.8±0.1 99.3±0.1 99.7±0.1 99.8±0.1
C10 86.7±0.2 87.0±0.2 84.6±0.2 87.6±0.4 89.1±0.2 85.4±0.3 86.8±0.2
C11 78.4±0.5 79.7±0.5 80.7±0.4 76.2±0.6 81.4±0.4 79.2±0.5 78.2±0.6 80.2±0.5
C12 96.0±0.2 97.5±0.1 97.7±0.1 95.3±0.2 98.3±0.1 96.6±0.2 97.3±0.1
C13 70.9±0.8 73.7±1.1 74.1±1.0 71.1±1.1 72.3±0.9 75.0±0.7 70.2±0.8 74.1±1.1
C14 83.5±0.3 84.0±0.5 80.1±0.3 84.0±0.3 81.4±0.4 83.2±0.4
Table 18: Classification performance per dataset (above 10K). The top performance score is bolded first, and then all scores are rounded. We report average AUROC with 95% CIs.
ID CatB-T CatB-T-U TabPFN TabSTAR TabSTAR-U XGB-T XGB-T-U
C01 90.3±0.4 90.5±0.4 90.3±0.3 90.8±0.3 91.2±0.3 90.2±0.4 90.4±0.3
C02 97.4±0.4 98.3±0.3 97.9±0.5 98.4±0.2 97.6±0.3 98.2±0.3
C03 81.2±0.6 81.5±0.8 82.4±0.3 83.0±0.3 83.8±0.5 82.1±0.3 82.3±0.3
C07 90.6±0.3 91.0±0.4 89.8±0.4 91.5±0.3 91.9±0.3 89.3±0.5 89.9±0.5
C08 93.2±1.0 93.1±1.2 91.3±1.1 93.5±1.5 95.1±0.9 94.4±0.7 94.4±0.9
C09 99.8±0.1 99.8±0.1 99.8±0.1 99.3±0.1 99.5±0.1 99.8±0.1 99.8±0.1
C11 80.7±0.4 81.8±0.4 81.4±0.4 79.2±0.5 81.0±0.5 80.2±0.5 81.4±0.4
C12 97.7±0.1 98.5±0.1 98.3±0.1 99.1±0.1 97.3±0.1 98.4±0.1
C13 74.1±1.0 76.7±0.8 72.3±0.9 75.0±0.7 77.9±1.0 74.1±1.1 76.9±0.9
C14 84.0±0.5 84.8±0.4 84.0±0.3 85.1±0.3 83.2±0.4 84.2±0.4
Table 19: Regression performance per dataset. The top performance score is bolded first, and then all scores are rounded. We report average $R^{2}$ with 95% CIs.
ID CARTE CatB CatB-T RF TabPFN TabSTAR XGB XGB-T
R01 100±0.0 100±0.0 100±0.0 100±0.0 100±0.0 99.9±0.1 100±0.0
R02 98.3±1.0 98.3±0.9 98.0±1.0 98.4±0.9 98.0±1.1 97.9±1.1 98.3±0.9
R03 71.8±0.5 73.8±0.4 74.1±0.3 68.4±0.6 74.9±0.4 70.9±0.8 69.1±0.5 73.7±0.5
R04 86.2±7.4 85.1±7.1 85.1±6.8 86.1±5.2 81.5±8.7 85.8±7.5 87.5±3.7
R05 93.2±0.8 89.5±1.3 90.0±1.2 86.7±1.3 92.7±0.9 93.2±1.2 87.1±1.4 90.1±1.3
R06 97.7±0.8 98.1±0.5 98.2±0.5 97.2±1.0 98.5±0.6 98.1±0.3 97.1±0.7 97.9±0.7
R07 71.6±1.3 66.9±0.8 67.8±0.8 61.5±1.2 62.2±1.3 71.1±1.8 61.9±1.5 68.8±0.8
R08 93.2±0.8 93.0±0.7 93.2±0.7 92.2±0.7 93.9±0.7 92.8±0.6 92.3±0.7 93.1±0.7
R09 89.2±0.5 89.2±0.6 89.3±0.4 88.6±0.7 89.8±0.5 88.8±0.5 88.2±0.7 89.3±0.5
R10 99.5±0.2 99.0±0.6 98.9±0.6 98.1±0.3 99.1±0.3 99.5±0.1 98.6±0.3 99.4±0.2
R11 94.2±0.9 94.0±0.9 92.5±1.3 94.5±0.9 93.9±1.1 93.3±1.1 94.2±0.8
R12 98.5±0.3 98.5±0.2 98.0±0.3 98.5±0.2 98.3±0.2 98.2±0.3 98.5±0.2
R13 52.8±2.1 57.6±2.7 58.2±2.6 51.8±3.5 53.8±2.7 53.4±3.0 49.0±3.3 57.5±2.7
R14 96.6±0.2 97.3±0.1 97.3±0.1 96.8±0.1 96.2±0.2 96.4±0.1 97.1±0.1 97.1±0.1
R15 79.5±0.9 79.9±1.1 77.3±1.2 79.4±1.0 78.9±1.5 76.8±1.2 80.3±1.0
R16 98.7±0.0 98.7±0.0 98.7±0.0 97.7±0.1 98.8±0.0 98.7±0.0 97.7±0.1 97.8±0.1
R17 95.3±2.0 94.8±2.4 95.2±2.0 94.2±2.5 95.1±2.1 94.7±2.3 94.5±2.6 95.4±1.9
R18 98.1±0.4 98.2±0.4 98.2±0.4 97.7±0.4 93.1±1.3 97.4±0.6 97.5±0.8 98.3±0.4
R19 85.5±0.8 85.8±0.9 85.3±0.9 84.0±1.1 83.8±1.1 83.9±1.0 82.4±2.0
R20 96.4±0.5 96.0±0.6 96.2±0.5 95.7±0.6 93.9±1.0 95.6±0.9 95.3±0.9 96.2±0.5
R21 92.3±1.2 92.5±1.1 92.6±1.1 91.8±1.1 93.3±1.1 92.3±1.0 91.6±1.0 92.6±1.0
R22 45.2±5.5 45.2±5.9 39.1±6.4 43.8±5.5 39.7±5.5 36.1±7.1 46.8±5.8
R23 85.0±1.8 85.4±1.2 85.5±1.1 81.8±1.5 84.8±1.6 83.0±1.6 83.3±1.2 85.1±1.2
R24 86.0±0.6 84.0±0.8 85.8±0.7 79.7±1.0 70.3±1.6 81.8±1.8 82.9±0.7 85.9±0.9
R25 94.3±0.6 95.4±0.5 95.3±0.5 94.8±0.5 86.7±1.0 94.8±0.5 94.6±0.6 95.4±0.5
R26 99.8±0.1 99.4±0.1 99.4±0.1 99.0±0.2 99.8±0.1 99.6±0.1 99.1±0.2 99.5±0.1
R27 81.9±1.6 85.2±1.3 85.3±1.4 84.8±1.3 82.2±1.6 82.0±1.6 83.2±1.2 85.5±1.2
R28 52.5±2.6 53.4±2.7 53.3±2.8 46.0±2.7 61.7±2.1 51.5±2.9 45.3±2.3 53.5±2.7
R29 94.4±0.7 94.4±0.7 93.0±1.0 95.7±0.6 94.3±0.9 93.4±0.8 94.5±0.8
R30 23.8±4.2 22.7±4.1 24.6±3.9 23.5±3.5 20.9±4.8 15.7±5.0 17.5±4.2 25.5±3.7
R31 92.1±0.3 92.1±0.4 92.0±0.4 89.9±0.5 93.2±0.3 91.7±0.4 90.3±0.5 92.0±0.4
R32 28.8±6.3 28.2±5.8 28.6±5.9 26.2±5.4 19.4±5.6 22.7±7.3 31.6±6.2
R33 46.6±1.5 47.4±1.7 47.8±1.7 40.2±1.6 44.9±1.6 46.0±1.8 40.3±2.1 47.9±1.6
R34 90.9±2.5 91.4±2.1 88.7±3.1 92.3±1.8 89.1±2.9 88.9±3.6 91.6±2.1
R35 85.9±0.6 84.4±0.5 84.5±0.5 81.2±0.4 85.2±0.7 85.7±0.8 81.6±0.8 84.4±0.3
R36 96.3±0.9 95.2±0.9 95.5±0.9 94.3±1.0 91.2±1.1 95.7±0.8 94.4±0.9 95.5±0.9
Table 20: Regression performance per dataset (above 10K). The top performance score is bolded first, and then all scores are rounded. We report average $R^{2}$ with 95% CIs.
ID CatB-T CatB-T-U TabPFN TabSTAR TabSTAR-U XGB-T XGB-T-U
R01 100±0.0 100±0.0 100±0.0 100±0.0 100±0.0 100±0.0
R02 98.3±0.9 98.8±0.9 98.4±0.9 98.0±1.1 98.4±1.0 98.3±0.9 98.8±0.9
R03 74.1±0.3 75.1±0.4 74.9±0.4 70.9±0.8 72.3±0.6 73.7±0.5 74.6±0.6
R04 85.1±7.1 91.3±0.8 86.1±5.2 81.5±8.7 91.0±0.7 87.5±3.7 91.3±0.6
R07 67.8±0.8 85.2±0.3 62.2±1.3 71.1±1.8 79.9±1.8 68.8±0.8 86.1±0.7
R08 93.2±0.7 93.8±0.6 93.9±0.7 92.8±0.6 93.3±0.6 93.1±0.7 93.9±0.5
R09 89.3±0.4 89.5±0.4 89.8±0.5 88.8±0.5 89.2±0.6 89.3±0.5 89.5±0.4
R12 98.5±0.2 98.9±0.1 98.5±0.2 98.3±0.2 98.4±0.2 98.5±0.2 98.9±0.2
R14 97.3±0.1 97.8±0.1 96.2±0.2 96.4±0.1 96.8±0.2 97.1±0.1 97.5±0.1
R15 79.9±1.1 86.5±0.7 79.4±1.0 78.9±1.5 83.9±1.4 80.3±1.0 87.0±0.7
R16 98.7±0.0 98.8±0.0 98.8±0.0 98.7±0.0 98.8±0.0 97.8±0.1 98.0±0.1
R17 95.2±2.0 97.6±0.6 95.1±2.1 94.7±2.3 97.6±0.6 95.4±1.9 97.7±0.6
R18 98.2±0.4 99.0±0.2 93.1±1.3 97.4±0.6 97.7±0.6 98.3±0.4 98.9±0.3
R20 96.2±0.5 96.3±0.4 93.9±1.0 95.6±0.9 95.3±1.0 96.2±0.5 96.3±0.5
R23 85.5±1.1 86.3±0.7 84.8±1.6 83.0±1.6 85.2±1.7 85.1±1.2 86.1±0.8
R24 85.8±0.7 97.1±0.3 70.3±1.6 81.8±1.8 95.5±0.7 85.9±0.9 97.3±0.3
R25 95.3±0.5 95.9±0.4 86.7±1.0 94.8±0.5 95.1±0.5 95.4±0.5 95.8±0.5
R31 92.0±0.4 93.0±0.3 93.2±0.3 91.7±0.4 92.7±0.4 92.0±0.4 92.9±0.3
R33 47.8±1.7 55.9±1.1 44.9±1.6 46.0±1.8 57.2±1.3 47.9±1.6 55.7±1.3
R36 95.5±0.9 95.5±0.9 91.2±1.1 95.7±0.8 95.5±0.5 95.5±0.9 95.5±0.9
F.2 Head-to-head comparisons
We compare the performance of TabSTAR against each of the models in head-to-head comparisons. We report win rate, which can be seen as a private case of the normalized metric with only two models. We exclude failed runs when comparing to CARTE and TabPFN-v2. Table 21 shows the performance of TabSTAR against all models competing up to 10K examples, for both regression and classification, with 95% CIs over the win rate. Table 22 does the same for TabSTAR-Unlimit.
Table 21: Win rates of TabSTAR (up to 10K) against baselines. Win rate with 95% CI.
CARTE CatB CatB-T RF TabPFN XGB XGB-T
Classification 93.3±5.2 73.6±7.3 67.9±7.8 88.6±5.3 57.3±9.3 89.3±5.1 71.4±7.5
Regression 35.7±5.9 33.6±4.9 32.2±4.8 66.9±4.9 40.9±5.2 69.4±4.8 30.3±4.8
Table 22: Win rates of TabSTAR-U (above 10K) against baselines. Win rate with 95% CI.
CARTE CatB-T CatB-T-U TabPFN TabSTAR XGB-T XGB-T-U
Classification 98.6±2.8 86.0±6.8 67.0±9.3 75.7±10.1 94.0±4.7 85.0±7.0 71.0±8.9
Regression 63.5±7.5 53.0±6.9 22.0±5.8 61.1±7.0 82.0±5.3 59.5±6.8 28.5±6.3
F.3 Running Times and Compute Information
In this section we present the running times of the different models (see Table 23). We focus on the condition with up to 10,000 examples, excluding CatBoost-Tuned and XGBoost-Tuned as they run with a 4 hours budget using 8 CPUs from AMD EPYC 7742 64-Core Processor .
Compute Information
TabSTAR runs on NVIDIA RTX A40 with 48GB memory. TabPFN-v2 runs directly on their API client. CatBoost, XGboost and Random Forest run using Intel(R) Core(TM) i9-10920X CPU @ 3.50GHz CPU cores. CARTE use inferior GPUs of type NVIDIA GeForce GTX 1080 Ti with 11GB memory, to allow for higher parallelization of their incredibly slow runs.
Running Times
TabSTAR’s average running time for a downstream task ranges from 30 seconds (C04) to 7,692 seconds (R24). We observe that these running times are significantly lower than 14,400 seconds (4 hours), used for CatBoost-Tuned and XGBoost-Tuned, although a single run of any tree model without 5-fold cross-validation is faster. The running times of TabPFN-v2 are considerably better. However, since they use ICL, they are limited to run only up to 10,000 examples and their inference time is just as costly as their fitting time. For CARTE, while their runs use inferior hardware and limited parallelization, the reported times are for a single run. In practice, they run 6 times per dataset split because of their lack of default hyperparameters, making it a poor choice for scale.
Table 23: Average model training running time in seconds, per dataset, for up to 10K examples.
ID CARTE CatBoost RF TabPFN-v2 TabSTAR XGBoost
C01 13,341 44 19 42 302 31
C02 420 346 1,145 491
C03 7,297 38 19 61 324 27
C04 11 6 9 30 9
C05 1,227 10 8 14 68 9
C06 3,068 68 31 112 1,126 63
C07 6,559 61 34 83 685 52
C08 4,631 40 41 49 681 49
C09 20 21 134 544 31
C10 22 22 30 519 23
C11 2,428 19 12 33 252 18
C12 4,411 169 16 765 39
C13 1,774 31 27 58 929 37
C14 500 140 5,254 308
R01 9,439 139 446 1,987 94
R02 9 16 54 188 6
R03 2,174 30 103 357 553 24
R04 10 10 84 115 7
R05 10,639 28 26 62 158 15
R06 3,185 22 43 70 307 13
R07 4,637 22 109 112 352 28
R08 4,252 57 366 396 1,136 58
R09 8,239 7 22 55 198 5
R10 1,505 28 37 33 168 21
R11 11 18 36 99 8
R12 13 22 78 199 9
R13 2,299 24 54 38 639 21
R14 6,396 28 30 167 772 15
R15 25 51 66 223 22
R16 13,380 34 118 156 798 35
R17 3,247 94 478 204 270 143
R18 9,663 34 631 309 341 32
R19 11 14 66 101 6
R20 6,173 22 144 154 637 20
R21 2,215 14 33 25 177 14
R22 16 21 17 149 15
R23 14,401 15 51 58 167 10
R24 25,903 90 170 314 7,692 106
R25 2,404 50 943 261 786 54
R26 636 10 11 13 71 8
R27 2,615 31 59 40 304 22
R28 12,265 45 163 128 389 46
R29 11 10 32 138 5
R30 2,477 9 18 24 45 7
R31 5,102 41 187 240 451 39
R32 8 11 13 32 7
R33 4,512 28 100 136 539 30
R34 26 24 18 61 12
R35 4,816 30 55 48 396 28
R36 15,389 18 81 79 949 23
Appendix G Extended Analysis
In this section we expand on the analysis results discussed in §6.
G.1 Evaluation Datasets for Analysis
All the experiments are evaluated over 20 datasets from the benchmark in Appendix D. Each experiment reports the performance with AUROC for classification and $R^{2}$ for regression. For conciseness, each tables reports both regression and classification tasks, distinguishable by their ID.
G.2 The Role of the Encoder Unfreezing (Q1)
Table 24 shows the results for each variant of the experiment presented in §6. It is evident that unfreezing the textual encoder yields a significant performance across datasets. Furthermore, while finetuning a single layer gives a significant boost, it underperforms compared to 6 unfrozen layers.
G.3 The Effect of Pretraining (Q2)
We pretrain three TabSTAR variants on nested dataset subsets of size 16, 64, and 256. The 64-dataset variant contains the original 16 plus 48 new datasets, and the 256-dataset variant builds on those 64 by adding another 192. This cumulative design minimizes variance between variants so that performance differences reflect only the effect of increasing data volume.
While LoRA [41] is a very efficient technique, its performance was much worse for a randomly initialized model. Therefore, we perform full finetuning of the non-pretrained model, as explained in Appendix B.2. Table 25 shows the dataset level results. It is evident that for most of the datasets, improvement is observed when scaling, with the 256 datasets variant winning for almost all datasets.
G.4 Numerical Verbalization (Q3)
We show the full results for the experiment in §6, with Table 26 illustrating the verbalizations in each variant. Note that we do not include an exact value verbalization, since it would increase the number of unique text inputs and place extra memory demands. The two variants which integrate numerical information into the verbalization dominate the experiment, although the improvement seems to be marginal for some datasets. Interestingly, some datasets significantly underperform, with the R27 dataset completely failing the task. The addition of the quantile information on top of the bin seems to have limited impact, although marginally winning on the average performance.
Table 24: Downstream performance for Q1: The Role of the Encoder Unfreezing. Results for 20 datasets with 95% CI, for varying number of unfrozen layers. The top performance score is bolded first, and then all scores are rounded. We report AUROC for classification and $R^{2}$ for regression.
ID 0 1 3 6 9
C01 87.8±0.3 90.4±0.4 90.8±0.4 91.0±0.4 90.8±0.4
C02 94.9±1.2 97.8±0.3 97.8±0.3 98.1±0.3 97.7±0.3
C03 77.6±0.9 82.4±0.2 82.8±0.3 83.0±0.5 83.0±0.4
C05 87.0±0.7 90.2±0.9 90.8±0.6 91.5±0.9 89.4±0.8
C07 82.1±1.4 89.1±0.9 90.6±0.5 91.3±0.4 91.0±0.4
C11 77.4±0.5 78.4±0.6 78.7±0.7 78.8±0.5 78.3±0.6
C12 94.4±0.2 98.3±0.1 98.3±0.1 98.3±0.1 98.2±0.1
C13 66.5±0.9 71.7±1.1 72.9±0.7 74.1±0.7 73.0±0.8
R02 97.0±1.7 97.9±1.2 98.0±1.1 97.9±1.2 98.0±1.1
R03 67.2±0.8 69.8±0.8 70.9±0.9 71.2±0.8 70.8±0.6
R05 80.8±2.6 91.8±1.5 93.2±0.5 93.1±0.7 92.9±0.8
R09 88.1±0.7 88.7±0.7 89.0±0.5 89.0±0.6 88.9±0.8
R12 97.0±0.6 98.2±0.3 98.2±0.2 98.3±0.3 98.2±0.3
R13 37.6±3.6 45.7±2.1 51.6±3.2 51.4±2.9 52.1±1.8
R18 96.3±1.1 97.1±0.6 97.4±0.6 97.3±0.6 97.0±0.6
R23 79.3±2.4 81.6±1.6 82.8±1.6 83.2±2.2 82.1±1.9
R27 81.3±1.6 81.3±1.6 81.5±1.6 81.3±1.7 81.4±1.6
R30 14.2±4.6 15.6±5.6 19.2±3.1 19.1±4.0 17.3±4.5
R33 36.0±2.0 43.6±1.5 44.3±1.4 46.0±1.8 43.5±2.2
R34 84.3±2.8 87.4±2.9 88.4±3.4 87.9±3.5 88.7±2.7
Table 25: Downstream performance for Q2: The Effect of Pretraining. Results for 20 datasets with 95% CI, for varying number of pretraining datasets. The top performance score is bolded first, and then all scores are rounded. We report AUROC for classification and $R^{2}$ for regression.
ID 0 16 64 256
C01 90.8±0.4 90.7±0.4 90.7±0.3 91.0±0.4
C02 98.0±0.4 97.4±0.8 97.8±0.3 98.1±0.3
C03 69.2±8.9 83.1±0.4 83.2±0.3 83.0±0.5
C05 90.7±0.8 90.3±1.1 90.6±0.5 91.5±0.9
C07 87.9±0.5 87.6±0.8 90.9±0.3 91.3±0.4
C11 78.2±0.6 77.4±0.9 78.0±0.4 78.8±0.5
C12 98.3±0.1 98.2±0.1 98.2±0.1 98.3±0.1
C13 74.0±0.5 73.5±0.8 73.7±1.0 74.1±0.7
R02 97.2±1.6 97.6±1.3 97.9±1.2 97.9±1.2
R03 67.3±1.9 68.5±1.0 71.2±0.7 71.2±0.8
R05 88.0±2.8 90.6±2.2 92.2±1.0 93.1±0.7
R09 88.1±1.0 88.4±0.7 88.8±0.5 89.0±0.6
R12 97.7±0.3 97.9±0.3 98.1±0.2 98.3±0.3
R13 49.4±2.7 49.1±3.6 48.0±3.0 51.4±2.9
R18 95.0±2.2 94.9±1.2 96.7±0.6 97.3±0.6
R23 82.2±1.6 81.2±2.3 81.8±2.3 83.2±2.2
R27 80.9±1.7 81.3±1.6 81.5±1.6 81.3±1.7
R30 13.1±4.1 18.5±4.9 18.5±3.9 19.1±4.0
R33 45.4±2.2 42.8±3.3 45.0±1.5 46.0±1.8
R34 83.8±4.3 86.0±3.3 88.0±3.4 87.9±3.5
Table 26: Illustrative verbalization of a numerical feature (Age) for the Q3: Numerical Verbalization experiment.
Value Name Name + Bin TabSTAR
17 Age: Numeric Age: Lower than 18 Age: Lower than 18 (Quantile 0%)
20 Age: Numeric Age: 18–23 Age: 18–23 (Quantile 0–10%)
25 Age: Numeric Age: 23–27 Age: 23–27 (Quantile 10–20%)
29 Age: Numeric Age: 27–31 Age: 27–31 (Quantile 20–30%)
33 Age: Numeric Age: 31–35 Age: 31–35 (Quantile 30–40%)
38 Age: Numeric Age: 35–40 Age: 35–40 (Quantile 40–50%)
42 Age: Numeric Age: 40–45 Age: 40–45 (Quantile 50–60%)
48 Age: Numeric Age: 45–51 Age: 45–51 (Quantile 60–70%)
55 Age: Numeric Age: 51–58 Age: 51–58 (Quantile 70–80%)
63 Age: Numeric Age: 58–67 Age: 58–67 (Quantile 80–90%)
83 Age: Numeric Age: 67-87 Age: 67–87 (Quantile 90–100%)
93 Age: Numeric Age: Higher than 87 Age: Higher than 87 (Quantile 100%)
Age: Unknown Value Age: Unknown Value Age: Unknown Value
Table 27: Downstream performance for Q3: Numerical Verbalization. Results for 20 datasets with 95% CI, for different verbalizations. The top performance score is bolded first, and then all scores are rounded. We report AUROC for classification and $R^{2}$ for regression.
ID Name Name + Bin TabSTAR
C01 91.0±0.4 91.2±0.4 91.0±0.4
C02 98.1±0.2 97.9±0.3 98.1±0.3
C03 83.3±0.4 83.3±0.3 83.0±0.5
C05 90.8±0.9 91.1±1.2 91.5±0.9
C07 91.2±0.2 91.4±0.4 91.3±0.4
C11 78.2±0.5 78.2±0.6 78.8±0.5
C12 98.2±0.1 98.4±0.1 98.3±0.1
C13 71.6±0.7 73.8±0.9 74.1±0.7
R02 98.0±1.1 98.0±1.1 97.9±1.2
R03 67.4±0.7 71.3±0.6 71.2±0.8
R05 93.1±1.1 93.2±0.8 93.1±0.7
R09 88.9±0.5 88.9±0.6 89.0±0.6
R12 98.2±0.2 98.3±0.2 98.3±0.3
R13 50.1±3.3 49.7±3.4 51.4±2.9
R18 97.1±0.6 97.1±0.7 97.3±0.6
R23 81.9±2.4 82.7±1.9 83.2±2.2
R27 16.7±6.6 82.0±1.5 81.3±1.7
R30 16.4±4.4 17.7±5.3 19.1±4.0
R33 45.6±1.1 46.2±2.2 46.0±1.8
R34 88.0±3.6 88.6±2.5 87.9±3.5

### v2

@@
-TabSTAR: A Foundation Tabular Model With Semantically Target-Aware Representations
+TabSTAR: A Tabular Foundation Model for Tabular Data with Text Fields
 Abstract
-While deep learning has achieved remarkable success across many domains, it has historically underperformed on tabular learning tasks, which remain dominated by gradient boosting decision trees (GBDTs).
+While deep learning has achieved remarkable success across many domains, it has historically underperformed on tabular learning tasks, which remain dominated by gradient boosting decision trees.
 However, recent advancements are paving the way for Tabular Foundation Models, which can leverage real-world knowledge and generalize across diverse datasets, particularly when the data contains free-text.
 Although incorporating language model capabilities into tabular tasks has been explored, most existing methods utilize static, target-agnostic textual representations, limiting their effectiveness.
-We introduce TabSTAR: a Foundation Tabular Model with Semantically Target-Aware Representations.
+We introduce TabSTAR: a Tabular Foundation Model with Semantically Target-Aware Representations.
 TabSTAR is designed to enable transfer learning on tabular data with textual features, with an architecture free of dataset-specific parameters.
@@
 TabSTAR achieves state-of-the-art performance for both medium- and large-sized datasets across known benchmarks of classification tasks with text features, and its pretraining phase exhibits scaling laws in the number of datasets, offering a pathway for further performance improvements.
+[^1] Code is available at https://github.com/alanarazi7/TabSTAR.
 1 Introduction
@@
 Addressing these limitations is crucial for developing tabular models that leverage semantic information, transfer knowledge from many datasets, and generalize across domains.
-In this paper, we introduce TabSTAR: a novel Tabular Foundation Model with Semantically Target-Aware Representations, [^2] Code is available at https://github.com/alanarazi7/TabSTAR.
-designed explicitly for end-to-end handling of purely textual features.
+In this paper, we introduce TabSTAR: a novel Tabular Foundation Model with Semantically Target-Aware Representations, designed explicitly for end-to-end handling of purely textual features.
 By integrating an unfrozen text encoder at its core, TabSTAR can optimize free-text feature representations, demonstrating their clear superiority over alternative frozen embedding approaches.
@@
 TabSTAR’s training is highly efficient [^3] Pretraining within 48 hours on a single A40 GPU.
-Finetuning with PEFT for low memory footprint.
+Finetuning with PEFT for a low memory footprint.
 and its performance steadily improves with more pretraining data.
-Empirically, TabSTAR achieves state-of-the-art (SOTA) performance on classification benchmarks containing substantial textual content, demonstrating significant advances over GBDTs and leading TFMs.
+Empirically, TabSTAR achieves state-of-the-art (SOTA) performance on classification datasets containing textual features, surpassing leading TFMs as well as GBDTs tuned for 4 hours.
 2 Related Work
-This section reviews prior work in five areas relevant to our approach.
+This section reviews prior work in supervised tabular learning.
 We begin with deep learning methods tailored for tabular data, which were applied to a single dataset.
@@
 Finally, we review recent work on applying large language models (LLMs) to tabular data and elaborate on existing AutoML [34] multimodal solutions.
+As we focus on supervised learning, we do not cover self-supervised methods [82, 6] and their downstream applications.
 Deep Learning on a Single Tabular Dataset
@@
 Recent works such as TabM [29] and RealMLP [39] focused on refining MLPs without an attention component.
-Despite these innovations, single-dataset deep learning models have not yet convincingly outperformed GBDTs [62, 30, 61].
-Furthermore, none of them addressed the challenge of modeling tabular datasets with rich textual features.
+Despite these innovations, single‐dataset deep learning models have not yet convincingly outperformed GBDTs [65, 31, 64].
+Furthermore, none of them addressed the challenge of modeling tabular data with rich textual features.
 Cross-Dataset Transfer Learning
-Deep learning was proven to shine when performing transfer learning in many machine learning domains [87].
+Deep learning has been proven to shine when performing transfer learning in many machine learning domains [91].
 Motivated by this success, [52, 89] proved that cross-dataset learning can boost single-dataset performance, but were limited to strict requirements such as partial overlap of feature names.
@@
 TFMs represent an emerging paradigm in tabular learning.
-While the definition is still evolving, we adopt the framing proposed by [74], which identifies key desired characteristics of TFMs: large-scale pretraining with adaptability to downstream tasks, mixed-type column support, cross-domain generalization, use of textual metadata, [^4] Contextual information such as the dataset description, column names and category names.
+While the definition is still evolving, we adopt the framing proposed by [77], which identifies key desired characteristics of TFMs: large-scale pretraining with adaptability to downstream tasks, mixed-type column support, cross-domain generalization, use of textual metadata, [^4] Contextual information such as the dataset description, column names, and category names.
 and column-order invariance.
-TabPFN [35] is recognized as the first TFM, and its successor TabPFN-v2 [37] currently sets the SOTA in tabular learning, becoming a popular approach for TFMs [58, 20].
+TabPFN [36] is recognized as the first TFM, and its successor TabPFN-v2 [38] currently sets the SOTA in tabular learning, becoming a popular approach for TFMs [61, 21, 54].
 TabPFN-v2 was the first model to consistently outperform GBDTs on medium-sized datasets, by pretraining Bayesian Prior-Data Fitted Networks (PFNs) [58] on 130 million synthetic datasets.
 Using ICL at inference time, it accepts up to 10,000 examples as input and predicts without updating its weights.
-Nonetheless, similarly to GBDTs, TabPFN-v2 uses off-the-shelf embeddings for text features, limiting its effectiveness.
-CM2 [85], CARTE [45] and TP-BERTa [80] represent a shift toward semantic tabular modeling, leveraging textual signals and external knowledge at a greater scale.
+TabPFN-v2 inspired TabICL [61] which improved scalability, and TabDPT [54], trained on real data.
+However, they all rely on off-the-shelf embeddings for text features like GBDTs, limiting their performance.
+CM2 [86], CARTE [47], and TP-BERTa [84] represent a shift toward semantic tabular modeling, leveraging textual signals and external knowledge at a greater scale.
 Unlike prior methods, these models transfer knowledge via language representations.
-CM2 pretrained on over 2,000 datasets, but did not focus on free-text features and used static word embeddings without further finetuning them.
+CM2 was pretrained on over 2,000 datasets, but did not focus on free-text features and used static word embeddings without further finetuning them.
 CARTE encodes tables as star-shaped graphs, jointly representing features by their names and values, and applies attention over the graph to capture contextual relations.
@@
 During the past years, several research attempts have tried to combine LLMs and tabular data.
-One line of work is on using LLMs directly for tabular prediction, by converting tabular data into serialized text.
+One line of work focuses on using LLMs directly for tabular prediction by converting tabular data into serialized text.
 TabLLM [35] assessed LLMs under few-shot scenarios, while Tabula-8b [24] finetuned the Llama 3-8B model extensively on tabular data.
@@
 While current generations of LLMs weren’t adopted for tabular learning, their emergent knowledge from their pretraining could be crucial when textual features are present [77, 20].
-Additionally, LLMs can be used in multiple aspects of tabular learning, as they seem to be promising synthetic data generators [9, 65], useful data cleaners [6] and clever feature engineers [36].
+Additionally, LLMs can be used in multiple aspects of tabular learning, as they seem to be promising synthetic data generators [10, 68], useful data cleaners [7], and clever feature engineers [37].
 Multimodal AutoML
@@
 In this section, we introduce TabSTAR: a Tabular Foundation Model with Semantically Target-Aware Representations.
-Our training framework consists of two stages: (1) Pretraining, where it is pretrained over a corpus of tabular datasets [^7] Ranging from metadata-rich, text-heavy datasets to numeric-only tables lacking column names.
+Our training framework consists of two stages: (1) Pretraining, where the model is pretrained over a corpus of tabular datasets [^7] Ranging from metadata-rich, text-heavy datasets to numeric-only tables lacking column names.
 in a multi-task regime, mixing classification with regression tasks, then (2) Finetuning, where the pretrained model is further trained with LoRA [42] on a single downstream task.
 TabSTAR is designed to enable effective cross-dataset learning by applying supervised learning on the target variable in both stages.
-At its core, is uses an unfrozen encoder-only language model, which can invoke world knowledge acquired during the language model pretraining.
+At its core, it uses an unfrozen encoder-only language model, which can potentially invoke world knowledge acquired during the language model pretraining.
 [^8] Note that the language-model pretraining occurs before TabSTAR’s pretraining.
@@
 The encoder is combined with a tabular-specific architecture tailored to structured data, mitigating the known limitations of language models in tabular settings [75, 77].
-TabSTAR’s architecture comprises five core modules: (1) Verbalization, mapping every feature into a textual representation composed of both the column name and value, with a special treatment to numerical features for full numerical precision; (2) Encoding, transforming semantic and numerical data into meaningful embeddings of the same dimension; (3) Fusion, integrating textual and numerical representations; (4) Interaction, modeling dependencies and relationships between different elements through self-attention and cross-element interactions; and (5) Prediction, where Interaction’s outputs are projected into a real value for regression or a probability distribution for classification.
-Figure 1 illustrates the architecture while Appendix A elaborates, and Appendix B discusses the training.
+TabSTAR’s architecture comprises five core modules: (1) Verbalization, mapping every feature into a textual representation composed of both the column name and value, with a special treatment to numerical features for full numerical precision; (2) Encoding, transforming semantic and numerical inputs into meaningful embeddings of the same dimension; (3) Fusion, integrating textual and numerical representations on each feature independently; (4) Interaction, modeling dependencies and relationships between different features through cross-feature self-attention; and (5) Prediction, where outputs are projected into a real value for regression or a probability distribution for classification.
+Figure 1 illustrates the architecture, Appendix A elaborates it, and Appendix B discusses the training.
 A key innovation of TabSTAR is the introduction of target-aware tokens, a novel approach that integrates the target variable’s identity as an input to the model.
@@
 Age 45 “Age: 40–50 (Quantile 50–60%)” $0.27$
-Department Cardiology “Department: Cardiology” 0-
-Report Mild chest discomfort. “Report: Mild chest discomfort.” 0-
-Decision Hospitalized “Target. Decision: Hospitalized” 0-
-Decision Released “Target. Decision: Released” 0-
+Department Cardiology “Department: Cardiology”
+Report Mild chest discomfort. “Report: Mild chest discomfort.”
+Decision Hospitalized “Target. Decision: Hospitalized”
+Decision Released “Target. Decision: Released”
 Verbalization
@@
 The numerical input is standardized using z-scores, with outlier clipping at ±3 standard deviations.
-In addition, they are verbalized using quantile-based binning into 10 bins, a novel approach to mitigate the precision loss inherent in language models [72].
+In addition, they are verbalized using quantile discretization into 10 bins, a novel approach to mitigate the precision loss inherent in language models [75].
 Appendix A.1 shows a precise example and §6 discusses different verbalization strategies.
@@
 We employ a pretrained e5-small-v2 [80] embedding model for semantic encoding, chosen for its strong performance on the MTEB benchmark [57] with a relatively modest parameter count.
-By unfreezing half of its layers, the representations are optimized for predicting the target variable, which leads to a significant impact on TabSTAR performance (see §6).
+By unfreezing the upper half of its layers, the representations are optimized for predicting the target variable, which leads to a significant impact on TabSTAR performance (see §6).
 Each verbalization element is encoded independently into a semantic representation, with attention applied between tokens within each sequence element.
 In parallel, we encode standardized numerical values by projecting them into the same dimension using a small MLP.
-For the patient in Table 2, this results in a numerical embedding for Age alongside semantic representations for each of the five verbalizations.
+For the patient in Table 2, this results in a numerical embedding for Age alongside semantic representations for each of the 5 verbalizations.
 Fusion
 To obtain a unified representation for each sequence element, we apply a fusion block consisting of a single encoder-only transformer layer.
-For each numerical feature, the block attends over its numerical and semantic embeddings, producing a fused representation.
+Crucially, each numerical feature is fused independently, as the block attends only to its numerical and semantic embeddings.
 In our running example, the representation of Age now jointly captures both its semantic context (the fact that the value represents age) as well as its numerical value (the patient’s age, 45, or 0.27 after standardization).
@@
 Each input element is now a token, with feature tokens and target tokens all attending to each other.
-Unlike standard language models which integrate positional encoding, the Interaction module’s inputs are order-invariant, a desiredatum for TFMs, as defined by [74].
+Unlike standard language models, which integrate positional encoding, the Interaction module’s inputs are order-invariant, a desideratum for TFMs, as defined by [77].
 The encoder produces contextualized representations for each target value.
 In our example, this yields dedicated embeddings for the Release and Hospitalization target values.
-The role of these representations is to carry information about the likelihood of each value to be the true value.
+The role of these representations is to carry information about how likely each value is to be the true value.
 Prediction
@@
 We manually curate a pretraining corpus of 350 high-quality tabular datasets (253 classification, 97 regression), in a tedious process in which we uncover numerous duplications in the most popular tabular repositories, OpenML [78] and Kaggle, [^10] https://www.kaggle.com/datasets as elaborated by [76].
-Our datasets are sourced from popular benchmarks [24, 45, 22, 21, 30, 60, 52, 31], which have almost non-existing representation of textual tabular datasets.
-Thus, we furthermore increase our corpus, focusing on classification datasets with rich semantic content.
+We begin by sourcing datasets from popular benchmarks [25, 47, 23, 22, 31, 63, 55, 32], but observe that the presence of textual tabular datasets in them is very rare.
+Thus, we furthermore expand our corpus, focusing on classification datasets with rich semantic content.
 See Appendix C for more details.
 Benchmark
-Tabular datasets with free-text have seen little prior research, and accordingly, benchmarks are incredibly rare.
-Therefore, we consider all the possible datasets from AutoML Multimodal Benchmark [60], from the analysis about free-text and high-cardinality features by [31] and from the CARTE paper [45].
-After a deduplication process we end up with 50 datasets.
-However, there are two important limitations: first, the benchmark is heavily biased towards regression, with 36 datasets in total.
-Secondly, 29 out of these 36 datasets were solely contributed by the CARTE benchmark, which focuses more heavily on high-cardinality features as it was pretrained over knowledge graphs.
-While our main motivation is classification tasks with textual features, we decide nevertheless to evaluate on the full set of 50 datasets although it heavily biases towards regression problems and high-cardinality features, rather than classification and free-text (see Appendix D).
+Tabular datasets with free-text have seen little prior research, and accordingly, benchmarks are rare.
+To address this, we compile all available datasets from three sources: (1) the AutoML Multimodal Benchmark [63], (2) the CARTE paper [47], and (3) the analysis of free-text and high-cardinality features by [32].
+After deduplication, our final benchmark includes 50 datasets.
+Despite its breadth, this collection has two key limitations: First, the benchmark is heavily skewed towards regression tasks, with 36 datasets.
+Secondly, 29 out of these 36 datasets were solely contributed by the CARTE benchmark, which focuses more heavily on high-cardinality features, rather than longer texts, as it was pretrained over knowledge graphs.
+While our main motivation is classification tasks with textual features, we decide nevertheless to evaluate on the full set of 50 datasets, although it is biased toward regression problems and high-cardinality features (see Appendix D).
 Baselines
-We compare TabSTAR against a diverse set of baselines.
-For tree-based methods, we evaluate CatBoost [57], XGBoost (XGB) [14], and Random Forest (RF) [11] with the default configuration proposed by [25].
-For CatBoost and XGBoost we consider a tuned version, where hyperparameters are optimized separately for each task using random search with 5-fold cross-validation under a 4-hour budget on 8 CPU cores.
-Among TFMs, we evaluate TabPFN-v2 [37] and CARTE [45].
-For CARTE, we tune only the learning rate for each task, following the original paper.
+We compare TabSTAR against GBDTs: Random Forest [12], LightGBM [46], XGBoost [15], and CatBoost [60]; MLPs: RealMLP [39] and TabM; and TFMs: [29], CARTE [47], TabDPT [54], TabPFN-v2 [38], and TabICL [61] which only supports classification tasks.
+For several models [^11] All MLPs and GBDTs, except the naive Random Forest baseline, used as a weak baseline.
+we consider both default variants as well as tuned ones, where hyperparameters are optimized separately for each task using random search with 5-fold cross-validation under a 4-hour budget on 8 CPU cores.
 Since the public TabPFN-v2 model does not support text, we use their closed-sourced API client.
 [^12] https://github.com/PriorLabs/tabpfn-client For models lacking native support for textual features, we embed text using e5-small-v2 [80], allowing a fair comparison.
-For more details about the hyperparameters for each baseline as well as exclusion of models such as TP-BERTa due to potential leakage concerns, see Appendix E.
+For more details about the hyperparameters for each baseline as well as exclusion of baselines due to potential leakage concerns, see Appendix E.
 Experimental Setup
-Each of the 50 datasets in the benchmarks is evaluated with 10 random train-test splits (90% training, 10% testing), resulting in 500 runs per model.
-While 30 of the datasets have more than 10,000 examples, the evaluated TFMs have strict limitations.
-TabPFN-v2 employs ICL and thus receives as input at most 10,000 examples.
-Although CARTE imposes no size cap, [^12] In their own paper, CARTE was evaluated only over up to 2,048 examples, without scaling guarantees.
-it suffers from inefficiency and no preset configuration, requiring six learning-rate trials and totalling 6,000 slow GPU runs.
-Because of these important limitations, we consider two experiment conditions: (1) 10K: Each model is trained [^13] While TabPFN-v2 isn’t technically trained, we adopt this term for conciseness.
-over at most 10,000 training examples, and (2) Unlimited: We add TabSTAR-Unlimit, CatBoost-Tuned-Unlimit, and XGBoost-Tuned-Unlimit and evaluate them on the full version of the 30 datasets, [^14] We technically cap the amount of examples to 100,000 for computational efficiency.
-while retaining 10K baselines as weaker reference points.
-We exclude the untuned GBDTs and keep the same number of models as in the 10K condition.
+Each of the 50 datasets in the benchmark is evaluated with 10 random train-test splits (90% training, 10% testing), resulting in 500 runs per model.
+While 30 of the datasets have more than 10,000 examples, most TFMs can’t effectively scale beyond it: TabPFN-v2, for example, employs ICL and thus receives as input at most 10,000 examples.
+While CARTE imposes no strict size cap, its tuning is slow, inefficient, and impractical for larger datasets.
+[^13] In their own paper, CARTE was evaluated only over up to 2,048 examples, without scaling guarantees.
+Because of these important limitations, we consider two experiment conditions: (1) 10K: Each model is trained [^14] While ICL-based TFMs aren’t technically trained, we adopt this term for conciseness.
+over at most 10,000 training examples, and (2) Unlimited: We evaluate TabSTAR and the most competitive, scalable baselines on the full version of the 30 datasets.
+[^15] We technically cap the amount of examples to 100,000 for computational efficiency.
 The TabSTAR Training
@@
 Crucially, the whole collection was carefully curated to prevent any data leakage from duplicate or overly similar datasets.
-For finetuning, while dataset-specific hyperparameter tuning can boost performance, we believe that robust out-of-the-box defaults are essential for TFMs and their evaluation, following Tab-PFN-v2’s approach.
-Therefore, we use a default hyperparameters configuration that was found robust over a disjoint set of tabular datasets, as detailed in Appendix B.2.
+As a result, each dataset is evaluated by finetuning the single pretrained variant, which excludes it from its pretraining.
+For finetuning, while dataset-specific hyperparameter tuning can boost performance, we believe that robust out-of-the-box defaults are essential for TFMs and their evaluation.
+Therefore, we use a default hyperparameter configuration that was found robust over a disjoint set of tabular datasets, as detailed in Appendix B.2.
 5 Results
@@
 Following [38], we normalize scores per dataset split to the $[0,1]$ range, using the best and worst model performance as anchors.
-[^15] For a single run, the best model gets 1, the worst gets 0 and the rest are linearly scaled accordingly.
+[^16] For a single run, the best model gets 1, the worst gets 0, and the rest are linearly scaled accordingly.
 The normalized scores are averaged across all runs, with 95% CIs.
-Performance for all models on both conditions are shown in Figure 2 (classification) and Figure 3 (regression).
-Besides their example limit, TabPFN-v2 cannot process 4 datasets (more than 10 target classes or more than 500,000 cells) and CARTE cannot handle 15 (a bug in their PowerTransformer implementation).
+Performance for all models on both conditions is shown in Figure 2 (classification) and Figure 3 (regression).
+For certain datasets, we are unable to execute TabPFN-v2, CARTE, and TabDPT due to model-specific implementation issues or scalability constraints.
 Reported averages for these models are computed only over the datasets where evaluation is feasible.
-Appendix F expands on dataset level performance, head-to-head comparisons and running times.
+Appendix F.1 elaborates on technical limitations, Appendix F.2 on dataset-level performance, and Appendix F.3 on head-to-head comparisons.
 Figure 2: Comparison of normalized scores with 95% CIs between TabSTAR and baseline models in classification tasks, evaluated on up to 10,000 examples (left) and above 10,000 (right).
@@
 This is evident both when restricting the dataset size to 10,000 examples and when using larger datasets in the unlimited condition.
-For the 10K condition, TabSTAR achieves a 0.809 score, performing better than TabPFN-v2 (0.783) and significantly better than GBDTs (0.756 CatBoost-Tuned, 0.744 XGB-Tuned).
-When analyzing head-to-head comparisons (Appendix F.2), TabSTAR outperforms TabPFN-v2 (7/11 datasets), XGB-Tuned (10/14) and CatBoost-Tuned (11/14).
-For the Unlimited condition, TabSTAR-Unlimit achieves a 0.874 score, significantly above the second-best CatBoost-Tuned-Unlimit with 0.734.
-Importantly, all Unlimit variants surpass the 10K ones, emphasizing the importance of scaling.
+For the 10K condition, TabSTAR achieves a 0.83 score, followed by TabM-Tuned (0.73), CatBoost-Tuned (0.69), and TabPFN-v2 (0.67).
+When analyzing head-to-head comparisons (see Appendix F.3), TabSTAR outperforms TabPFN-v2 (8/11 datasets), TabM-Tuned (10/14), and CatBoost-Tuned (12/14).
+For the unlimited condition, TabSTAR-Unlimited achieves a 0.84 score, followed by TabM-Tuned-Unlimited (0.79), and much above the rest.
+Importantly, unlimited variants significantly surpass the 10K ones, emphasizing the importance of models with no scaling restrictions.
 Figure 3: Comparison of normalized scores with 95% CIs between TabSTAR and baseline models in regression tasks, evaluated on up to 10,000 examples (left) and above 10,000 (right).
@@
 This emphasizes the need for better modeling of textual tabular learning, especially since TabPFN-v2 has shown remarkable performance in non-textual tabular datasets, and CARTE set the SOTA for small datasets.
-When analyzing the Unlimited variants, TabSTAR scales well, and surpasses other TFMs which cannot scale, but the gap from GBDTs remains significant.
+In the unlimited setting, TabSTAR scales well and surpasses other TFMs which cannot scale, but the gap from GBDTs remains significant.
 §7 discusses this limitation and suggests promising directions for future generations of TabSTAR to achieve SOTA in regression as well.
+Cost Analysis
+Table 3 reports the computational requirements of TabSTAR and competing baselines, measured by runtime and peak memory usage, for both training and inference.
+To highlight the importance of GPU acceleration when text embeddings are part of the preprocessing, we include a variant of CatBoost restricted to CPU only.
+We exclude TabPFN-v2, whose performance is probably comparable to TabICL, since it is only accessible via an API, and CARTE, whose extensive tuning requirements and prohibitive training costs forced evaluation on less capable hardware.
+We find that TabSTAR incurs modest inference costs, comparable to GPU-accelerated GBDTs, since GPU processing of text embeddings dominates runtime.
+This is evident from the CatBoost-CPU variant, whose inference is nearly ten times slower than TabSTAR.
+Moreover, TabICL and TabDPT are orders of magnitude slower and consume far more memory, emphasizing the scalability limitations for ICL-based models.
+During training, TabSTAR is considerably slower than the baselines.
+However, while fitting a single GBDT is much faster, this advantage quickly diminishes in practice when hyperparameter tuning is needed or GPU acceleration for text embeddings is not available.
+Appendix F.4 provides dataset-level runtimes and F.5 analyzes costs as a function of the feature count.
+Table 3: Cost Analysis. Median training and inference times and peak memory usage on GPU and CPU, aggregated over the 50 datasets of the benchmark with up to 10,000 examples.
+Train Inference
+Model Time (s) GPU (GB) CPU (GB) Time (s) GPU (GB) CPU (GB)
+CatBoost-CPU 360.8 2.2 34.0 2.1
+CatBoost 68.5 1.3 1.5 2.0 1.3 1.5
+LightGBM 39.1 1.3 1.5 2.0 1.3 1.5
+RandomForest 86.1 1.3 1.5 2.5 1.3 1.5
+RealMLP 136.4 1.4 1.7 2.5 1.3 1.7
+TabDPT 10.4 2.2 1.8 161.4 33.8 6.0
+TabICL 42.6 1.2 1.7 13.4 25.4 2.1
+TabM 19.1 1.8 4.4 1.5 1.8 4.3
+TabSTAR 493.2 4.7 1.8 3.0 1.3 1.8
+XGBoost 38.4 1.3 1.5 2.5 1.3 1.5
 6 Analysis
@@
 To answer these questions, we pretrain several variants of TabSTAR for each analysis, limiting ourselves to a subset of the tabular datasets used for the main experiment (see §4).
-Specifically, each variant is pretrained over 256 datasets [^16] Except for variants of Q2, which analyze the effect of number of datasets on pretraining.
+Specifically, each variant is pretrained over 256 datasets [^17] Except for variants of Q2, which analyze the effect of the number of datasets on pretraining.
 including 30 datasets from our benchmark, and evaluated over the remaining 20 datasets (12 regression, 8 classification).
 This reduced setup allows leveraging transfer learning and exploiting our corpus, without the burden of training multiple folds per variant.
-Results are reported with the same normalized metric used in §5, scaling the performance to the $[0,1]$ range.
 Appendix G.1 lists the 20 datasets used for evaluation along with per-dataset results.
-Figure 4: Performance as a function of the number of encoder layers unfrozen: Validation loss during TabSTAR’s pretraining (left) and normalized scores with 95% CIs on the downstream tasks (right).
-Unfreezing even a single encoder layer significantly improves the performance of TabSTAR.
 Q1: The Role of the Encoder Unfreezing
-To investigate whether unfreezing layers of the textual encoder impacts performance, we conduct experiments where we unfreeze varying numbers of the 12 encoder layers during both TabSTAR’s pretraining and finetuning stages.
-[^17] For each variant, the number of unfrozen layers remains the same in both pretraining and finetuning.
+We examine the effect of unfreezing different numbers of textual encoder layers during pretraining.
+In finetuning, we keep all base layers frozen and instead apply LoRA adapters to the same layers that were unfrozen in pretraining.
+[^18] For simplicity, we will refer to them as unfrozen to differentiate them from layers without LoRA.
 Figure 4 shows the validation loss during TabSTAR pretraining (left) and the normalized score on the downstream tasks (right) as a function of the number of unfrozen encoder layers.
@@
 See Appendix G.2 for more details.
-Table 3: Normalized score with 95% CIs by the number of datasets used during TabSTAR pretraining.
-Pretraining Datasets 0 16 64 256
-Classification 0.352 ± 0.086 0.450 ± 0.084 0.558 ± 0.086 0.786 ± 0.076
-Regression 0.338 ± 0.073 0.395 ± 0.068 0.642 ± 0.066 0.811 ± 0.055
+Figure 4: Performance as a function of the number of encoder layers unfrozen: Validation loss during TabSTAR’s pretraining (left) and normalized scores with 95% CIs on the downstream tasks (right).
+Unfreezing even a single encoder layer significantly improves the performance of TabSTAR.
 Q2: The Effect of Pretraining
@@
 To further investigate the effect of the number of pretraining datasets on downstream task performance, we train two additional versions: one pretrained on 16 datasets and another on 64 datasets.
-As shown in Table 3, increasing the number of pretraining datasets consistently improved performance in both classification and regression tasks.
-Notably, the substantial gain in regression tasks suggests that TabSTAR’s downstream performance on §5 could improve with more pretraining data (see Appendix G.3).
+Figure 5: Average performance on downstream tasks as a function of the number of pretraining datasets (in log scale). We use AUROC for classification (left), and $R^{2}$ for regression (right).
+As shown in Figure 5, increasing the number of pretraining datasets consistently improved performance in both classification and regression tasks.
+Notably, the substantial gain in regression tasks suggests that TabSTAR’s downstream performance on §5 could improve with more pretraining data, potentially reaching SOTA performance with enough scale.
+See Appendix G.3 for more details.
 Q3: Numerical Verbalization
 A key challenge in integrating language models with numerical data is determining how to best represent numerical values within a linguistic framework.
-While some semantic tabular methods omit numerical features from the verbalization [79, 82], TP-BERTa [80] introduced Relative Magnitude Tokenization [80], which encode numerical information through non-semantic special bin tokens.
-In constrast, TabSTAR injects semantic numerical information into the verbalization of numerical features, as illustrated in Table 2.
+While some semantic tabular methods omit numerical features from the verbalization [83, 86], TP-BERTa [84] introduced Relative Magnitude Tokenization [84], which encodes numerical information through non-semantic special bin tokens.
+In contrast, TabSTAR injects semantic numerical information into the verbalization of numerical features, as illustrated in Table 2.
 To quantify the effect of our novel verbalization, we explore two thinner variants: (1) Name + Bin, which excludes the quantile information, and (2) Name, which omits numeric information entirely and verbalizes the feature name only.
@@
 For model scaling, we envision a family of model sizes, common for LLMs [30, 74, 51], that will allow a trade-off between quality and costs.
-Data scaling might leverage self-supervised learning [59] over large-scale table corpora [17], or realistic synthetic tabular data generators [9], which have proven successful [37, 2].
+Data scaling might leverage self-supervised learning [62, 82] over large-scale table corpora [18], or realistic synthetic tabular data generators [10], which have proven successful [38, 2].
 At scale, it could potentially unlock few-shot learning capabilities and develop automatic feature-engineering skills [37].
-Beyond scaling, TabSTAR’s semantic approach has tremendous potential to explicitly include world knowledge, by leveraging LLMs which to date have had limited impact on tabular learning.
-As a few motivating examples, LLMs could improve TabSTAR’s numerical verbalization binning approach by providing semantically informed thresholds, or provide explicit, contextual world knowledge that could be injected as a strong prior in small data scenarios.
+Beyond scaling, TabSTAR’s semantic approach has tremendous potential to explicitly include world knowledge by leveraging LLMs, which to date have had a limited impact on tabular learning.
+As a few motivating examples, LLMs could improve TabSTAR’s numerical verbalization binning approach by providing semantically informed thresholds, or by providing explicit world knowledge that could be injected as a strong prior in small data scenarios.
 While these directions seem like plausible research paths, they come with a risk of data leakage due to the memorization properties of LLMs [9].
@@
 While TabSTAR sets a new bar in classification, its regression results lag behind GBDTs, which outperform other TFMs as well.
-This gap could be narrowed through additional scaling, and also by exploring regression-via-classification techniques like [37, 2].
+This gap could be narrowed through additional scaling, by exploring regression-via-classification techniques like [38, 2], and by enriching the numerical encoder with distribution-aware statistics per feature as done by [61].
+In addition, similar to other TFMs, TabSTAR may encounter memory bottlenecks on datasets with hundreds of features, and its training speed lags behind untuned GBDTs and ICL-based TFMs.
+Yet TabSTAR achieves GBDT-level efficiency during inference, as opposed to TabPFN-v2 and TabICL.
 Furthermore, TabSTAR has not been extensively evaluated in few-shot scenarios and in purely numerical datasets.
-[^19] Partly because of the computational burden of tuning baselines, and the lack of objective leaderboards [73].
-In addition, it demands more compute compared to GBDTs, and it may struggle with memory constraints on datasets containing hundreds of features.
 Despite these limitations, TabSTAR offers a promising pathway toward improving performance on tabular datasets with textual fields, common in industries with high social impact (e.g., healthcare, education), or with significant economic value (e.g., banking, manufacturing).
-We believe TabSTAR paves the way for a new generation of semantically enriched tabular models, and we welcome the research community’s innovations built on this foundation.
+In addition, its architecture lends itself to multimodality and could be extended to tabular datasets that combine numerical, textual, and image features.
+We believe TabSTAR paves the way for a new generation of semantically enriched, Multimodal TFMs, and invite the research community to advance this vision.
 Acknowledgments and Disclosure of Funding
-We thank Omri Feldman for brainstorming since the very beginning; Elad Hoffer and Ofir Lindenbaum for consulting and feedback; David Holzmüller and Myung Kim for supporting evaluations; and Noah Hollmann, Léo Grinsztajn, and the Prior Labs team for providing extensive access to TabPFN-v2.
+Roi Reichart and Eilam Shapira have been partially supported by a VATAT grant on data science.
+We thank Omri Feldman for brainstorming since the very beginning; Elad Hoffer and Ofir Lindenbaum for consulting and feedback; David Holzmüller, Lennart Purucker, Myung Kim, and Gaël Varoquaux for assisting with evaluations and benchmarks; and Frank Hutter, Noah Hollmann, Léo Grinsztajn, and the rest of the Prior Labs team for providing extensive access to TabPFN-v2’s API version.
 Appendix A Architecture
@@
 Numerical Features
-[…]
+Numerical features are represented by both a numerical and a semantic representation.
+For the numerical representation, given a value $x$, we compute the clipped z-score $z^{\prime}=\operatorname{clip}\bigl((x-\mu)/\sigma,\,-3,\,3\bigr)$ where $\mu,\sigma$ are the training set mean and the standard deviation, and missing values are set to 0.
+For the semantic representation, we build $B=10$ quantile bins over the training distribution to map the value accordingly.
+Table 5 shows an illustrative example for the feature Age from our running example in Table 2.
 Semantic Features
@@
 Interaction
-The fused sequence of $e$ tokens is processed by a standard Transformer Encoder with model dimension $d=384$, $L=6$ layers, $6$ attention heads per layer, feed-forward size $4d$, dropout $0.1$, ReLU activation and using a pre-norm configuration.
+The fused sequence of $e$ tokens is processed by a standard Transformer Encoder with model dimension $d=384$, $L=6$ layers, $6$ attention heads per layer, feed‐forward size $4d$, dropout $0.1$, ReLU activation and using a pre-norm configuration.
 Unlike in language modeling, feature ordering is irrelevant, so no positional encodings are used.
@@
 The Fusion Block’s Mechanism
-For the fusion block, we consider two simpler alternatives to the attention mechanism, both of then underperforming: (1) Concatenation, by concatenating the semantic and numerical $d$-dimensional vectors into a $2d$-dimensional vector, and projecting them back via an MLP, and (2) Multiplication, by multiplying the semantic representation directly with the numerical value [^22] After rescaling it to be centered around 1 rather than 0, using a learned scaling factor.
-in a straightforward, parameter-free manner as in [79, 82].
+For the fusion block, we consider two simpler alternatives to the attention mechanism, both of them underperforming: (1) Concatenation, by concatenating the semantic and numerical $d$-dimensional vectors into a $2d$-dimensional vector, and projecting them back via an MLP, and (2) Multiplication, by multiplying the semantic representation directly with the numerical value [^22] After rescaling it to be centered around 1 rather than 0, using a learned scaling factor.
+in a straightforward, parameter‐free manner as in [83, 86].
 The Number of Interaction Layers
@@
 We observe no substantial gain, and thus we prioritize the simpler solution, as row-level attention is sensitive to the batch size and adds complexity to inference time.
+Verbalization
+We experiment with various verbalization strategies.
+For features, including the column name alongside the value consistently outperforms using the value alone.
+For target-aware tokens, we find that explicitly verbalizing all possible target values yields better performance than omitting them.
+In contrast, adding a dataset-level description prefix to improve contextualization offers no measurable advantage.
+Interestingly, TabSTAR maintains high performance even when no semantic metadata is present.
+We see this robustness as a strength.
 Appendix B Training
@@
 Therefore, we observe no effect from changing the mini-batch size when keeping the global batch size fixed to 128.
-We tune 1,597,440 out of TabSTAR’s 47,263,874 parameters (3.4%).
+We tune 1,597,440 out of TabSTAR’s 47,263,874 parameters (3.4%), spanning all blocks of the architecture.
+The frozen layers of the text encoder do not receive LoRA adapters.
 Finetuning hyperparameters are selected by pretraining TabSTAR over 256 datasets and performing grid-search over a held-out set of 25 downstream tasks disjoint from the 50 datasets in the benchmark evaluated in §4.
@@
 CARTE natively supports textual inputs, and the TabPFN-v2 API client [^28] https://github.com/PriorLabs/tabpfn-client does as well, although its implementation details remain undisclosed.
-On the other hand, GBDTs do not natively support free-text features, [^29] CatBoost includes a built-in text module, but it underperforms compared to dense text embeddings., so we preprocess these features into fixed-size embeddings using skrub’s TextEncoder, which internally applies a frozen e5-small-v2 encoder to each semantic column.
+As all other baselines do not natively support free-text features, [^29] CatBoost includes a built-in text module, but it underperforms compared to dense text embeddings.
+we preprocess these features into fixed-size embeddings using skrub’s TextEncoder, which internally applies a frozen e5-small-v2 encoder to each semantic column.
 This aligns with the encoder used in TabSTAR, enabling a fair comparison across models.
@@
 This compressed representation performs comparably to the full embedding space, while offering improved inference efficiency.
+Hyperparameter Tuning
+For the tuned models we use the Optuna package [^30] https://pypi.org/project/optuna/ with random search, with a budget of 4 hours for every run and parallelizing trials on 8 CPU cores.
+We use 5-fold cross-validation and take the best configuration selected based on this mean score.
+We use it then to retrain the model on the full training data.
 E.2.1 TabPFN-v2
@@
 While the intrinsic details of their textual handling remain undocumented, it’s reasonable to assume that it resembles the processing we apply to GBDTs, as their model leverages ICL and their architecture has no textual encoder.
-E.2.2 CARTE
+E.2.2 TabICL
+We run TabICL using its package [^32] https://pypi.org/project/tabicl/ with the default configuration.
+Since TabICL can only be evaluated on classification tasks, we exclude it from the regression analysis.
+E.2.3 TabDPT
+We run TabDPT using its package [^33] https://pypi.org/project/tabdpt/ with the default configuration.
+While it achieves state-of-the-art performance on some regression tasks, we encounter severe performance issues on others.
+Due to this inconsistency, we exclude TabDPT from the reported results and are in contact with the authors to investigate potential implementation issues.
+E.2.4 CARTE
 […]
-E.2.3 CatBoost and CatBoost-Tuned
-We run CatBoost using the catboost package [^33] https://pypi.org/project/catboost/ and run the default configuration suggested by [25] by setting $early\_stopping\_rounds=50$, $od\_pval=0.001$, $iterations=2000$.
-For the tuned version, we use the Optuna package [^34] https://pypi.org/project/optuna/ with random search, with a budget of 4 hours for every run and parallelizing trials on 8 CPU cores.
-We use 5-fold cross-validation and take the best configuration selected based on this mean score.
-We use it then to retrain the model on the full training data.
-For hyperparameter space, we follow the hyperparameter search suggested by [37] as detailed in Table 15.
+E.2.5 RealMLP
+We run RealMLP using its official implementation in the pytabkit package.
+[^36] https://pypi.org/project/pytabkit/ Following the advice of its authors, we disable label smoothing and optimize for cross_entropy for binary classification and 1-auc_ovr for multiclass classification, and keep all other default hyperparameters.
+For the tuned version, we follow the hyperparameter search space suggested by [39] as detailed in Table 15.
+Table 15: RealMLP-Tuned Hyperparameters Search Space
+Hyperparameter Search Space
+num_emb_type Choice([None, PBLD, PL, PLR])
+add_front_scale Choice([True, False], p=[0.6, 0.4])
+lr $\log\mathcal{U}(2e-2,3e-1)$
+p_drop Choice([0.0, 0.15, 0.3], p=[0.3, 0.5, 0.2])
+act Choice([ReLU, SELU, Mish])
+hidden_sizes Choice([[256, 256, 256], [64, 64, 64, 64], [512]], p=[0.6, 0.2, 0.2])
+wd Choice([0.0, 2e-2])
+plr_sigma $\log\mathcal{U}(0.05,0.5)$ if classification else 0.1
+E.2.6 CatBoost
+We run CatBoost using the catboost package [^37] https://pypi.org/project/catboost/ and run the default configuration suggested by [26] by setting early_stopping_rounds=50early\_stopping\_rounds=50, od_pval=0.001od\_pval=0.001, $iterations=2000$.
+For the tuned version, we follow the hyperparameter search suggested by [38] as detailed in Table 16.
 Table 16: CatBoost-Tuned Hyperparameters Search Space
 Hyperparameter Search Space
-lr $\log\mathcal{U}(e^{-5},1)$
+learning_rate $\log\mathcal{U}(e^{-5},1)$
 random_strength $\mathcal{U}\{1,2,\ldots,20\}$
@@
 iterations $\mathcal{U}\{100,101,\ldots,4000\}$
-E.2.4 XGBoost and XGBoost-Tuned
-We run XGBoost using the xgboost package [^35] https://pypi.org/project/xgboost/ and follow the same procedure as for CatBoost, except additional preprocessing (e.g., transforming categorical variables into one-hot encodings).
-For the default configuration, we follow the suggestion of [25] and use: $booster="gbtree"$, $early\_stopping\_rounds=50$, $n\_estimators=2000$.
+E.2.7 XGBoost
+We run XGBoost using the xgboost package.
+[^38] https://pypi.org/project/xgboost/ For the default configuration, we follow the suggestion of [26] and use: booster=gbtreebooster="gbtree", early_stopping_rounds=50early\_stopping\_rounds=50, n_estimators=2000n\_estimators=2000.
 For the tuned variant, we follow the hyperparameter search space suggested by [38], as shown in Table 17.
@@
 n_estimators $\mathcal{U}\{100,101,\ldots,4000\}$
-E.2.5 Random Forest
+E.2.8 LightGBM
+We run LightGBM using the lightgbm package.
+[^39] https://pypi.org/project/lightgbm/ and its default implementation.
+For the tuned variant, we follow the hyperparameter search space suggested by [38], as shown in Table 18.
+Table 18: LightGBM-Tuned Hyperparameters Search Space
+Hyperparameter Search Space
+num_leaves $\mathcal{U}\{5,6,\ldots,50\}$
+max_depth $\mathcal{U}\{3,4,\ldots,20\}$
+learning_rate $\log\mathcal{U}(e^{-3},1)$
+n_estimators $\mathcal{U}\{50,51,\ldots,2000\}$
+min_child_weight $\{10^{-5},10^{-3},10^{-2},10^{-1},1,10,10^{2},10^{3},10^{4}\}$
+subsample $\mathcal{U}(0.2,0.8)$
+colsample_bytree $\mathcal{U}(0.2,0.8)$
+reg_alpha $\{0,10^{-1},1,2,5,7,10,50,100\}$
+reg_lambda $\{0,10^{-1},1,5,10,20,50,100\}$
+E.2.9 Random Forest
 We treat Random Forest as a weak baseline to establish a lower-bound reference for each dataset split.
-We run it with the sklearn package  [^36] https://scikit-learn.org/ and use its default configuration with $n\_estimators=100$.
+We run it with the sklearn package  [^40] https://scikit-learn.org/ and use its default configuration with n_estimators=100n\_estimators=100.
 Appendix F Extended Main Results
@@
 The final reported performance for each model is the average over these normalized runs, and we compute 95% confidence intervals using the standard normal approximation: $\hat{\mu}\pm 1.96\,\frac{\hat{\sigma}}{\sqrt{n}}$.
-In this section, we often use abbreviated model names for conciseness: We refer to CatBoost variants as CatB and CatB-T, with the latter being the Tuned version.
-Similarly, we use XGB and XGB-T for XGBoost.
-We maintain the abbreviation of Random Forest (RF) and shorten TabPFN-v2 to simply TabPFN.
-For models that run on an unlimited number of training examples, we add a -U suffix, i.e.
-TabSTAR-U, CatB-T-U, and XGB-T-U.
-F.1 Dataset Level Performance
+F.1 Technical Limitations
+As discussed in §5, TabPFN-v2 is unable to run on 4 datasets: C12, because it is a multiclass problem with more than 10 classes, and C02, C14 and R01 because they support inference for up to 500,000 cells.
+Attempts to run the model over a subset of the examples led to a significantly worse performance, and thus we decide not to report them to allow a fair comparison.
+Additionally, CARTE is unable to run over 15 of the datasets in the benchmark due to a known bug [^41] https://github.com/soda-inria/carte/issues/23 in their implementation for the PowerTransformation, which struggles in the presence of features with too less unique values.
+Furthermore, TabDPT’s performance over a few regression datasets is especially low.
+Since we focus on classification tasks, we exclude this model’s performance on this task.
+Finally, RealMLP, fails to run on R01 in the unlimited setting due to memory constraints.
+F.2 Dataset Level Performance
 We report AUROC for classification and $R^{2}$ for regression, with 95% CIs computed over the 10 runs for each dataset.
-Tables  17 and 18 summarize classification performance on datasets with up to 10K and over 10K examples, respectively.
-Tables 19 and 20 to the same for regression tasks.
+Tables  19 and 20 summarize classification performance on datasets with up to 10K and over 10K examples, respectively.
+Tables LABEL:tab:reg_10k_dataset_performance and 22 to the same for regression tasks.
 For conciseness, datasets are referred by their ID from Appendix D.
-As discussed in §5, TabPFN-v2 is unable to run over 4 datasets: C12, because it is a multiclass problem with more than 10 classes, and C02, C14 and R01 because they support inference for up to 500,000 cells.
-Attempts to run the model over a subset of the examples led to a significantly worse performance, and thus we decide not to report them to allow a fair comparison.
-Furthermore, CARTE is unable to run over 15 of the datasets in the benchmark due a known bug [^37] https://github.com/soda-inria/carte/issues/23 in their implementation for the PowerTransformation, which struggles in the presence of features with too less unique values.
-Table 17: Classification performance per dataset (up to 10K). The top performance score is bolded first, and then all scores are rounded. We report average AUROC with 95% CIs.
-ID CARTE CatB CatB-T RF TabPFN TabSTAR XGB XGB-T
-C01 88.4±0.3 90.2±0.3 90.3±0.4 88.8±0.3 90.3±0.3 90.8±0.3 89.3±0.4 90.2±0.4
-C02 97.3±0.5 97.4±0.4 96.3±0.5 97.9±0.5 97.2±0.3 97.6±0.3
-C03 82.5±0.3 82.0±0.3 81.2±0.6 77.3±0.3 82.4±0.3 83.0±0.3 80.5±0.3 82.1±0.3
-C04 84.5±1.9 85.4±1.8 82.4±3.0 88.3±1.5 83.7±2.3 80.7±2.9 85.4±2.0
-C05 88.5±0.9 90.6±1.0 90.9±0.5 88.0±1.1 91.2±0.7 91.3±0.8 88.1±1.1 90.3±0.6
-C06 80.9±1.5 81.3±1.3 82.6±1.1 73.6±1.2 87.7±0.5 87.0±0.6 81.5±1.0 83.7±0.9
-C07 90.1±0.4 90.3±0.3 90.6±0.3 85.1±0.7 89.8±0.4 91.5±0.3 87.2±0.6 89.3±0.5
-C08 90.8±1.6 93.0±0.9 93.2±1.0 90.2±1.3 91.3±1.1 93.5±1.5 91.9±1.1 94.4±0.7
-C09 99.8±0.1 99.8±0.1 99.6±0.2 99.8±0.1 99.3±0.1 99.7±0.1 99.8±0.1
-C10 86.7±0.2 87.0±0.2 84.6±0.2 87.6±0.4 89.1±0.2 85.4±0.3 86.8±0.2
-C11 78.4±0.5 79.7±0.5 80.7±0.4 76.2±0.6 81.4±0.4 79.2±0.5 78.2±0.6 80.2±0.5
-C12 96.0±0.2 97.5±0.1 97.7±0.1 95.3±0.2 98.3±0.1 96.6±0.2 97.3±0.1
-C13 70.9±0.8 73.7±1.1 74.1±1.0 71.1±1.1 72.3±0.9 75.0±0.7 70.2±0.8 74.1±1.1
-C14 83.5±0.3 84.0±0.5 80.1±0.3 84.0±0.3 81.4±0.4 83.2±0.4
-Table 18: Classification performance per dataset (above 10K). The top performance score is bolded first, and then all scores are rounded. We report average AUROC with 95% CIs.
-ID CatB-T CatB-T-U TabPFN TabSTAR TabSTAR-U XGB-T XGB-T-U
-C01 90.3±0.4 90.5±0.4 90.3±0.3 90.8±0.3 91.2±0.3 90.2±0.4 90.4±0.3
-C02 97.4±0.4 98.3±0.3 97.9±0.5 98.4±0.2 97.6±0.3 98.2±0.3
-C03 81.2±0.6 81.5±0.8 82.4±0.3 83.0±0.3 83.8±0.5 82.1±0.3 82.3±0.3
-C07 90.6±0.3 91.0±0.4 89.8±0.4 91.5±0.3 91.9±0.3 89.3±0.5 89.9±0.5
-C08 93.2±1.0 93.1±1.2 91.3±1.1 93.5±1.5 95.1±0.9 94.4±0.7 94.4±0.9
-C09 99.8±0.1 99.8±0.1 99.8±0.1 99.3±0.1 99.5±0.1 99.8±0.1 99.8±0.1
-C11 80.7±0.4 81.8±0.4 81.4±0.4 79.2±0.5 81.0±0.5 80.2±0.5 81.4±0.4
-C12 97.7±0.1 98.5±0.1 98.3±0.1 99.1±0.1 97.3±0.1 98.4±0.1
-C13 74.1±1.0 76.7±0.8 72.3±0.9 75.0±0.7 77.9±1.0 74.1±1.1 76.9±0.9
-C14 84.0±0.5 84.8±0.4 84.0±0.3 85.1±0.3 83.2±0.4 84.2±0.4
-Table 19: Regression performance per dataset. The top performance score is bolded first, and then all scores are rounded. We report average $R^{2}$ with 95% CIs.
-ID CARTE CatB CatB-T RF TabPFN TabSTAR XGB XGB-T
-R01 100±0.0 100±0.0 100±0.0 100±0.0 100±0.0 99.9±0.1 100±0.0
-R02 98.3±1.0 98.3±0.9 98.0±1.0 98.4±0.9 98.0±1.1 97.9±1.1 98.3±0.9
-R03 71.8±0.5 73.8±0.4 74.1±0.3 68.4±0.6 74.9±0.4 70.9±0.8 69.1±0.5 73.7±0.5
-R04 86.2±7.4 85.1±7.1 85.1±6.8 86.1±5.2 81.5±8.7 85.8±7.5 87.5±3.7
-R05 93.2±0.8 89.5±1.3 90.0±1.2 86.7±1.3 92.7±0.9 93.2±1.2 87.1±1.4 90.1±1.3
-R06 97.7±0.8 98.1±0.5 98.2±0.5 97.2±1.0 98.5±0.6 98.1±0.3 97.1±0.7 97.9±0.7
-R07 71.6±1.3 66.9±0.8 67.8±0.8 61.5±1.2 62.2±1.3 71.1±1.8 61.9±1.5 68.8±0.8
-R08 93.2±0.8 93.0±0.7 93.2±0.7 92.2±0.7 93.9±0.7 92.8±0.6 92.3±0.7 93.1±0.7
-R09 89.2±0.5 89.2±0.6 89.3±0.4 88.6±0.7 89.8±0.5 88.8±0.5 88.2±0.7 89.3±0.5
-R10 99.5±0.2 99.0±0.6 98.9±0.6 98.1±0.3 99.1±0.3 99.5±0.1 98.6±0.3 99.4±0.2
-R11 94.2±0.9 94.0±0.9 92.5±1.3 94.5±0.9 93.9±1.1 93.3±1.1 94.2±0.8
-R12 98.5±0.3 98.5±0.2 98.0±0.3 98.5±0.2 98.3±0.2 98.2±0.3 98.5±0.2
-R13 52.8±2.1 57.6±2.7 58.2±2.6 51.8±3.5 53.8±2.7 53.4±3.0 49.0±3.3 57.5±2.7
-R14 96.6±0.2 97.3±0.1 97.3±0.1 96.8±0.1 96.2±0.2 96.4±0.1 97.1±0.1 97.1±0.1
-R15 79.5±0.9 79.9±1.1 77.3±1.2 79.4±1.0 78.9±1.5 76.8±1.2 80.3±1.0
-R16 98.7±0.0 98.7±0.0 98.7±0.0 97.7±0.1 98.8±0.0 98.7±0.0 97.7±0.1 97.8±0.1
-R17 95.3±2.0 94.8±2.4 95.2±2.0 94.2±2.5 95.1±2.1 94.7±2.3 94.5±2.6 95.4±1.9
-R18 98.1±0.4 98.2±0.4 98.2±0.4 97.7±0.4 93.1±1.3 97.4±0.6 97.5±0.8 98.3±0.4
-R19 85.5±0.8 85.8±0.9 85.3±0.9 84.0±1.1 83.8±1.1 83.9±1.0 82.4±2.0
-R20 96.4±0.5 96.0±0.6 96.2±0.5 95.7±0.6 93.9±1.0 95.6±0.9 95.3±0.9 96.2±0.5
-R21 92.3±1.2 92.5±1.1 92.6±1.1 91.8±1.1 93.3±1.1 92.3±1.0 91.6±1.0 92.6±1.0
-R22 45.2±5.5 45.2±5.9 39.1±6.4 43.8±5.5 39.7±5.5 36.1±7.1 46.8±5.8
-R23 85.0±1.8 85.4±1.2 85.5±1.1 81.8±1.5 84.8±1.6 83.0±1.6 83.3±1.2 85.1±1.2
-R24 86.0±0.6 84.0±0.8 85.8±0.7 79.7±1.0 70.3±1.6 81.8±1.8 82.9±0.7 85.9±0.9
-R25 94.3±0.6 95.4±0.5 95.3±0.5 94.8±0.5 86.7±1.0 94.8±0.5 94.6±0.6 95.4±0.5
-R26 99.8±0.1 99.4±0.1 99.4±0.1 99.0±0.2 99.8±0.1 99.6±0.1 99.1±0.2 99.5±0.1
-R27 81.9±1.6 85.2±1.3 85.3±1.4 84.8±1.3 82.2±1.6 82.0±1.6 83.2±1.2 85.5±1.2
-R28 52.5±2.6 53.4±2.7 53.3±2.8 46.0±2.7 61.7±2.1 51.5±2.9 45.3±2.3 53.5±2.7
-R29 94.4±0.7 94.4±0.7 93.0±1.0 95.7±0.6 94.3±0.9 93.4±0.8 94.5±0.8
-R30 23.8±4.2 22.7±4.1 24.6±3.9 23.5±3.5 20.9±4.8 15.7±5.0 17.5±4.2 25.5±3.7
-R31 92.1±0.3 92.1±0.4 92.0±0.4 89.9±0.5 93.2±0.3 91.7±0.4 90.3±0.5 92.0±0.4
-R32 28.8±6.3 28.2±5.8 28.6±5.9 26.2±5.4 19.4±5.6 22.7±7.3 31.6±6.2
-R33 46.6±1.5 47.4±1.7 47.8±1.7 40.2±1.6 44.9±1.6 46.0±1.8 40.3±2.1 47.9±1.6
-R34 90.9±2.5 91.4±2.1 88.7±3.1 92.3±1.8 89.1±2.9 88.9±3.6 91.6±2.1
-R35 85.9±0.6 84.4±0.5 84.5±0.5 81.2±0.4 85.2±0.7 85.7±0.8 81.6±0.8 84.4±0.3
-R36 96.3±0.9 95.2±0.9 95.5±0.9 94.3±1.0 91.2±1.1 95.7±0.8 94.4±0.9 95.5±0.9
-Table 20: Regression performance per dataset (above 10K). The top performance score is bolded first, and then all scores are rounded. We report average $R^{2}$ with 95% CIs.
-ID CatB-T CatB-T-U TabPFN TabSTAR TabSTAR-U XGB-T XGB-T-U
-R01 100±0.0 100±0.0 100±0.0 100±0.0 100±0.0 100±0.0
-R02 98.3±0.9 98.8±0.9 98.4±0.9 98.0±1.1 98.4±1.0 98.3±0.9 98.8±0.9
-R03 74.1±0.3 75.1±0.4 74.9±0.4 70.9±0.8 72.3±0.6 73.7±0.5 74.6±0.6
-R04 85.1±7.1 91.3±0.8 86.1±5.2 81.5±8.7 91.0±0.7 87.5±3.7 91.3±0.6
-R07 67.8±0.8 85.2±0.3 62.2±1.3 71.1±1.8 79.9±1.8 68.8±0.8 86.1±0.7
-R08 93.2±0.7 93.8±0.6 93.9±0.7 92.8±0.6 93.3±0.6 93.1±0.7 93.9±0.5
-R09 89.3±0.4 89.5±0.4 89.8±0.5 88.8±0.5 89.2±0.6 89.3±0.5 89.5±0.4
-R12 98.5±0.2 98.9±0.1 98.5±0.2 98.3±0.2 98.4±0.2 98.5±0.2 98.9±0.2
-R14 97.3±0.1 97.8±0.1 96.2±0.2 96.4±0.1 96.8±0.2 97.1±0.1 97.5±0.1
-R15 79.9±1.1 86.5±0.7 79.4±1.0 78.9±1.5 83.9±1.4 80.3±1.0 87.0±0.7
-R16 98.7±0.0 98.8±0.0 98.8±0.0 98.7±0.0 98.8±0.0 97.8±0.1 98.0±0.1
-R17 95.2±2.0 97.6±0.6 95.1±2.1 94.7±2.3 97.6±0.6 95.4±1.9 97.7±0.6
-R18 98.2±0.4 99.0±0.2 93.1±1.3 97.4±0.6 97.7±0.6 98.3±0.4 98.9±0.3
-R20 96.2±0.5 96.3±0.4 93.9±1.0 95.6±0.9 95.3±1.0 96.2±0.5 96.3±0.5
-R23 85.5±1.1 86.3±0.7 84.8±1.6 83.0±1.6 85.2±1.7 85.1±1.2 86.1±0.8
-R24 85.8±0.7 97.1±0.3 70.3±1.6 81.8±1.8 95.5±0.7 85.9±0.9 97.3±0.3
-R25 95.3±0.5 95.9±0.4 86.7±1.0 94.8±0.5 95.1±0.5 95.4±0.5 95.8±0.5
-R31 92.0±0.4 93.0±0.3 93.2±0.3 91.7±0.4 92.7±0.4 92.0±0.4 92.9±0.3
-R33 47.8±1.7 55.9±1.1 44.9±1.6 46.0±1.8 57.2±1.3 47.9±1.6 55.7±1.3
-R36 95.5±0.9 95.5±0.9 91.2±1.1 95.7±0.8 95.5±0.5 95.5±0.9 95.5±0.9
-F.2 Head-to-head comparisons
+Table 19: Classification performance per dataset (up to 10K).
+Best score before rounding is bolded.
+We report average AUROC with 95% CIs.
+Models: CARTE (CRT), CatBoost (CTB), TabICL (ICL), LightGBM (LGB), RealMLP (MLP), TabDPT (DPT), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB).
+Tuned models are marked with a ’+’.
+ID CRT CTB CTB+ DPT ICL LGB LGB+ MLP MLP+ PFN RF STR TBM TBM+ XGB XGB+
+C01 88.4 $\pm$0.3 90.2 $\pm$0.3 90.3 $\pm$0.4 90.2 $\pm$0.3 90.7 $\pm$0.4 89.5 $\pm$0.4 89.6 $\pm$0.4 90.2 $\pm$0.5 90.2 $\pm$0.3 90.3 $\pm$0.3 88.8 $\pm$0.3 90.8 $\pm$0.3 90.5 $\pm$0.3 90.5 $\pm$0.3 89.3 $\pm$0.4 90.2 $\pm$0.4
+C02 97.3 $\pm$0.5 97.4 $\pm$0.4 92.3 $\pm$0.6 95.2 $\pm$0.4 97.2 $\pm$0.5 97.3 $\pm$0.5 96.4 $\pm$0.5 96.6 $\pm$0.5 96.3 $\pm$0.5 97.9 $\pm$0.5 96.8 $\pm$0.5 97.2 $\pm$0.4 97.2 $\pm$0.3 97.6 $\pm$0.3
+C03 82.5 $\pm$0.3 82.0 $\pm$0.3 81.2 $\pm$0.6 85.4 $\pm$0.3 81.4 $\pm$0.3 79.8 $\pm$0.3 79.9 $\pm$0.3 82.6 $\pm$0.4 82.7 $\pm$0.3 82.4 $\pm$0.3 77.3 $\pm$0.3 83.0 $\pm$0.3 83.3 $\pm$0.3 83.8 $\pm$0.3 80.5 $\pm$0.3 82.1 $\pm$0.3
+C04 84.5 $\pm$1.9 85.4 $\pm$1.8 83.3 $\pm$2.8 86.9 $\pm$2.3 83.0 $\pm$2.3 82.8 $\pm$2.1 82.4 $\pm$3.1 82.8 $\pm$2.2 88.3 $\pm$1.5 82.4 $\pm$3.0 83.7 $\pm$2.3 85.0 $\pm$3.1 85.1 $\pm$2.6 80.7 $\pm$2.9 85.4 $\pm$2.0
+C05 88.5 $\pm$0.9 90.6 $\pm$1.0 90.9 $\pm$0.5 92.2 $\pm$0.6 92.9 $\pm$0.5 88.2 $\pm$1.3 88.4 $\pm$0.6 91.0 $\pm$0.8 91.4 $\pm$0.6 91.2 $\pm$0.7 88.0 $\pm$1.1 91.3 $\pm$0.8 91.1 $\pm$0.7 90.7 $\pm$0.6 88.1 $\pm$1.1 90.3 $\pm$0.6
+C06 80.9 $\pm$1.5 81.3 $\pm$1.3 82.6 $\pm$1.1 73.5 $\pm$2.2 74.3 $\pm$1.8 81.9 $\pm$1.1 82.8 $\pm$0.9 75.6 $\pm$1.4 77.5 $\pm$1.4 87.7 $\pm$0.5 73.6 $\pm$1.2 87.0 $\pm$0.6 79.0 $\pm$1.7 79.2 $\pm$1.5 81.5 $\pm$1.0 83.7 $\pm$0.9
+C07 90.1 $\pm$0.4 90.3 $\pm$0.3 90.6 $\pm$0.3 89.9 $\pm$0.2 86.7 $\pm$0.3 88.9 $\pm$0.4 89.3 $\pm$0.4 90.0 $\pm$0.3 90.2 $\pm$0.2 89.8 $\pm$0.4 85.1 $\pm$0.7 91.5 $\pm$0.3 91.1 $\pm$0.3 91.1 $\pm$0.3 87.2 $\pm$0.6 89.3 $\pm$0.5
+C08 90.8 $\pm$1.6 93.0 $\pm$0.9 93.2 $\pm$1.0 93.2 $\pm$1.4 95.1 $\pm$0.7 93.1 $\pm$0.7 93.2 $\pm$0.9 92.1 $\pm$0.9 91.9 $\pm$1.3 91.3 $\pm$1.1 90.2 $\pm$1.3 93.5 $\pm$1.5 92.1 $\pm$1.3 92.1 $\pm$1.5 91.9 $\pm$1.1 94.4 $\pm$0.7
+C09 82.4 $\pm$1.3 82.6 $\pm$1.4 81.3 $\pm$1.2 84.3 $\pm$1.0 81.1 $\pm$1.2 82.9 $\pm$1.3 82.6 $\pm$1.4 81.3 $\pm$2.5 82.5 $\pm$1.1 76.4 $\pm$1.4 94.0 $\pm$0.8 81.8 $\pm$4.1 83.6 $\pm$1.0 79.4 $\pm$1.6 83.2 $\pm$1.2
+C10 86.7 $\pm$0.2 87.0 $\pm$0.2 87.2 $\pm$0.2 87.7 $\pm$0.2 85.7 $\pm$0.3 85.9 $\pm$0.3 87.7 $\pm$0.2 87.6 $\pm$0.2 87.6 $\pm$0.4 84.6 $\pm$0.2 89.1 $\pm$0.2 87.5 $\pm$0.3 87.5 $\pm$0.3 85.4 $\pm$0.3 86.8 $\pm$0.2
+C11 78.4 $\pm$0.5 79.7 $\pm$0.5 80.7 $\pm$0.4 79.6 $\pm$0.4 81.5 $\pm$0.5 78.9 $\pm$0.5 79.2 $\pm$0.4 80.4 $\pm$0.4 79.6 $\pm$0.5 81.4 $\pm$0.4 76.2 $\pm$0.6 79.2 $\pm$0.5 81.2 $\pm$0.3 81.5 $\pm$0.3 78.2 $\pm$0.6 80.2 $\pm$0.5
+C12 96.0 $\pm$0.2 97.5 $\pm$0.1 97.7 $\pm$0.1 93.8 $\pm$0.3 97.8 $\pm$0.1 96.7 $\pm$0.2 96.7 $\pm$0.1 97.6 $\pm$0.1 97.6 $\pm$0.1 95.3 $\pm$0.2 98.3 $\pm$0.1 97.8 $\pm$0.1 97.9 $\pm$0.1 96.6 $\pm$0.2 97.3 $\pm$0.1
+C13 70.9 $\pm$0.8 73.7 $\pm$1.1 74.1 $\pm$1.0 70.7 $\pm$0.7 74.0 $\pm$0.8 72.6 $\pm$1.0 72.7 $\pm$1.0 73.0 $\pm$0.7 72.4 $\pm$1.2 72.3 $\pm$0.9 71.1 $\pm$1.1 75.0 $\pm$0.7 74.3 $\pm$0.6 74.4 $\pm$0.7 70.2 $\pm$0.8 74.1 $\pm$1.1
+C14 83.5 $\pm$0.3 84.0 $\pm$0.5 82.4 $\pm$0.3 80.4 $\pm$0.3 81.9 $\pm$0.4 82.3 $\pm$0.4 81.9 $\pm$0.5 83.2 $\pm$0.4 80.1 $\pm$0.3 84.0 $\pm$0.3 84.8 $\pm$0.2 85.4 $\pm$0.3 81.4 $\pm$0.4 83.2 $\pm$0.4
+Table 20: Classification performance per dataset (above 10K).
+Best score before rounding is bolded.
+We report average AUROC with 95% CIs.
+Models: CARTE (CRT), CatBoost (CTB), TabICL (ICL), LightGBM (LGB), RealMLP (MLP), TabDPT (DPT), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB).
+Unlimited models are marked with a ’!’.
+ID CRT CTB CTB! DPT ICL ICL! LGB LGB! MLP MLP! PFN STR STR! TBM TBM! XGB XGB!
+C01 88.4 $\pm$0.3 90.3 $\pm$0.4 90.5 $\pm$0.4 90.2 $\pm$0.3 90.7 $\pm$0.4 90.9 $\pm$0.4 89.6 $\pm$0.4 89.8 $\pm$0.4 90.2 $\pm$0.3 90.6 $\pm$0.4 90.3 $\pm$0.3 90.8 $\pm$0.3 91.2 $\pm$0.3 90.5 $\pm$0.3 90.8 $\pm$0.4 90.2 $\pm$0.4 90.4 $\pm$0.3
+C02 97.4 $\pm$0.4 98.3 $\pm$0.3 92.3 $\pm$0.6 95.2 $\pm$0.4 96.9 $\pm$0.5 97.3 $\pm$0.5 98.3 $\pm$0.3 96.6 $\pm$0.5 98.5 $\pm$0.2 97.9 $\pm$0.5 98.4 $\pm$0.2 97.2 $\pm$0.4 98.5 $\pm$0.3 97.6 $\pm$0.3 98.2 $\pm$0.3
+C03 82.5 $\pm$0.3 81.2 $\pm$0.6 81.5 $\pm$0.8 85.4 $\pm$0.3 81.4 $\pm$0.3 81.2 $\pm$0.3 79.9 $\pm$0.3 80.4 $\pm$0.3 82.7 $\pm$0.3 83.4 $\pm$0.3 82.4 $\pm$0.3 83.0 $\pm$0.3 83.8 $\pm$0.5 83.8 $\pm$0.3 84.1 $\pm$0.4 82.1 $\pm$0.3 82.3 $\pm$0.3
+C07 90.1 $\pm$0.4 90.6 $\pm$0.3 91.0 $\pm$0.4 89.9 $\pm$0.2 86.7 $\pm$0.3 87.1 $\pm$0.3 89.3 $\pm$0.4 90.0 $\pm$0.3 90.2 $\pm$0.2 90.5 $\pm$0.4 89.8 $\pm$0.4 91.5 $\pm$0.3 91.9 $\pm$0.3 91.1 $\pm$0.3 91.6 $\pm$0.3 89.3 $\pm$0.5 89.9 $\pm$0.5
+C08 90.8 $\pm$1.6 93.2 $\pm$1.0 93.1 $\pm$1.2 93.2 $\pm$1.4 95.1 $\pm$0.7 95.3 $\pm$0.6 93.2 $\pm$0.9 93.4 $\pm$0.9 91.9 $\pm$1.3 91.9 $\pm$1.3 91.3 $\pm$1.1 93.5 $\pm$1.5 95.1 $\pm$0.9 92.1 $\pm$1.5 93.9 $\pm$1.1 94.4 $\pm$0.7 94.4 $\pm$0.9
+C09 82.6 $\pm$1.4 85.9 $\pm$1.2 81.3 $\pm$1.2 84.3 $\pm$1.0 86.3 $\pm$1.1 82.9 $\pm$1.3 85.6 $\pm$1.2 81.3 $\pm$2.5 84.8 $\pm$0.9 82.5 $\pm$1.1 94.0 $\pm$0.8 96.2 $\pm$0.3 83.6 $\pm$1.0 86.1 $\pm$1.3 83.2 $\pm$1.2 85.5 $\pm$1.2
+C11 78.4 $\pm$0.5 80.7 $\pm$0.4 81.8 $\pm$0.4 79.6 $\pm$0.4 81.5 $\pm$0.5 82.3 $\pm$0.4 79.2 $\pm$0.4 80.5 $\pm$0.5 79.6 $\pm$0.5 81.0 $\pm$0.3 81.4 $\pm$0.4 79.2 $\pm$0.5 81.0 $\pm$0.5 81.5 $\pm$0.3 82.8 $\pm$0.4 80.2 $\pm$0.5 81.4 $\pm$0.4
+C12 96.0 $\pm$0.2 97.7 $\pm$0.1 98.5 $\pm$0.1 93.8 $\pm$0.3 97.8 $\pm$0.1 98.6 $\pm$0.1 96.7 $\pm$0.1 79.7 $\pm$1.1 97.6 $\pm$0.1 98.5 $\pm$0.1 98.3 $\pm$0.1 99.1 $\pm$0.1 97.9 $\pm$0.1 98.7 $\pm$0.1 97.3 $\pm$0.1 98.4 $\pm$0.1
+C13 70.9 $\pm$0.8 74.1 $\pm$1.0 76.7 $\pm$0.8 70.7 $\pm$0.7 74.0 $\pm$0.8 76.7 $\pm$0.8 72.7 $\pm$1.0 75.8 $\pm$0.9 72.4 $\pm$1.2 76.4 $\pm$0.6 72.3 $\pm$0.9 75.0 $\pm$0.7 77.9 $\pm$1.0 74.4 $\pm$0.7 77.6 $\pm$1.0 74.1 $\pm$1.1 76.9 $\pm$0.9
+C14 84.0 $\pm$0.5 84.8 $\pm$0.4 82.4 $\pm$0.3 80.4 $\pm$0.3 81.3 $\pm$0.3 82.3 $\pm$0.4 83.5 $\pm$0.3 83.2 $\pm$0.4 84.3 $\pm$0.3 84.0 $\pm$0.3 85.1 $\pm$0.3 85.4 $\pm$0.3 86.2 $\pm$0.3 83.2 $\pm$0.4 84.2 $\pm$0.4
+Table 21: Regression performance per dataset (up to 10K).
+Best score before rounding is bolded.
+We report average $R^{2}$ with 95% CIs.
+Models: CARTE (CRT), CatBoost (CTB), LightGBM (LGB), RealMLP (MLP), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB).
+Tuned models are marked with a ’+’.
+ID CRT CTB CTB+ LGB LGB+ MLP MLP+ PFN RF STR TBM TBM+ XGB XGB+
+R01 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 99.9 $\pm$0.1 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 99.8 $\pm$0.1 99.8 $\pm$0.1 99.9 $\pm$0.1 100.0 $\pm$0.0
+R02 98.3 $\pm$1.0 98.3 $\pm$0.9 98.1 $\pm$1.0 98.1 $\pm$1.0 98.2 $\pm$1.0 98.2 $\pm$1.0 98.4 $\pm$0.9 98.0 $\pm$1.0 98.0 $\pm$1.1 99.0 $\pm$0.8 99.0 $\pm$0.8 97.9 $\pm$1.1 98.3 $\pm$0.9
+R03 71.8 $\pm$0.5 73.8 $\pm$0.4 74.1 $\pm$0.3 72.2 $\pm$0.4 72.4 $\pm$0.4 69.7 $\pm$0.7 70.1 $\pm$0.8 74.9 $\pm$0.4 68.4 $\pm$0.6 70.9 $\pm$0.8 75.3 $\pm$0.5 75.4 $\pm$0.5 69.1 $\pm$0.5 73.7 $\pm$0.5
+R04 86.2 $\pm$7.4 85.1 $\pm$7.1 85.7 $\pm$6.5 84.6 $\pm$6.2 88.7 $\pm$4.4 90.3 $\pm$3.2 86.1 $\pm$5.2 85.1 $\pm$6.8 81.5 $\pm$8.7 84.0 $\pm$3.1 84.1 $\pm$3.1 85.8 $\pm$7.5 87.5 $\pm$3.7
+R05 93.2 $\pm$0.8 89.5 $\pm$1.3 90.0 $\pm$1.2 87.4 $\pm$1.5 88.2 $\pm$1.4 90.9 $\pm$1.2 91.2 $\pm$1.5 92.7 $\pm$0.9 86.7 $\pm$1.3 93.2 $\pm$1.2 92.1 $\pm$1.1 92.3 $\pm$1.1 87.1 $\pm$1.4 90.1 $\pm$1.3
+R06 97.7 $\pm$0.8 98.1 $\pm$0.5 98.2 $\pm$0.5 97.7 $\pm$0.7 97.8 $\pm$0.7 97.5 $\pm$0.9 97.8 $\pm$0.9 98.5 $\pm$0.6 97.2 $\pm$1.0 98.1 $\pm$0.3 97.8 $\pm$0.8 97.9 $\pm$0.8 97.1 $\pm$0.7 97.9 $\pm$0.7
+R07 71.6 $\pm$1.3 66.9 $\pm$0.8 67.8 $\pm$0.8 62.4 $\pm$1.0 63.4 $\pm$1.0 68.9 $\pm$0.8 67.4 $\pm$1.3 62.2 $\pm$1.3 61.5 $\pm$1.2 71.1 $\pm$1.8 67.9 $\pm$0.7 69.6 $\pm$0.9 61.9 $\pm$1.5 68.8 $\pm$0.8
+R08 93.2 $\pm$0.8 93.0 $\pm$0.7 93.2 $\pm$0.7 92.9 $\pm$0.7 93.0 $\pm$0.7 92.9 $\pm$0.8 92.7 $\pm$0.8 93.9 $\pm$0.7 92.2 $\pm$0.7 92.8 $\pm$0.6 93.4 $\pm$0.7 93.3 $\pm$0.7 92.3 $\pm$0.7 93.1 $\pm$0.7
+R09 89.2 $\pm$0.5 89.2 $\pm$0.6 89.3 $\pm$0.4 88.7 $\pm$0.5 88.9 $\pm$0.5 88.6 $\pm$0.4 89.2 $\pm$0.5 89.8 $\pm$0.5 88.6 $\pm$0.7 88.8 $\pm$0.5 89.5 $\pm$0.7 89.7 $\pm$0.7 88.2 $\pm$0.7 89.3 $\pm$0.5
+R10 99.5 $\pm$0.2 99.0 $\pm$0.6 98.9 $\pm$0.6 98.5 $\pm$0.3 98.5 $\pm$0.3 99.3 $\pm$0.2 99.4 $\pm$0.2 99.1 $\pm$0.3 98.1 $\pm$0.3 99.5 $\pm$0.1 98.9 $\pm$0.7 99.1 $\pm$0.4 98.6 $\pm$0.3 99.4 $\pm$0.2
+R11 94.2 $\pm$0.9 94.0 $\pm$0.9 93.6 $\pm$0.9 93.6 $\pm$1.0 94.2 $\pm$1.0 94.4 $\pm$1.0 94.5 $\pm$0.9 92.5 $\pm$1.3 93.9 $\pm$1.1 94.5 $\pm$0.8 94.6 $\pm$0.8 93.3 $\pm$1.1 94.2 $\pm$0.8
+R12 98.5 $\pm$0.3 98.5 $\pm$0.2 98.3 $\pm$0.3 98.4 $\pm$0.2 98.4 $\pm$0.3 98.4 $\pm$0.2 98.5 $\pm$0.2 98.0 $\pm$0.3 98.3 $\pm$0.2 98.6 $\pm$0.2 98.7 $\pm$0.2 98.2 $\pm$0.3 98.5 $\pm$0.2
+R13 52.8 $\pm$2.1 57.6 $\pm$2.7 58.2 $\pm$2.6 54.7 $\pm$2.8 55.0 $\pm$2.6 55.4 $\pm$2.5 55.1 $\pm$2.5 53.8 $\pm$2.7 51.8 $\pm$3.5 53.4 $\pm$3.0 58.7 $\pm$1.9 58.4 $\pm$2.1 49.0 $\pm$3.3 57.5 $\pm$2.7
+R14 96.6 $\pm$0.2 97.3 $\pm$0.1 97.3 $\pm$0.1 97.0 $\pm$0.1 97.0 $\pm$0.1 96.8 $\pm$0.1 96.7 $\pm$0.2 96.2 $\pm$0.2 96.8 $\pm$0.1 96.4 $\pm$0.1 96.3 $\pm$0.1 96.3 $\pm$0.1 97.1 $\pm$0.1 97.1 $\pm$0.1
+R15 79.5 $\pm$0.9 79.9 $\pm$1.1 77.8 $\pm$0.9 77.7 $\pm$0.9 76.2 $\pm$1.4 77.4 $\pm$1.5 79.4 $\pm$1.0 77.3 $\pm$1.2 78.9 $\pm$1.5 80.6 $\pm$0.7 79.9 $\pm$0.7 76.8 $\pm$1.2 80.3 $\pm$1.0
+R16 98.7 $\pm$0.0 98.7 $\pm$0.0 98.7 $\pm$0.0 98.7 $\pm$0.0 98.7 $\pm$0.0 97.6 $\pm$0.1 97.8 $\pm$0.1 98.8 $\pm$0.0 97.7 $\pm$0.1 98.7 $\pm$0.0 97.9 $\pm$0.1 97.9 $\pm$0.1 97.7 $\pm$0.1 97.8 $\pm$0.1
+R17 95.3 $\pm$2.0 94.8 $\pm$2.4 95.2 $\pm$2.0 94.6 $\pm$2.3 95.2 $\pm$2.0 94.4 $\pm$2.4 95.4 $\pm$1.9 95.1 $\pm$2.1 94.2 $\pm$2.5 94.7 $\pm$2.3 96.7 $\pm$1.1 96.7 $\pm$1.1 94.5 $\pm$2.6 95.4 $\pm$1.9
+R18 98.1 $\pm$0.4 98.2 $\pm$0.4 98.2 $\pm$0.4 98.1 $\pm$0.4 98.1 $\pm$0.4 98.0 $\pm$0.5 97.7 $\pm$0.5 93.1 $\pm$1.3 97.7 $\pm$0.4 97.4 $\pm$0.6 98.2 $\pm$0.3 98.1 $\pm$0.4 97.5 $\pm$0.8 98.3 $\pm$0.4
+R19 85.5 $\pm$0.8 85.8 $\pm$0.9 85.2 $\pm$0.9 85.2 $\pm$0.9 82.7 $\pm$0.8 83.3 $\pm$1.3 84.0 $\pm$1.1 85.3 $\pm$0.9 83.8 $\pm$1.1 85.0 $\pm$1.0 85.1 $\pm$0.9 83.9 $\pm$1.0 82.4 $\pm$2.0
+R20 96.4 $\pm$0.5 96.0 $\pm$0.6 96.2 $\pm$0.5 95.7 $\pm$0.5 95.8 $\pm$0.5 95.3 $\pm$0.8 96.0 $\pm$0.5 93.9 $\pm$1.0 95.7 $\pm$0.6 95.6 $\pm$0.9 95.9 $\pm$0.7 95.8 $\pm$0.7 95.3 $\pm$0.9 96.2 $\pm$0.5
+R21 92.3 $\pm$1.2 92.5 $\pm$1.1 92.6 $\pm$1.1 92.3 $\pm$1.2 92.3 $\pm$1.2 92.0 $\pm$1.1 91.7 $\pm$1.3 93.3 $\pm$1.1 91.8 $\pm$1.1 92.3 $\pm$1.0 91.6 $\pm$1.4 92.1 $\pm$1.2 91.6 $\pm$1.0 92.6 $\pm$1.0
+R22 45.2 $\pm$5.5 45.2 $\pm$5.9 42.4 $\pm$5.6 44.7 $\pm$5.5 43.6 $\pm$6.4 44.4 $\pm$6.3 43.8 $\pm$5.5 39.1 $\pm$6.4 39.7 $\pm$5.5 47.9 $\pm$5.5 48.1 $\pm$5.4 36.1 $\pm$7.1 46.8 $\pm$5.8
+R23 85.0 $\pm$1.8 85.4 $\pm$1.2 85.5 $\pm$1.1 84.3 $\pm$1.5 85.5 $\pm$1.2 83.3 $\pm$1.8 85.6 $\pm$1.5 84.8 $\pm$1.6 81.8 $\pm$1.5 83.0 $\pm$1.6 86.4 $\pm$0.9 86.6 $\pm$0.8 83.3 $\pm$1.2 85.1 $\pm$1.2
+R24 86.0 $\pm$0.6 84.0 $\pm$0.8 85.8 $\pm$0.7 78.8 $\pm$0.9 79.5 $\pm$0.9 84.6 $\pm$0.8 82.8 $\pm$0.9 70.3 $\pm$1.6 79.7 $\pm$1.0 81.8 $\pm$1.8 84.9 $\pm$0.9 85.5 $\pm$0.8 82.9 $\pm$0.7 85.9 $\pm$0.9
+R25 94.3 $\pm$0.6 95.4 $\pm$0.5 95.3 $\pm$0.5 95.2 $\pm$0.6 95.2 $\pm$0.5 93.2 $\pm$0.6 94.1 $\pm$0.7 86.7 $\pm$1.0 94.8 $\pm$0.5 94.8 $\pm$0.5 94.3 $\pm$0.4 94.5 $\pm$0.4 94.6 $\pm$0.6 95.4 $\pm$0.5
+R26 99.8 $\pm$0.1 99.4 $\pm$0.1 99.4 $\pm$0.1 99.2 $\pm$0.2 99.3 $\pm$0.2 99.9 $\pm$0.1 99.9 $\pm$0.0 99.8 $\pm$0.1 99.0 $\pm$0.2 99.6 $\pm$0.1 98.1 $\pm$0.7 98.7 $\pm$0.8 99.1 $\pm$0.2 99.5 $\pm$0.1
+R27 81.9 $\pm$1.6 85.2 $\pm$1.3 85.3 $\pm$1.4 84.4 $\pm$1.4 84.8 $\pm$1.3 83.3 $\pm$1.5 76.7 $\pm$11.2 82.2 $\pm$1.6 84.8 $\pm$1.3 82.0 $\pm$1.6 83.8 $\pm$1.5 84.0 $\pm$1.7 83.2 $\pm$1.2 85.5 $\pm$1.2
+R28 52.5 $\pm$2.6 53.4 $\pm$2.7 53.3 $\pm$2.8 50.0 $\pm$3.0 51.0 $\pm$2.9 52.3 $\pm$3.2 52.6 $\pm$3.0 61.7 $\pm$2.1 46.0 $\pm$2.7 51.5 $\pm$2.9 54.7 $\pm$2.7 55.6 $\pm$2.6 45.3 $\pm$2.3 53.5 $\pm$2.7
+R29 94.4 $\pm$0.7 94.4 $\pm$0.7 94.8 $\pm$0.7 94.8 $\pm$0.8 94.3 $\pm$0.7 94.4 $\pm$0.8 95.7 $\pm$0.6 93.0 $\pm$1.0 94.3 $\pm$0.9 94.9 $\pm$0.8 95.0 $\pm$0.7 93.4 $\pm$0.8 94.5 $\pm$0.8
+R30 23.8 $\pm$4.2 22.7 $\pm$4.1 24.6 $\pm$3.9 22.3 $\pm$3.7 23.9 $\pm$4.2 15.7 $\pm$3.9 18.8 $\pm$5.4 20.9 $\pm$4.8 23.5 $\pm$3.5 15.7 $\pm$5.0 22.9 $\pm$5.0 23.4 $\pm$4.8 17.5 $\pm$4.2 25.5 $\pm$3.7
+R31 92.1 $\pm$0.3 92.1 $\pm$0.4 92.0 $\pm$0.4 91.6 $\pm$0.4 91.6 $\pm$0.4 91.5 $\pm$0.3 91.5 $\pm$0.4 93.2 $\pm$0.3 89.9 $\pm$0.5 91.7 $\pm$0.4 92.4 $\pm$0.5 92.6 $\pm$0.5 90.3 $\pm$0.5 92.0 $\pm$0.4
+R32 28.8 $\pm$6.3 28.2 $\pm$5.8 25.0 $\pm$6.4 27.3 $\pm$5.7 19.8 $\pm$8.8 19.8 $\pm$7.1 26.2 $\pm$5.4 28.6 $\pm$5.9 19.4 $\pm$5.6 25.8 $\pm$5.3 27.3 $\pm$6.0 22.7 $\pm$7.3 31.6 $\pm$6.2
+R33 46.6 $\pm$1.5 47.4 $\pm$1.7 47.8 $\pm$1.7 44.8 $\pm$1.8 45.2 $\pm$1.6 40.4 $\pm$2.1 44.6 $\pm$1.6 44.9 $\pm$1.6 40.2 $\pm$1.6 46.0 $\pm$1.8 48.4 $\pm$1.1 48.4 $\pm$1.2 40.3 $\pm$2.1 47.9 $\pm$1.6
+R34 90.9 $\pm$2.5 91.4 $\pm$2.1 88.9 $\pm$2.9 89.9 $\pm$2.4 88.6 $\pm$3.0 90.7 $\pm$1.4 92.3 $\pm$1.8 88.7 $\pm$3.1 89.1 $\pm$2.9 91.1 $\pm$1.7 90.9 $\pm$1.8 88.9 $\pm$3.6 91.6 $\pm$2.1
+R35 85.9 $\pm$0.6 84.4 $\pm$0.5 84.5 $\pm$0.5 83.7 $\pm$0.6 83.7 $\pm$0.6 83.1 $\pm$0.9 84.2 $\pm$0.8 85.2 $\pm$0.7 81.2 $\pm$0.4 85.7 $\pm$0.8 84.5 $\pm$0.6 84.5 $\pm$0.5 81.6 $\pm$0.8 84.4 $\pm$0.3
+R36 96.3 $\pm$0.9 95.2 $\pm$0.9 95.5 $\pm$0.9 94.7 $\pm$1.0 94.9 $\pm$1.0 95.4 $\pm$1.1 95.7 $\pm$0.9 91.2 $\pm$1.1 94.3 $\pm$1.0 95.7 $\pm$0.8 96.2 $\pm$0.6 96.4 $\pm$0.6 94.4 $\pm$0.9 95.5 $\pm$0.9
+Table 21: Regression performance per dataset (up to 10K).
+Best score before rounding is bolded.
+We report average $R^{2}$ with 95% CIs.
+Models: CARTE (CRT), CatBoost (CTB), LightGBM (LGB), RealMLP (MLP), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB).
+Tuned models are marked with a ’+’.
+Table 22: Regression performance per dataset (above 10K).
+Best score before rounding is bolded.
+We report average $R^{2}$ with 95% CIs.
+Models: CARTE (CRT), CatBoost (CTB), LightGBM (LGB), RealMLP (MLP), TabM (TBM), TabPFNv2 (PFN), RandomForest (RF), TabSTAR (STR), XGBoost (XGB).
+Unlimited models are marked with a ’!’.
+ID CRT CTB CTB! LGB LGB! MLP MLP! PFN STR STR! TBM TBM! XGB XGB!
+R01 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0 99.8 $\pm$0.1 100.0 $\pm$0.0 100.0 $\pm$0.0 100.0 $\pm$0.0
+R02 98.3 $\pm$0.9 98.8 $\pm$0.9 98.1 $\pm$1.0 98.8 $\pm$1.0 98.2 $\pm$1.0 98.8 $\pm$1.0 98.4 $\pm$0.9 98.0 $\pm$1.1 98.4 $\pm$1.0 99.0 $\pm$0.8 98.5 $\pm$0.9 98.3 $\pm$0.9 98.8 $\pm$0.9
+R03 71.8 $\pm$0.5 74.1 $\pm$0.3 75.1 $\pm$0.4 72.4 $\pm$0.4 72.7 $\pm$0.5 70.1 $\pm$0.8 71.0 $\pm$0.6 74.9 $\pm$0.4 70.9 $\pm$0.8 72.3 $\pm$0.6 75.4 $\pm$0.5 76.4 $\pm$0.5 73.7 $\pm$0.5 74.6 $\pm$0.6
+R04 85.1 $\pm$7.1 91.3 $\pm$0.8 84.6 $\pm$6.2 91.0 $\pm$1.3 90.3 $\pm$3.2 91.4 $\pm$0.8 86.1 $\pm$5.2 81.5 $\pm$8.7 91.0 $\pm$0.7 84.1 $\pm$3.1 91.0 $\pm$0.7 87.5 $\pm$3.7 91.3 $\pm$0.6
+R07 71.6 $\pm$1.3 67.8 $\pm$0.8 85.2 $\pm$0.3 63.4 $\pm$1.0 70.1 $\pm$0.5 67.4 $\pm$1.3 85.0 $\pm$0.3 62.2 $\pm$1.3 71.1 $\pm$1.8 79.9 $\pm$1.8 69.6 $\pm$0.9 84.0 $\pm$1.3 68.8 $\pm$0.8 86.1 $\pm$0.7
+R08 93.2 $\pm$0.8 93.2 $\pm$0.7 93.8 $\pm$0.6 93.0 $\pm$0.7 93.6 $\pm$0.5 92.7 $\pm$0.8 93.5 $\pm$0.7 93.9 $\pm$0.7 92.8 $\pm$0.6 93.3 $\pm$0.6 93.3 $\pm$0.7 93.9 $\pm$0.6 93.1 $\pm$0.7 93.9 $\pm$0.5
+R09 89.2 $\pm$0.5 89.3 $\pm$0.4 89.5 $\pm$0.4 88.9 $\pm$0.5 89.2 $\pm$0.5 89.2 $\pm$0.5 89.3 $\pm$0.5 89.8 $\pm$0.5 88.8 $\pm$0.5 89.2 $\pm$0.6 89.7 $\pm$0.7 89.3 $\pm$0.4 89.3 $\pm$0.5 89.5 $\pm$0.4
+R12 98.5 $\pm$0.2 98.9 $\pm$0.1 98.4 $\pm$0.2 98.9 $\pm$0.2 98.4 $\pm$0.2 99.0 $\pm$0.1 98.5 $\pm$0.2 98.3 $\pm$0.2 98.4 $\pm$0.2 98.7 $\pm$0.2 99.0 $\pm$0.2 98.5 $\pm$0.2 98.9 $\pm$0.2
+R14 96.6 $\pm$0.2 97.3 $\pm$0.1 97.8 $\pm$0.1 97.0 $\pm$0.1 97.5 $\pm$0.1 96.7 $\pm$0.2 97.4 $\pm$0.1 96.2 $\pm$0.2 96.4 $\pm$0.1 96.8 $\pm$0.2 96.3 $\pm$0.1 97.2 $\pm$0.1 97.1 $\pm$0.1 97.5 $\pm$0.1
+R15 79.9 $\pm$1.1 86.5 $\pm$0.7 77.7 $\pm$0.9 82.0 $\pm$0.6 77.4 $\pm$1.5 85.7 $\pm$0.6 79.4 $\pm$1.0 78.9 $\pm$1.5 83.9 $\pm$1.4 79.9 $\pm$0.7 87.2 $\pm$0.8 80.3 $\pm$1.0 87.0 $\pm$0.7
+R16 98.7 $\pm$0.0 98.7 $\pm$0.0 98.8 $\pm$0.0 98.7 $\pm$0.0 98.8 $\pm$0.0 97.8 $\pm$0.1 98.0 $\pm$0.1 98.8 $\pm$0.0 98.7 $\pm$0.0 98.8 $\pm$0.0 97.9 $\pm$0.1 98.1 $\pm$0.1 97.8 $\pm$0.1 98.0 $\pm$0.1
+R17 95.3 $\pm$2.0 95.2 $\pm$2.0 97.6 $\pm$0.6 95.2 $\pm$2.0 97.6 $\pm$0.7 95.4 $\pm$1.9 97.5 $\pm$0.6 95.1 $\pm$2.1 94.7 $\pm$2.3 97.6 $\pm$0.6 96.7 $\pm$1.1 97.2 $\pm$0.7 95.4 $\pm$1.9 97.7 $\pm$0.6
+R18 98.1 $\pm$0.4 98.2 $\pm$0.4 99.0 $\pm$0.2 98.1 $\pm$0.4 98.8 $\pm$0.3 97.7 $\pm$0.5 98.4 $\pm$0.4 93.1 $\pm$1.3 97.4 $\pm$0.6 97.7 $\pm$0.6 98.1 $\pm$0.4 98.7 $\pm$0.3 98.3 $\pm$0.4 98.9 $\pm$0.3
+R20 96.4 $\pm$0.5 96.2 $\pm$0.5 96.3 $\pm$0.4 95.8 $\pm$0.5 95.8 $\pm$0.5 96.0 $\pm$0.5 95.6 $\pm$0.5 93.9 $\pm$1.0 95.6 $\pm$0.9 95.3 $\pm$1.0 95.8 $\pm$0.7 95.9 $\pm$0.6 96.2 $\pm$0.5 96.3 $\pm$0.5
+R23 85.0 $\pm$1.8 85.5 $\pm$1.1 86.3 $\pm$0.7 85.5 $\pm$1.2 85.8 $\pm$0.7 85.6 $\pm$1.5 86.6 $\pm$0.8 84.8 $\pm$1.6 83.0 $\pm$1.6 85.2 $\pm$1.7 86.6 $\pm$0.8 88.3 $\pm$0.4 85.1 $\pm$1.2 86.1 $\pm$0.8
+R24 86.0 $\pm$0.6 85.8 $\pm$0.7 97.1 $\pm$0.3 79.5 $\pm$0.9 85.4 $\pm$0.6 82.8 $\pm$0.9 96.5 $\pm$0.4 70.3 $\pm$1.6 81.8 $\pm$1.8 95.5 $\pm$0.7 85.5 $\pm$0.8 97.8 $\pm$0.3 85.9 $\pm$0.9 97.3 $\pm$0.3
+R25 94.3 $\pm$0.6 95.3 $\pm$0.5 95.9 $\pm$0.4 95.2 $\pm$0.5 95.7 $\pm$0.5 94.1 $\pm$0.7 95.0 $\pm$0.5 86.7 $\pm$1.0 94.8 $\pm$0.5 95.1 $\pm$0.5 94.5 $\pm$0.4 95.9 $\pm$0.4 95.4 $\pm$0.5 95.8 $\pm$0.5
+R31 92.1 $\pm$0.3 92.0 $\pm$0.4 93.0 $\pm$0.3 91.6 $\pm$0.4 92.5 $\pm$0.3 91.5 $\pm$0.4 92.5 $\pm$0.4 93.2 $\pm$0.3 91.7 $\pm$0.4 92.7 $\pm$0.4 92.6 $\pm$0.5 93.6 $\pm$0.3 92.0 $\pm$0.4 92.9 $\pm$0.3
+R33 46.6 $\pm$1.5 47.8 $\pm$1.7 55.9 $\pm$1.1 45.2 $\pm$1.6 49.8 $\pm$1.3 44.6 $\pm$1.6 54.5 $\pm$1.3 44.9 $\pm$1.6 46.0 $\pm$1.8 57.2 $\pm$1.3 48.4 $\pm$1.2 57.8 $\pm$1.2 47.9 $\pm$1.6 55.7 $\pm$1.3
+R36 96.3 $\pm$0.9 95.5 $\pm$0.9 95.5 $\pm$0.9 94.9 $\pm$1.0 94.9 $\pm$1.0 95.7 $\pm$0.9 95.8 $\pm$0.9 91.2 $\pm$1.1 95.7 $\pm$0.8 95.5 $\pm$0.5 96.4 $\pm$0.6 96.1 $\pm$0.9 95.5 $\pm$0.9 95.5 $\pm$0.9
+F.3 Head-to-head comparisons
 We compare the performance of TabSTAR against each of the models in head-to-head comparisons.
@@
 Table 23 shows the performance of TabSTAR against all models competing up to 10K examples, for both regression and classification, with 95% CIs over the win rate.
-Table 22 does the same for TabSTAR-Unlimit.
+Table 24 does the same for TabSTAR-Unlimited.
 Table 23: Win rates of TabSTAR (up to 10K) against baselines. Win rate with 95% CI.
-CARTE CatB CatB-T RF TabPFN XGB XGB-T
-Classification 93.3±5.2 73.6±7.3 67.9±7.8 88.6±5.3 57.3±9.3 89.3±5.1 71.4±7.5
-Regression 35.7±5.9 33.6±4.9 32.2±4.8 66.9±4.9 40.9±5.2 69.4±4.8 30.3±4.8
-Table 22: Win rates of TabSTAR-U (above 10K) against baselines. Win rate with 95% CI.
-CARTE CatB-T CatB-T-U TabPFN TabSTAR XGB-T XGB-T-U
-Classification 98.6±2.8 86.0±6.8 67.0±9.3 75.7±10.1 94.0±4.7 85.0±7.0 71.0±8.9
-Regression 63.5±7.5 53.0±6.9 22.0±5.8 61.1±7.0 82.0±5.3 59.5±6.8 28.5±6.3
-F.3 Running Times and Compute Information
-In this section we present the running times of the different models (see Table 23).
-We focus on the condition with up to 10,000 examples, excluding CatBoost-Tuned and XGBoost-Tuned as they run with a 4 hours budget using 8 CPUs from AMD EPYC 7742 64-Core Processor .
-Compute Information
-TabSTAR runs on NVIDIA RTX A40 with 48GB memory.
-TabPFN-v2 runs directly on their API client.
-CatBoost, XGboost and Random Forest run using Intel(R) Core(TM) i9-10920X CPU @ 3.50GHz CPU cores.
-CARTE use inferior GPUs of type NVIDIA GeForce GTX 1080 Ti with 11GB memory, to allow for higher parallelization of their incredibly slow runs.
+Model Classification Regression
+CARTE 93.3 ± 5.2 35.7 ± 5.9
+CatBoost 80.7 ± 6.6 33.6 ± 4.9
+CatBoost-Tuned 75.0 ± 7.2 32.2 ± 4.8
+LightGBM 92.1 ± 4.5 49.7 ± 5.2
+LightGBM-Tuned 90.7 ± 4.8 45.3 ± 5.1
+RandomForest 95.7 ± 3.4 66.9 ± 4.9
+RealMLP 82.1 ± 6.4 51.4 ± 5.2
+RealMLP-Tuned 83.6 ± 6.2 49.4 ± 5.2
+TabDPT 73.6 ± 7.3 66.7 ± 4.9
+TabICL 70.0 ± 7.6
+TabM 64.3 ± 8.0 39.2 ± 5.0
+TabM-Tuned 63.3 ± 8.0 37.4 ± 5.0
+TabPFN-v2 66.4 ± 8.9 40.9 ± 5.2
+XGBoost 96.4 ± 3.1 69.4 ± 4.8
+XGBoost-Tuned 78.6 ± 6.8 30.3 ± 4.8
+Table 24: Win rates of TabSTAR-Unlimited (above 10K) against baselines. Win rate with 95% CI.
+Model Classification Regression
+CARTE 98.6 ± 2.8 63.5 ± 7.5
+CatBoost-Tuned 96.0 ± 3.9 53.0 ± 6.9
+CatBoost-Tuned-Unlimited 77.0 ± 8.3 22.0 ± 5.8
+LightGBM-Tuned 98.0 ± 2.8 69.5 ± 6.4
+LightGBM-Tuned-Unlimited 92.0 ± 5.3 44.0 ± 6.9
+RealMLP-Tuned 97.0 ± 3.4 72.0 ± 6.2
+RealMLP-Tuned-Unlimited 86.0 ± 6.8 38.9 ± 7.0
+TabDPT 85.0 ± 7.0
+TabICL 85.0 ± 7.0
+TabICL-Unlimited 82.0 ± 7.6
+TabM-Tuned 76.8 ± 8.4 65.2 ± 6.7
+TabM-Tuned-Unlimited 53.0 ± 9.8 22.5 ± 5.8
+TabPFN-v2 90.0 ± 7.1 61.1 ± 7.0
+TabSTAR 94.0 ± 4.7 82.0 ± 5.3
+XGBoost-Tuned 95.0 ± 4.3 59.5 ± 6.8
+XGBoost-Tuned-Unlimited 81.0 ± 7.7 28.5 ± 6.3
+F.4 Running Times and Compute Information
+Hardware
+All baselines are evaluated using a single NVIDIA A100-SXM4 GPU with 40GB memory, and 8 CPU cores of type AMD EPYC 7742 64-Core Processor.
+The only exclusions are TabPFN-v2, which runs on its API version, and CARTE, for which we employ several NVIDIA GeForce GTX 1080 Ti GPUs with 11GB memory, as their high latency and lack of default hyperparameters forced us to rely on more accessible hardware.
 Running Times
-TabSTAR’s average running time for a downstream task ranges from 30 seconds (C04) to 7,692 seconds (R24).
-We observe that these running times are significantly lower than 14,400 seconds (4 hours), used for CatBoost-Tuned and XGBoost-Tuned, although a single run of any tree model without 5-fold cross-validation is faster.
-The running times of TabPFN-v2 are considerably better.
-However, since they use ICL, they are limited to run only up to 10,000 examples and their inference time is just as costly as their fitting time.
-For CARTE, while their runs use inferior hardware and limited parallelization, the reported times are for a single run.
-In practice, they run 6 times per dataset split because of their lack of default hyperparameters, making it a poor choice for scale.
+Table 25 presents the average running times of the different models per dataset, with up to 10,000 examples, excluding the tuned variants, TabPFN-v2 and CARTE.
+TabSTAR’s average running time for a downstream task ranges from 59s (C04) to 9,332s (R24).
+Although higher than a single run of other baselines, these times remain far below the 14,400s (4 hours) typically required for their tuned counterparts.
+TabICL achieves considerably faster running times; however, their ICL-based approach limits them to smaller datasets, and inference is as costly as training (see §5).
+TabPFN-v2, though unreported, is subject to the same limitation.
 Table 25: Average model training running time in seconds, per dataset, for up to 10K examples.
-ID CARTE CatBoost RF TabPFN-v2 TabSTAR XGBoost
-C01 13,341 44 19 42 302 31
-C02 420 346 1,145 491
-C03 7,297 38 19 61 324 27
-C04 11 6 9 30 9
-C05 1,227 10 8 14 68 9
-C06 3,068 68 31 112 1,126 63
-C07 6,559 61 34 83 685 52
-C08 4,631 40 41 49 681 49
-C09 20 21 134 544 31
-C10 22 22 30 519 23
-C11 2,428 19 12 33 252 18
-C12 4,411 169 16 765 39
-C13 1,774 31 27 58 929 37
-C14 500 140 5,254 308
-R01 9,439 139 446 1,987 94
-R02 9 16 54 188 6
-R03 2,174 30 103 357 553 24
-R04 10 10 84 115 7
-R05 10,639 28 26 62 158 15
-R06 3,185 22 43 70 307 13
-R07 4,637 22 109 112 352 28
-R08 4,252 57 366 396 1,136 58
-R09 8,239 7 22 55 198 5
-R10 1,505 28 37 33 168 21
-R11 11 18 36 99 8
-R12 13 22 78 199 9
-R13 2,299 24 54 38 639 21
-R14 6,396 28 30 167 772 15
-R15 25 51 66 223 22
-R16 13,380 34 118 156 798 35
-R17 3,247 94 478 204 270 143
-R18 9,663 34 631 309 341 32
-R19 11 14 66 101 6
-R20 6,173 22 144 154 637 20
-R21 2,215 14 33 25 177 14
-R22 16 21 17 149 15
-R23 14,401 15 51 58 167 10
-R24 25,903 90 170 314 7,692 106
-R25 2,404 50 943 261 786 54
-R26 636 10 11 13 71 8
-R27 2,615 31 59 40 304 22
-R28 12,265 45 163 128 389 46
-R29 11 10 32 138 5
-R30 2,477 9 18 24 45 7
-R31 5,102 41 187 240 451 39
-R32 8 11 13 32 7
-R33 4,512 28 100 136 539 30
-R34 26 24 18 61 12
-R35 4,816 30 55 48 396 28
-R36 15,389 18 81 79 949 23
+ID CatBoost LightGBM RandomForest RealMLP TabICL TabSTAR XGBoost
+C01 116 35 48 148 37 397 36
+C02 251 57 75 179 58 1,135 58
+C03 106 51 61 147 49 553 48
+C04 29 17 19 25 20 59 18
+C05 26 13 16 63 13 132 13
+C06 170 79 83 126 77 938 81
+C07 163 78 98 186 80 1,044 78
+C08 89 72 86 180 74 1,057 69
+C09 33 25 33 128 29 414 25
+C10 49 35 43 126 36 572 35
+C11 43 22 29 121 20 521 20
+C12 614 39 37 145 28 729 36
+C13 70 55 68 163 55 551 50
+C14 1,938 307 333 418 311 4,068 329
+R01 488 161 1,204 368 2,908 161
+R02 28 10 45 128 283 10
+R03 111 62 282 171 592 61
+R04 21 9 17 138 240 7
+R05 65 20 68 90 366 23
+R06 60 24 119 129 501 23
+R07 66 39 285 145 673 44
+R08 213 105 951 222 1,448 112
+R09 12 8 56 115 604 7
+R10 102 45 104 72 255 47
+R11 26 10 45 98 197 10
+R12 34 10 42 133 335 11
+R13 78 44 145 94 484 46
+R14 67 24 91 159 1,296 27
+R15 38 18 102 132 390 17
+R16 108 76 316 193 898 80
+R17 62 54 1,110 160 261 53
+R18 124 74 1,705 201 553 68
+R19 22 13 38 107 198 12
+R20 70 52 401 176 705 49
+R21 53 29 86 62 222 30
+R22 57 31 62 52 171 33
+R23 41 19 137 140 281 18
+R24 461 396 650 520 9,332 392
+R25 191 104 2,583 223 750 110
+R26 29 12 22 30 205 11
+R27 103 49 164 92 334 49
+R28 175 92 460 158 647 97
+R29 34 8 21 100 308 7
+R30 26 14 47 56 114 14
+R31 148 78 505 200 797 77
+R32 20 11 23 31 82 11
+R33 89 55 271 161 608 47
+R34 66 22 66 50 122 24
+R35 99 50 146 101 411 50
+R36 58 38 223 146 823 39
+F.5 TabSTAR’s Cost Analysis by Number of Features
+While §5 reports average computational costs, these values vary considerably across datasets, depending on factors such as the number of examples, the feature count, and the amount of textual fields.
+In this subsection, we examine how TabSTAR’s computational cost scales with the number of features.
+Table 26 summarizes both latency and memory consumption of TabSTAR, aggregated over a group of up to 10 features, a group of 30-50 features and a group of more than 100 features.
+These datasets were randomly sampled from the pretraining corpus; consequently, the reported times may differ from those observed on the benchmark datasets.
+We observe that latency can increase by a factor of 5, and GPU memory consumption by a factor of 4.
+This behavior stems from the transformer’s architecture, whose computational complexity is quadratic in the sequence length.
+This raises concerns for training TabSTAR over datasets with hundreds of features, and calls for improvements to expand to datasets of thousands of features.
+Additionally, TabSTAR’s memory consumption in classification tasks is affected by the number of classes due to its target-aware tokens.
+We assume that for TabSTAR, the effective number of features equals the actual number of features plus the number of classes.
+For highly multiclass datasets, this can become a limiting factor.
+Table 26: Cost Analysis for TabSTAR per number of features. Median training and inference times and peak memory usage on GPU and CPU, aggregated across 45 datasets.
+Train Inference
+Feature Group Time (s) GPU (GB) CPU (GB) Time (s) GPU (GB) CPU (GB)
+Up to 10 85.6 1.9 4.4 0.5 1.6 4.4
+30-50 228.8 3.1 4.5 1.0 1.7 4.4
+100+ 417.9 7.8 4.5 2.5 2.3 4.4
+F.6 Number of Trials for Tuned models
+The tuned models are optimized with a budget of 4 hours using 8 CPU cores.
+Each trial is executed on a single core, ensuring that at least one trial could be completed within the allocated time.
+Table 27 presents the number of hyperparameter trials per dataset.
+Table 27: Average number of hyperparameter trials per dataset. Tuned models: CatBoost (CTB), LightGBM (LGB), RealMLP (MLP), XGBoost (XGB). Unlimited models are marked with a ’!’.
+Dataset CTB CTB! LGB LGB! MLP MLP! XGB XGB!
+C01 27.1 24.4 770.2 812.0 56.0 26.7 97.7 90.0
+C02 18.1 9.4 417.5 51.3 31.9 8.0 47.2 19.6
+C03 18.0 15.8 312.0 338.1 37.4 25.2 42.4 39.5
+C04 105.8 11283.8 642.1 458.5
+C05 122.1 3373.3 98.3 228.3
+C06 18.1 406.9 67.7 44.6
+C07 17.3 18.5 335.6 376.6 31.8 23.7 53.4 37.7
+C08 101.1 67.8 2911.4 2394.6 43.1 37.3 508.3 566.1
+C09 169.0 81.3 3969.5 475.7 50.5 9.2 883.9 207.2
+C10 73.7 1598.1 55.4 165.2
+C11 58.3 48.5 812.4 863.4 44.8 26.0 124.1 96.3
+C12 9.9 8.5 207.2 34.4 35.3 8.2 41.8 17.3
+C13 85.8 27.4 1848.3 329.6 40.8 8.7 234.6 105.7
+C14 8.6 8.4 50.9 36.3 14.9 9.7 13.7 11.6
+R01 14.2 12.2 278.5 169.6 11.3 32.8 15.4
+R02 138.9 118.6 4463.0 5214.6 48.0 30.7 742.4 1054.5
+R03 55.5 64.2 1033.6 1265.9 35.8 30.0 94.3 97.8
+R04 123.8 31.6 4309.9 714.9 44.1 8.7 1324.0 364.8
+R05 78.8 4428.5 87.9 307.0
+R06 88.8 2305.3 47.5 232.2
+R07 109.9 58.8 1370.0 344.1 42.0 8.5 145.6 72.5
+R08 30.5 30.1 568.6 122.9 21.4 10.4 65.7 47.1
+R09 194.6 208.7 6391.0 5975.1 50.4 31.2 589.0 680.6
+R10 53.7 2806.9 135.9 241.8
+R11 142.9 4669.1 58.7 662.6
+R12 126.8 45.6 2990.9 814.7 46.9 8.6 385.8 178.0
+R13 80.9 2814.6 91.2 204.8
+R14 59.9 67.6 1725.3 1542.6 39.5 15.2 200.6 177.0
+R15 107.4 69.2 2896.1 1177.4 52.5 14.0 337.5 242.7
+R16 65.8 46.3 1446.3 663.7 34.9 10.9 160.9 104.9
+R17 79.7 44.6 1517.3 316.3 45.3 9.1 254.0 104.4
+R18 39.1 43.8 1572.4 1041.4 29.4 16.6 258.8 144.5
+R19 117.5 4046.2 63.1 1063.2
+R20 62.6 89.1 1611.5 2176.4 36.4 30.1 250.5 258.8
+R21 101.6 3284.9 122.3 159.4
+R22 67.8 3584.3 259.9 127.2
+R23 91.8 105.4 3179.2 2773.0 53.4 27.6 420.1 314.5
+R24 52.4 38.1 867.2 417.7 35.2 10.4 104.7 60.3
+R25 40.4 33.5 806.9 357.6 23.6 10.8 122.4 72.0
+R26 166.5 12567.7 279.4 625.0
+R27 53.9 2034.2 85.3 155.5
+R28 40.9 670.8 37.5 59.3
+R29 135.3 7615.2 63.8 1024.3
+R30 194.3 5275.3 114.8 431.7
+R31 36.6 39.5 774.5 504.3 26.4 12.6 90.3 54.1
+R32 143.2 6605.7 191.7 337.3
+R33 69.3 30.8 1053.9 212.0 36.7 8.3 119.8 51.1
+R34 69.1 3962.2 148.9 314.9
+R35 65.8 3351.6 76.3 139.9
+R36 108.0 120.2 2633.9 4634.0 49.9 48.0 289.9 228.8
 Appendix G Extended Analysis
@@
 G.1 Evaluation Datasets for Analysis
-All the experiments are evaluated over 20 datasets from the benchmark in Appendix D.
-Each experiment reports the performance with AUROC for classification and $R^{2}$ for regression.
-For conciseness, each tables reports both regression and classification tasks, distinguishable by their ID.
+All experiments are conducted on 20 datasets from the benchmark described in Appendix D.
+Each experiment reports performance using AUROC for classification and $R^{2}$ for regression.
+Each tables reports both regression and classification tasks, distinguishable by their ID.
+Furthermore, each experiment compares its variants to other models, excluding TabPFN-v2 and CARTE, which could not be executed on all datasets.
+Although this comparison covers only a subset of the benchmark and should not be interpreted as conclusive, the results nonetheless offer valuable reference points.
 G.2 The Role of the Encoder Unfreezing (Q1)
 Table 28 shows the results for each variant of the experiment presented in §6.
-It is evident that unfreezing the textual encoder yields a significant performance across datasets.
+It is evident that unfreezing the textual encoder yields a significant performance boost across datasets, as seen in Figure 6: The frozen TabSTAR variant is much worse than any other baseline!
 Furthermore, while finetuning a single layer gives a significant boost, it underperforms compared to 6 unfrozen layers.
-G.3 The Effect of Pretraining (Q2)
-We pretrain three TabSTAR variants on nested dataset subsets of size 16, 64, and 256.
-The 64-dataset variant contains the original 16 plus 48 new datasets, and the 256-dataset variant builds on those 64 by adding another 192.
-This cumulative design minimizes variance between variants so that performance differences reflect only the effect of increasing data volume.
-While LoRA [41] is a very efficient technique, its performance was much worse for a randomly initialized model.
-Therefore, we perform full finetuning of the non-pretrained model, as explained in Appendix B.2.
-Table 25 shows the dataset level results.
-It is evident that for most of the datasets, improvement is observed when scaling, with the 256 datasets variant winning for almost all datasets.
-G.4 Numerical Verbalization (Q3)
-We show the full results for the experiment in §6, with Table 26 illustrating the verbalizations in each variant.
-Note that we do not include an exact value verbalization, since it would increase the number of unique text inputs and place extra memory demands.
-The two variants which integrate numerical information into the verbalization dominate the experiment, although the improvement seems to be marginal for some datasets.
-Interestingly, some datasets significantly underperform, with the R27 dataset completely failing the task.
-The addition of the quantile information on top of the bin seems to have limited impact, although marginally winning on the average performance.
+Figure 6: Comparison of normalized scores with 95% CIs for the encoder unfreezing experiment, comparing TabSTAR variants against baseline models in classification tasks.
 Table 28: Downstream performance for Q1: The Role of the Encoder Unfreezing.
@@
 R34 84.3±2.8 87.4±2.9 88.4±3.4 87.9±3.5 88.7±2.7
+G.3 The Effect of Pretraining (Q2)
+We pretrain three TabSTAR variants on nested dataset subsets of size 16, 64, and 256.
+The 64-dataset variant contains the original 16 plus 48 new datasets, and the 256-dataset variant builds on those 64 by adding another 192.
+This cumulative design minimizes variance between variants so that performance differences reflect only the effect of increasing data volume.
+While LoRA [42] is a powerful technique, it can’t be applied to a randomly initialized model. Therefore, we perform full finetuning of the non-pretrained model, as explained in Appendix B.2.
+Table 29 show the normalized results for classification and regression, and Table 30 shows the dataset level results.
+It is evident that for most of the datasets, improvement is observed when scaling, with the 256 datasets variant winning for almost all datasets.
+Figure 7 highlights this effect, showing that adding more datasets enables TabSTAR to outperform the baselines.
+Table 29: Normalized score with 95% CIs by the number of datasets used during TabSTAR pretraining.
+Pretraining Datasets 0 16 64 256
+Classification 0.352 ± 0.086 0.450 ± 0.084 0.558 ± 0.086 0.786 ± 0.076
+Regression 0.338 ± 0.073 0.395 ± 0.068 0.642 ± 0.066 0.811 ± 0.055
+Figure 7: Comparison of normalized scores with 95% CIs for the pretraining effect experiment, comparing TabSTAR variants against baseline models in classification tasks.
 Table 30: Downstream performance for Q2: The Effect of Pretraining.
@@
 R34 83.8±4.3 86.0±3.3 88.0±3.4 87.9±3.5
+G.4 Numerical Verbalization (Q3)
+We show the full results for the experiment in §6, with Table 31 illustrating the verbalizations in each variant.
+Note that we do not include an exact value verbalization, since it would increase the number of unique text inputs and place extra memory demands.
+The two variants which integrate numerical information into the verbalization dominate the experiment, although the improvement seems to be marginal for some datasets.
+Interestingly, some datasets significantly underperform, with the R27 dataset completely failing the task.
+The addition of the quantile information on top of the bin seems to have limited impact, although marginally winning on the average performance.
+Figure 8 presents a comparison with baseline models, highlighting the importance of incorporating richer numerical verbalizations beyond relying solely on the column name.
+Figure 8: Comparison of normalized scores with 95% CIs for the numerical verbalization experiment, comparing TabSTAR variants against baseline models in classification tasks.
 Table 31: Illustrative verbalization of a numerical feature (Age) for the Q3: Numerical Verbalization experiment.